### v5

# The Small-Gain Condition for Infinite Networks Modeled on $\ell^{\infty}$-Spaces

Abstract
In recent years, attempts have been made to extend nonlinear small-gain theorems for input-to-state stability (ISS) from finite networks to countably infinite networks with finite indegrees. Under specific assumptions about the interconnection gains and the ISS formulation, corresponding infinite-dimensional small-gain results have been proven. However, concerning these assumptions, the results are still too narrow to be considered a full extension of the state-of-the-art for finite networks. We take a step to closing this gap by developing a general technical framework within which the small-gain condition for both finite and infinite networks can be analyzed. This includes a thorough investigation of various monotone operators associated with a network and a specific ISS formulation. Our results extend and generalize the existing theory for finite networks, yield complete characterizations of the small-gain condition for specific ISS formulations, and show which obstacles still have to be overcome to obtain a complete theory for the most general infinite case.
Everyone who builds an infinite-dimensional version of a finite dimensional theory is being punished twice: first, because proofs are very hard, and second, because, on the surface, the final product looks not much different from the original.
A. Blumenthal, Y. Latushkin [5]
Keywords– input-to-state stability; infinite-dimensional systems; small-gain theory; monotone operators
1 Introduction
Given a stabilizing feedback controller for a nonlinear system, it is possible that small deviations from this controller lead to divergent behavior. Since any practical implementation of a controller necessarily deviates from the conceptual controller, this is a serious problem. As a remedy, Eduardo Sontag introduced the concept of input-to-state stability (ISS) in [22]. If the open-loop system obtained by including time-dependent bounded perturbations into the feedback loop is input-to-state-stable, practical stability is guaranteed and the bounded deviation of trajectories from the equilibrium can be estimated quantitatively over an infinite time horizon. As any general property of a system can be hard to verify in a concrete example, it is necessary to find sufficient conditions for ISS which can be checked by computation. Sontag also introduced the notion of an ISS Lyapunov function, whose existence is equivalent to ISS, at least in finite dimensions. This shifts the problem of verifying ISS directly to the construction (or proof of existence) of an ISS Lyapunov function, which is not necessarily easier though, especially for high-dimensional systems.
The small-gain approach tries to mitigate the difficulty by splitting the task of finding a Lyapunov function into two manageable subtasks. The idea is to decompose the system into a number of lower-dimensional, physically interconnected subsystems. Every subsystem has two types of inputs, the external perturbations acting on the overall system and the outputs of the other subsystems. If one can find an ISS Lyapunov function for each subsystem with respect to both types of inputs, there is a chance that these functions can be composed in a certain manner to an overall ISS Lyapunov function. This strategy works if the interconnections are sufficiently weak, which can be formalized in terms of a so-called small-gain condition. The latter can be formulated as a property of a monotone operator $\Gamma$ (called gain operator), built from the interconnection gains and acting on the positive cone of an ordered vector space. In finite dimensions, there are four equivalent ways to describe the small-gain condition: in terms of a path in the positive cone evolving in the decay set of $\Gamma$, in terms of an asymptotic stability property of the dynamical system generated by $\Gamma$, in terms of a monotone bounded invertibility property of the operator $\mathrm{id}-\Gamma$, and in terms of the componentwise action of $\Gamma$ in one time step. The last characterization is particularly useful, since it allows for computational verification in many cases. There is also a connection to the theory of positive linear operators. Indeed, under specific assumptions about the gains and the ISS formulation, $\Gamma$ becomes a positive linear operator and the small-gain condition is equivalent to the spectral radius condition $r(\Gamma)<1$.
An alternative view on the small-gain theory is that the network structure is given a priori by the setup of the control system, which naturally splits into many interacting components that one wants to control simultaneously, and which may be distributed over a large area. Examples can be found, e.g., in automated traffic systems, electric power grids, or aerospace and defense systems. In such applications, the number of subsystems may even vary in time and networks can become very large. As the number of subsystems in a network grows, measures of stability may deteriorate even if ISS is preserved. A successful approach to dealing with a system that consists of a very large number of components is to over-approximate it by some sort of infinite system. This strategy is used, for instance, in physics to model the dynamics of a fluid or a gas, consisting of an enormous number of interacting molecules. Motivated by such ideas, we study the ISS small-gain approach for networks consisting of an infinite number of interacting systems. First results in this direction have been obtained in [7, 12, 13, 14, 16, 17]. However, in all of these papers, the authors either restricted their attention to a narrow class of gains and/or formulations of the ISS property, or did not obtain a full generalization of the existing results for finite networks as presented in [8, 9, 11, 20].
In this paper, we aim at such a generalization. To this end, we introduce a very general class of gain operators associated with a network of countably many interconnected systems. We mainly impose one restriction, namely that each subsystem is only influenced by finitely many other subsystems. For such a setup, the paper [13] already provided a small-gain theorem in which the small-gain condition was formulated in terms of the existence of a so-called path of strict decay. However, the existence of such a path was only investigated for a particular formulation of the ISS property, leading to gain operators with a very convenient algebraic property. In this paper, we aspire much greater generality.
We abstain from an extensive literature review on the control of infinite networks or input-to-state stability of infinite-dimensional systems, since one can find such reviews in the papers cited above. We just mention [2, 3] as two prominent examples, which have connections to the topic of this paper, but address control problems for large-scale or infinite networks rather than mere stability problems. The paper also abstains from the discussion of concrete examples, since they would not provide a valuable contribution to the general theory without a much deeper analysis tailored to those examples, which can be part of future work. Finally, we note that our results can potentially contribute to the general theory of monotone operators on Banach spaces, but this needs some further abstraction, which can also be part of future work. Probably, the best way to describe this paper is to say that it provides a technical framework within which all known Lyapunov-based ISS small-gain results for finite or infinite networks can be explained and connected to each other, and which prepares the ground for more general results.
We now describe the structure of the paper and the contents of each section.
In Section 2, we define gain operators together with the main object needed to construct an overall ISS Lyapunov function from ISS Lyapunov functions of the subsystems, namely a path of strict decay.
[…]
Subsection 2.2: We define gain operators and introduce the notion of a path of strict decay. A gain operator $\Gamma$ acts as a monotone operator on the cone $\ell^{\infty}_{+}(\mathcal{I})$. A path $\sigma$ of strict decay evolves in the decay set $\Psi(\Gamma_{\rho}):=\{s\in\ell^{\infty}_{+}(\mathcal{I}):\Gamma_{\rho}(s)\leq s\}$ of an enlarged gain operator $\Gamma_{\rho}=(\mathrm{id}+\rho)\circ\Gamma$, where $\rho$ is a $\mathcal{K}_{\infty}$-function. By definition, such a path needs to satisfy a bunch of properties, including a local bi-Lipschitz condition. We prove that it is not necessary to show that all of these properties hold for a given candidate path $\sigma$. Under considerably weaker assumptions on $\sigma$, it is possible to obtain the desired path by certain algebraic manipulations performed on $\sigma$. This leads to the notion of a $C^{0}$-path (of strict decay), which is a continuous and increasing path in $\Psi(\Gamma_{\rho})$ satisfying coercivity estimates.
Subsection 2.3: We introduce further monotone operators derived from a gain operator together with the associated dynamical systems. This allows us to characterize the existence of a path with all the properties of a $C^{0}$-path except for continuity.
Subsection 2.4: We investigate stability properties of the dynamical system $\Sigma(\Gamma)$ generated by a gain operator $\Gamma$. We first characterize global attractivity of $\Sigma(\Gamma)$ in the weak-topology. Here, the classical small-gain condition $\Gamma(s)\not\geq s$ for $s>0$, also called no-joint-increase (NJI) condition, is an essential ingredient. We then prove that global attractivity in the norm-topology implies a uniform NJI condition, which requires that for each index $i\in\mathcal{I}$, the componentwise decay condition $\Gamma_{j}(s)<s_{j}$ holds for some index $j=j(i)$, uniformly close to $i$ in the topology of the interconnection graph. For finite networks, this condition is equivalent to the usual NJI condition, and in this sense, it is a proper generalization to the infinite case. We also prove a converse result, showing that the uniform NJI condition for $\Gamma_{\rho}$ together with the cofinality of $\Psi(\Gamma_{\rho})$ in $\ell^{\infty}_{+}(\mathcal{I})$ implies uniform global asymptotic stability (UGAS) of $\Sigma(\Gamma)$. Combining these results, we obtain a strong version of UGAS, which can be characterized in terms of the uniform NJI condition together with the norm-boundedness of the trajectories of an augmented gain operator. For finite networks, everything reduces to well-known results.
In Section 3, we investigate conditions for the existence of a $C^{0}$-path.
Subsection 3.1: We introduce a property of gain operators, called the maximum monotone bounded invertibility ($\oplus$-MBI) property. This property, in particular, implies the uniform cofinality of $\Psi(\Gamma)$ and weak global attractivity of $\Sigma(\Gamma)$. Moreover, it implies that each of the so-called projected gain operators $\bar{\Gamma}_{b}(s):=b\oplus\Gamma(s)$ with $b\in\ell^{\infty}_{+}(\mathcal{I})$ has a minimal and a maximal fixed point $s_{*}(b)$ and $s^{*}(b)$, respectively, such that the order interval $s_{*}(b),s^{*}(b)]$ is a compact global attractor of $\Sigma(\bar{\Gamma}_{b})$ in the weak-topology.
Subsection 3.2: We obtain as further necessary conditions for the existence of a $C^{0}$-path (i) the $\oplus$-MBI property for an enlarged gain operator $\Gamma_{\rho}$ and (ii) UGAS of $\Sigma(\Gamma_{\rho})$. These conditions even follow without the assumption that the path is increasing. Together, the two conditions imply the strong version of UGAS introduced in Subsection 2.4.
Subsection 3.3: We discuss three methods to construct a $C^{0}$-path. The first method considers $\sigma_{*}(r):=s_{*}(r\mathds{1})$, $\sigma_{*}:\mathbb{R}_{+}\rightarrow\ell^{\infty}_{+}(\mathcal{I})$, as a candidate path, which under the $\oplus$-MBI property already satisfies all axioms except for continuity. We show that $\sigma_{*}$ is continuous under the additional assumption that each of the projected gain operators $\bar{\Gamma}_{r\mathds{1}}$ has a globally attractive (hence unique) fixed point. We also obtain a converse result, showing that the continuity of $\sigma_{*}$, under mild technical assumptions, implies the uniqueness of the fixed point for each $\bar{\Gamma}_{r\mathds{1}}$. The second path construction method is based on the existence of a complete orbit $(s^{n})_{n\in\mathbb{Z}}$ of $\Gamma$ and the path is obtained by linear interpolation between the points $s^{n}$. We show that this method works if we assume that $\Sigma(\Gamma)$ is globally attractive and there exists a forward-invariant, cofinal and coercive subset $A\subset\Psi(\Gamma)$. It can easily be seen that the last assumption is also necessary for the interpolation method to work. The third method uses the maximal fixed points $s^{*}(r\mathds{1})$ in combination with local linear interpolation between the points of $\Sigma(\bar{\Gamma}_{r\mathds{1}})$-trajectories. Using this method, we establish that the $\oplus$-MBI property is equivalent to the existence of a path, satisfying all properties of a $C^{0}$-path except for norm-continuity from above, which has to be replaced by sequential weak-continuity. When the index set $\mathcal{I}$ is finite, a complete characterization of the existence of a path of strict decay in terms of the $\oplus$-MBI property follows. Moreover, we can conclude that the $\oplus$-MBI property for $\Gamma_{\rho}$ implies that each finite sub-network admits a path of strict decay.
In Section 4, we apply our results to three classes of gain operators.
Subsection 4.1: We study max-type gain operators, which have the property that they preserve the $\oplus$-operation on $\ell^{\infty}_{+}(\mathcal{I})$, i.e. $\Gamma(s^{1}\oplus s^{2})=\Gamma(s^{1})\oplus\Gamma(s^{2})$. For these operators, we obtain a fairly complete theory. Using the results of subsections 3.2 and 3.3, we can prove that the existence of a path of strict decay is equivalent to UGAS of $\Sigma(\Gamma_{\rho})$ for some $\rho\in\mathcal{K}_{\infty}$. We can also provide a characterization in terms of the uniform NJI condition and an asymptotic condition involving chains of interconnection gains.
Subsection 4.2: We consider gain operators, which are homogeneous, i.e. $\Gamma(\alpha s)=\alpha\Gamma(s)$ for $s\in\ell^{\infty}_{+}(\mathcal{I})$ and $\alpha\geq 0$, and subadditive, i.e. $\Gamma(s^{1}+s^{2})\leq\Gamma(s^{1})+\Gamma(s^{2})$ for $s^{1},s^{2}\in\ell^{\infty}_{+}(\mathcal{I})$. For this class of operators, we show that the existence of a path of strict decay is equivalent to UGAS of $\Sigma(\Gamma)$, which is in turn equivalent to a condition involving the generalized spectral radius of such operators. Moreover, we prove that under the restriction to linear gain functions for building enlarged gain operators also the $\oplus$-MBI property is an equivalent condition.
Subsection 4.3: We consider the case when the index set $\mathcal{I}$ is finite. Here, we recover and improve known results from the literature. The improvement mainly consists in the fact that the known results require a subadditivity property of the functions defining the ISS formulation. We prove our results assuming the considerably weaker property of uniform continuity.
In Section 5, we summarize the main contributions of the paper and present a list of open questions and problems for future work. Finally, the appendix, Section A, contains some useful technical lemmas.
2 Gain operators
In this section, we introduce the central objects of the paper and study their fundamental properties.
Notation: We write $\mathbb{N}=\{1,2,3,\ldots\}$ for the natural numbers, $\mathbb{N}_{0}=\{0\}\cup\mathbb{N}$, and $\mathbb{Z}$ for the integers. If $a,b\in\mathbb{Z}$ with $a\leq b$, we write $a;b]=\{n\in\mathbb{Z}:a\leq n\leq b\}$. We further write $\mathbb{R}$ for the reals and $\mathbb{R}_{+}=\{r\in\mathbb{R}:r\geq 0\}$. If $A$ is a finite set, $|A|$ denotes its cardinality. If $X,Y$ are topological spaces, we use the notation $C^{0}(X,Y)$ for the set of continuous mappings from $X$ into $Y$. For any $r\in\mathbb{R}$, we write $\lfloor r\rfloor=\max\{n\in\mathbb{Z}:n\leq r\}$ and $\lceil r\rceil=\min\{n\in\mathbb{Z}:r\leq n\}$.
2.1 Preliminaries
The spaces of $\mathcal{K}$-functions and $\mathcal{K}_{\infty}$-functions, respectively, are defined by
$\mathcal{K}$ $:=\left\{\gamma\in C^{0}(\mathbb{R}_{+},\mathbb{R}_{+}):\gamma(0)=0\mbox{ and }\gamma\mbox{ is strictly increasing}\right\},$
$\mathcal{K}_{\infty}$ $:=\left\{\gamma\in\mathcal{K}:\lim_{r\rightarrow\infty}\gamma(r)=\infty\right\}.$
We further introduce the classes of $\mathcal{L}$-functions and $\mathcal{K}\mathcal{L}$-functions:
$\mathcal{L}$ $:=\left\{\gamma\in C^{0}(\mathbb{R}_{+},\mathbb{R}_{+}):\gamma\mbox{ is strictly decreasing with }\lim_{t\rightarrow\infty}\gamma(t)=0\right\},$
$\mathcal{K}\mathcal{L}$ $:=\left\{\beta\in C^{0}(\mathbb{R}_{+}\times\mathbb{R}_{+},\mathbb{R}_{+}):\beta(\cdot,t)\in\mathcal{K},\ \forall t\geq 0,\ \beta(r,\cdot)\in\mathcal{L},\ \forall r>0\right\}.$
[…]
$\ell^{\infty}(\mathcal{I})$ $:=\bigl\{s=(s_{i})_{i\in\mathcal{I}}:s_{i}\in\mathbb{R},\ \sup_{i\in\mathcal{I}}|s_{i}|<\infty\bigr\},$
$\ell^{\infty}_{+}(\mathcal{I})$ $:=\left\{s\in\ell^{\infty}(\mathcal{I}):s_{i}\geq 0,\ \forall i\in\mathcal{I}\right\}.$
The norm on $\ell^{\infty}(\mathcal{I})$ is the usual norm of uniform convergence:
$\|s\|:=\sup_{i\in\mathcal{I}}|s_{i}|.$
The cone $\ell^{\infty}_{+}(\mathcal{I})$ is closed, convex, and has nonempty interior given by
$\mathrm{int}(\ell^{\infty}_{+}(\mathcal{I}))=\bigl\{s\in\ell^{\infty}_{+}(\mathcal{I})\ :\ \exists\alpha>0\mbox{ with }\inf_{i\in\mathcal{I}}s_{i}\geq\alpha\bigr\}.$
The partial order on $\ell^{\infty}(\mathcal{I})$, induced by the cone $\ell^{\infty}_{+}(\mathcal{I})$, is given by
$s^{1}\leq s^{2}\quad:\Leftrightarrow\quad s^{2}-s^{1}\in\ell^{\infty}_{+}(\mathcal{I}).$
Hence, $s\in\ell^{\infty}_{+}(\mathcal{I})$ if and only if $s\geq 0$. We write $s^{1}<s^{2}$ if $s^{1}\leq s^{2}$ and $s^{1}\neq s^{2}$. We further write $s^{1}\ll s^{2}$ if $s^{2}-s^{1}\in\mathrm{int}(\ell^{\infty}_{+}(\mathcal{I}))$. For two subsets $A,B\subset\ell^{\infty}_{+}(\mathcal{I})$, we write $A\leq B$ if $a\leq b$ for all $(a,b)\in A\times B$. For any $0\leq s^{1}\leq s^{2}$, we define the order interval $s^{1},s^{2}]:=\{s\in\ell^{\infty}_{+}(\mathcal{I}):s^{1}\leq s\leq s^{2}\}$.
The Banach space $\ell^{\infty}(\mathcal{I})$ can be identified with the topological dual of $\ell^{1}(\mathcal{I})$ (the space of absolutely summable sequences). In this identification, any $s\in\ell^{\infty}(\mathcal{I})$ acts linearly on elements of $\ell^{1}(\mathcal{I})$ by
$s(t):=\sum_{i\in\mathcal{I}}s_{i}t_{i}\mbox{\quad for all\ }t=(t_{i})_{i\in\mathcal{I}}\in\ell^{1}(\mathcal{I}).$ (1)
On every topological dual space $X^{*}$ of a normed space $X$, we have the weak-topology which is the coarsest topology such that for every $x\in X$ the linear functional $x^{*}\mapsto x^{*}(x)$, defined on $X^{*}$, is continuous. If a sequence $(s^{n})_{n\in\mathbb{N}}$ in $\ell^{\infty}(\mathcal{I})$ converges in the weak-topology to some $s$, we write $s^{n}\stackrel{{\scriptstyle\star}}{{\rightharpoonup}}s$. By Lemma A.1, this is the case if and only if $(s^{n})_{n\in\mathbb{N}}$ is norm-bounded and converges to $s$ componentwise. The weak-topology on $\ell^{\infty}(\mathcal{I})$ coincides with the norm-topology if and only if $\mathcal{I}$ is finite.
An operator $T:\ell^{\infty}_{+}(\mathcal{I})\rightarrow\ell^{\infty}_{+}(\mathcal{I})$ is called monotone if $s^{1}\leq s^{2}$ implies $T(s^{1})\leq T(s^{2})$ for all $s^{1},s^{2}\in\ell^{\infty}_{+}(\mathcal{I})$. The discrete-time dynamical system generated by $T$ is given by
$\Sigma(T):\quad s^{n+1}=T(s^{n}),\quad s^{0}\in\ell^{\infty}_{+}(\mathcal{I}),\ n\in\mathbb{N}_{0}.$
For any $s\in\ell^{\infty}_{+}(\mathcal{I})$, the sequence $(T^{n}(s))_{n\in\mathbb{N}_{0}}$ is called the trajectory of $\Sigma(T)$ starting in $s$. Under the assumption that $T(0)=0$, we call $\Sigma(T)$
  • uniformly globally stable (UGS) if there exists $\varphi\in\mathcal{K}_{\infty}$ such that $\|T^{n}(s)\|\leq\varphi(\|s\|)$ for all $s\in\ell^{\infty}_{+}(\mathcal{I})$ and $n\in\mathbb{N}_{0}$.
  • uniformly globally asymptotically stable (UGAS) if there exists $\beta\in\mathcal{K}\mathcal{L}$ such that $\|T^{n}(s)\|\leq\beta(\|s\|,n)$ for all $s\in\ell^{\infty}_{+}(\mathcal{I})$ and $n\in\mathbb{N}_{0}$.
  • globally attractive (GATT) if every trajectory of $\Sigma(T)$ converges to the origin in the norm-topology.
  • uniformly globally attractive (UGATT) if for all $r,\varepsilon>0$ there exists $n\in\mathbb{N}$ such that
    $\|s\|\leq r\mbox{\ and\ }k\geq n\quad\Rightarrow\quad\|T^{k}(s)\|\leq\varepsilon.$
[…]
It is clear that UGAS $\Rightarrow$ UGATT $\Rightarrow$ GATT $\Rightarrow$ GATT and UGAS $\Rightarrow$ UGS. From Lemma A.8, it follows that UGAS $\Leftrightarrow$ UGS $\wedge$ GATT. We say that the trajectories of $\Sigma(T)$ are norm-bounded if for every $s\in\ell^{\infty}_{+}(\mathcal{I})$ there exists $M>0$ with $\|T^{n}(s)\|\leq M$ for all $n\in\mathbb{N}_{0}$. This property is implied by all of the above properties.
A point $s\in\ell^{\infty}_{+}(\mathcal{I})$ is called a point of decay for $T$ if $T(s)\leq s$ and a fixed point if $T(s)=s$. The set of decay $\Psi(T)$ and the fixed point set $\mathrm{Fix}(T)$ of $T$ are defined as the sets of all points of decay and fixed points, respectively.
By $e^{i}$, $i\in\mathcal{I}$, we denote the unit vectors in $\ell^{\infty}(\mathcal{I})$, i.e. $e^{i}_{j}=\delta_{ij}$ with the Kronecker delta $\delta_{ij}$. By $\mathds{1}$, we denote the vector all of whose components are $1$s, i.e. $\mathds{1}=\sum_{i\in\mathcal{I}}e^{i}$.
A natural binary operation on $\ell^{\infty}_{+}(\mathcal{I})$ is the componentwise maximum,
$s^{1}\oplus s^{2}:=\left(\max\{s^{1}_{i},s^{2}_{i}\}\right)_{i\in\mathcal{I}}\mbox{\quad for all\ }s^{1},s^{2}\in\ell^{\infty}_{+}(\mathcal{I}).$
It can be extended to arbitrary norm-bounded families of vectors $s^{a}$, $a\in A$, by
$\bigoplus_{a\in A}s^{a}:=\Bigl(\sup_{a\in A}s^{a}_{i}\Bigr)_{i\in\mathcal{I}}.$
A set $A\subset\ell^{\infty}_{+}(\mathcal{I})$ is called cofinal if for every $s\in\ell^{\infty}_{+}(\mathcal{I})$ there is $\hat{s}\in A$ such that $s\leq\hat{s}$. It is called uniformly cofinal if there exists $\varphi\in\mathcal{K}_{\infty}$ such that for every $s\in\ell^{\infty}_{+}(\mathcal{I})$ there is $\hat{s}\in A$ with $s\leq\hat{s}$ and $\|\hat{s}\|\leq\varphi(\|s\|)$. A characterization of uniform cofinality is provided by Lemma A.10.
A sequence $(s^{n})_{n\in\mathbb{N}}$ in $\ell^{\infty}_{+}(\mathcal{I})$ is called increasing if $s^{n}\leq s^{n+1}$ and decreasing if $s^{n+1}\leq s^{n}$ for all $n\in\mathbb{N}$.
We call a set $A\subset\ell^{\infty}_{+}(\mathcal{I})$ coercive if there exists $\underline{\varphi}\in\mathcal{K}_{\infty}$ such that $s\geq\underline{\varphi}(\|s\|)\mathds{1}$ for all $s\in A$. This inequality can also be written as
$\inf_{i\in\mathcal{I}}s_{i}\geq\underline{\varphi}(\sup_{i\in\mathcal{I}}s_{i}).$
Hence, a set $A$ is coercive if the variation in the components of each $s\in A$ is limited by a $\mathcal{K}_{\infty}$-function.
For any subset $\mathcal{J}\subset\mathcal{I}$, we define
$\ell^{\infty}_{+}(\mathcal{I},\mathcal{J}):=\{s\in\ell^{\infty}_{+}(\mathcal{I}):s_{i}=0,\ \forall i\in\mathcal{I}\backslash\mathcal{J}\}$
which can be identified naturally and isometrically with $\ell^{\infty}_{+}(\mathcal{J})$. For every $s\in\ell^{\infty}_{+}(\mathcal{I})$, we let
$s_{|\mathcal{J}}:=\sum_{i\in\mathcal{J}}s_{i}e^{i}\in\ell^{\infty}_{+}(\mathcal{I},\mathcal{J}).$
2.2 Definition and elementary properties
We now define gain operators as specific monotone operators on $\ell^{\infty}_{+}(\mathcal{I})$. A gain operator is associated with an infinite network of dynamical systems, indexed by $\mathcal{I}$, where each system is directly influenced by only finitely many neighbors. The set of indices of the neighbors of system $i$ is denoted by $\mathcal{I}_{i}\subset\mathcal{I}$. Here, we assume that $i\notin\mathcal{I}_{i}$ and do not exclude the case $\mathcal{I}_{i}=\emptyset$. We also define an associated interconnection graph $\mathcal{G}$, which is a directed graph with vertex set $V(\mathcal{G})$ and edge set $E(\mathcal{G})$ given by
[…]
respectively. For a vertex $i$ and a positive integer $n\in\mathbb{N}$, we denote by $\mathcal{N}^{-}_{i}(n)$ the set of all vertices $j$ such that there exists a directed path from $j$ to $i$ of length at most $n$. In contrast, the set $\mathcal{N}^{+}_{i}(n)$ is the set of all vertices $j$ such that there exists a directed path from $i$ to $j$ of length at most $n$. In addition, we put $\mathcal{N}^{-}_{i}(0)=\mathcal{N}^{+}_{i}(0):=\{i\}$ and define
$\mathcal{N}^{-}_{i}:=\bigcup_{n\in\mathbb{N}_{0}}\mathcal{N}^{-}_{i}(n),\quad\mathcal{N}^{+}_{i}:=\bigcup_{n\in\mathbb{N}_{0}}\mathcal{N}^{+}_{i}(n).$
Observe that $\mathcal{N}^{-}_{i}(1)=\{i\}\cup\mathcal{I}_{i}$. Inductively, it follows that the sets $\mathcal{N}^{-}_{i}(n)$ are all finite. In contrast, the sets $\mathcal{N}^{+}_{i}(n)$ may be infinite.
A gain operator compatible with the directed graph $\mathcal{G}$ is built from two families of functions. The first family consists of functions $\gamma_{ij}\in\mathcal{K}_{\infty}$, $ji\in E(\mathcal{G})$, called interconnection gains. We assume that this family is pointwise equicontinuous, i.e. for all $r_{0}\in\mathbb{R}_{+}$ and $\varepsilon>0$ there exists $\delta>0$ such that $|r-r_{0}|\leq\delta$ implies $|\gamma_{ij}(r)-\gamma_{ij}(r_{0})|\leq\varepsilon$ for all $ji\in E(\mathcal{G})$ and $r\in\mathbb{R}_{+}$.
The second family has elements $\mu_{i}:\ell^{\infty}_{+}(\mathcal{I})\rightarrow[0,\infty]$, $i\in\mathcal{I}$, called monotone aggregation functions (MAFs), [^3] This term is adopted from the small-gain literature. satisfying the following properties:
  1. (M1)
    Uniform positive definiteness: There exists $\xi\in\mathcal{K}_{\infty}$ such that $\mu_{i}(0)=0$ and $\mu_{i}(s)\geq\xi(\|s\|)$ for all $s\in\ell^{\infty}_{+}(\mathcal{I})$ and $i\in\mathcal{I}$.
  2. (M2)
    Monotonicity: $0\leq s^{1}\leq s^{2}$ implies $\mu_{i}(s^{1})\leq\mu_{i}(s^{2})$ for all $i\in\mathcal{I}$.
  3. (M3)
    Finite continuity: For any finite set $\mathcal{J}\subset\mathcal{I}$ and $i\in\mathcal{I}$, the restriction of $\mu_{i}$ to $\ell^{\infty}_{+}(\mathcal{I},\mathcal{J})$ is finite-valued and continuous.
  4. (M4)
    Equicontinuity: For every norm-bounded set $A\subset\ell^{\infty}_{+}(\mathcal{I})$ and $\varepsilon>0$, there is $\delta>0$ such that for all $s^{0}\in A$ and $s\in\ell^{\infty}_{+}(\mathcal{I})$:
    $\|s-s^{0}\|\leq\delta\quad\Rightarrow\quad\sup_{i\in\mathcal{I}}|\mu_{i}(s_{|\mathcal{I}_{i}})-\mu_{i}(s^{0}_{|\mathcal{I}_{i}})|\leq\varepsilon.$
We now define gain operators.
Definition 2.1.
The gain operator $\Gamma:\ell^{\infty}_{+}(\mathcal{I})\rightarrow\ell^{\infty}_{+}(\mathcal{I})$ associated with the graph $\mathcal{G}$, the interconnection gains $\gamma_{ij}$, and the MAFs $\mu_{i}$, is defined by
$\Gamma(s)=(\Gamma_{i}(s))_{i\in\mathcal{I}},\quad\Gamma_{i}(s):=\mu_{i}([\gamma_{ij}(s_{j})]_{j\in\mathcal{I}_{i}}),$
where we use the natural identification of $\ell^{\infty}_{+}(\mathcal{I}_{i})$ with $\ell^{\infty}_{+}(\mathcal{I},\mathcal{I}_{i})$.
In the rest of the paper, we speak of the finite case when we consider gain operators with $|\mathcal{I}|<\infty$ and otherwise of the infinite case.
Example 2.1.
We consider two of the most relevant examples for MAFs. Let
$\mu_{\oplus}(s):=\sup_{i\in\mathcal{I}}s_{i}=\|s\|,\quad\mu_{\oplus}:\ell^{\infty}_{+}(\mathcal{I})\rightarrow[0,\infty].$
In case that $\mu_{i}\equiv\mu_{\oplus}$, we call $\Gamma$ a max-type operator. Here, (M1) holds with $\xi=\mathrm{id}$ and (M2), (M3) are trivial. If $\|s-s^{0}\|\leq\delta$ for some $s^{0},s\in\ell^{\infty}_{+}(\mathcal{I})$, the triangle inequality from below implies
$\sup_{i\in\mathcal{I}}\Bigl|\sup_{j\in\mathcal{I}_{i}}s_{j}-\sup_{j\in\mathcal{I}_{i}}s^{0}_{j}\Bigr|\leq\sup_{ji\in E(\mathcal{G})}|s_{j}-s^{0}_{j}|\leq\|s-s^{0}\|\leq\delta.$
Hence, also (M4) is satisfied, even on the whole cone $\ell^{\infty}_{+}(\mathcal{I})$ and not only on bounded subsets. We thus see that for max-type operators the only requirement is that the interconnection gains form a pointwise equicontinuous family. Now, consider
$\mu_{\otimes}(s):=\sum_{i\in\mathcal{I}}s_{i},\quad\mu_{\otimes}:\ell^{\infty}_{+}(\mathcal{I})\rightarrow[0,\infty].$
In case that $\mu_{i}\equiv\mu_{\otimes}$, we call $\Gamma$ a sum-type operator. Here, (M1) again holds with $\xi=\mathrm{id}$ and (M2), (M3) are trivial. Concerning (M4), $\|s-s^{0}\|\leq\delta$ implies
$\Bigl|\sum_{j\in\mathcal{I}_{i}}s_{j}-\sum_{j\in\mathcal{I}_{i}}s^{0}_{j}\Bigr|\leq\sum_{j\in\mathcal{I}_{i}}|s_{j}-s^{0}_{j}|\leq\delta\cdot|\mathcal{I}_{i}|,$
which can hold with equality in certain cases. Hence, for (M4) it is necessary and sufficient that the cardinality $|\mathcal{I}_{i}|$ is bounded over $i\in\mathcal{I}$.
Elementary properties of gain operators are summarized in the following result.
Proposition 2.1.
Any gain operator $\Gamma$ has the following properties:
  1. (a)
    $\Gamma$ is well-defined, i.e. $\Gamma(s)\in\ell^{\infty}_{+}(\mathcal{I})$ for all $s\in\ell^{\infty}_{+}(\mathcal{I})$.
  2. (b)
    $\Gamma(0)=0$ and $\Gamma$ is a monotone operator.
  3. (c)
    $\Gamma$ is continuous and uniformly continuous on bounded sets.
  4. (d)
    $\Gamma$ is sequentially continuous in the weak-topology.
Proof.
First, consider the family $\mathcal{F}:=\{f_{i}:i\in\mathcal{I}\}\subset C^{0}(\mathbb{R}_{+},\mathbb{R}_{+})$ with $f_{i}(r):\equiv\mu_{i}(r\mathds{1}_{|\mathcal{I}_{i}})$. By (M3), each $f_{i}$ is finite-valued and continuous. We claim that $\mathcal{F}$ is uniformly bounded on every compact interval $0,r_{0}]$. To show this, we use (M4) with $\varepsilon=1$, which gives $\delta>0$ such that $|r-r^{\prime}|\leq\delta$ with $r,r^{\prime}\in[0,r_{0}]$ implies
$\sup_{i\in\mathcal{I}}|f_{i}(r^{\prime})-f_{i}(r)|=\sup_{i\in\mathcal{I}}\left|\mu_{i}\left(r^{\prime}\mathds{1}_{|\mathcal{I}_{i}}\right)-\mu_{i}\left(r\mathds{1}_{|\mathcal{I}_{i}}\right)\right|\leq 1,$
leading to $f_{i}(r_{0})=|f_{i}(r_{0})-f_{i}(0)|\leq\lceil r_{0}/\delta\rceil$ for all $i\in\mathcal{I}$.
Now, we show that $\Gamma$ is well-defined. To this end, let $s\in\ell^{\infty}_{+}(\mathcal{I})$ and put $r:=\|s\|$. Then, by the monotonicity of each $\mu_{i}$ and $\gamma_{ij}$, we have
$\Gamma_{i}(s)=\mu_{i}([\gamma_{ij}(s_{j})]_{j\in\mathcal{I}_{i}})\leq\mu_{i}([\gamma_{ij}(r)]_{j\in\mathcal{I}_{i}})\mbox{\quad for all\ }i\in\mathcal{I}.$
With the same arguments as used above for the family $\mathcal{F}$, the equicontinuity of $\{\gamma_{ij}\}$ implies the existence of $R>0$ with $\gamma_{ij}(r)\leq R$ for all $ji\in E(\mathcal{G})$. We conclude the existence of $B>0$ with
$\|\Gamma(s)\|=\sup_{i\in\mathcal{I}}\Gamma_{i}(s)\leq\sup_{i\in\mathcal{I}}\mu_{i}\left(R\mathds{1}_{|\mathcal{I}_{i}}\right)=\sup_{i\in\mathcal{I}}f_{i}(R)\leq B,$
showing that $\Gamma$ is well-defined. From the monotonicity assumptions on $\gamma_{ij}$ and $\mu_{i}$, it follows that $\Gamma$ is a monotone operator. By (M1), it is clear that $\Gamma(0)=0$. Hence, (a) and (b) are proven.
We now prove the uniform continuity of $\Gamma$ on bounded sets. To this end, let $B\subset\ell^{\infty}_{+}(\mathcal{I})$ be bounded, say $\|s\|\leq r$ for all $s\in B$. Then also the set
$A:=\left\{(\gamma_{ij}(s_{j}))_{j\in\mathcal{I}_{i}}\in\ell^{\infty}_{+}(\mathcal{I},\mathcal{I}_{i})\ :\ s\in B,\ i\in\mathcal{I}\right\}$
is bounded, because the equicontinuity of $\{\gamma_{ij}\}$ implies $\gamma_{ij}(s_{j})\leq\gamma_{ij}(r)\leq R$ for all $ji\in E(\mathcal{G})$ and some $R>0$. For a given $\varepsilon>0$, we first choose $\delta=\delta(A,\varepsilon)$ according to (M4). We subsequently choose $\rho>0$ so that $\|s^{1}-s^{2}\|\leq\rho$ with $s^{1},s^{2}\in B$ implies $|\gamma_{ij}(s^{1}_{j})-\gamma_{ij}(s^{2}_{j})|\leq\delta$ for all $ji\in E(\mathcal{G})$, which is possible by the pointwise equicontinuity of $\{\gamma_{ij}\}$ (which is uniform on compact intervals). Altogether, $\|s^{1}-s^{2}\|\leq\rho$ with $s^{1},s^{2}\in B$ implies
$\|\Gamma(s^{1})-\Gamma(s^{2})\|=\sup_{i\in\mathcal{I}}\left|\mu_{i}([\gamma_{ij}(s^{1}_{j})]_{j\in\mathcal{I}_{i}})-\mu_{i}([\gamma_{ij}(s^{2}_{j})]_{j\in\mathcal{I}_{i}})\right|\leq\varepsilon.$
Hence, (c) is proved. To prove (d), consider a sequence $(s^{n})_{n\in\mathbb{N}}$ with $s^{n}\stackrel{{\scriptstyle\star}}{{\rightharpoonup}}s$ for some $s$. By Lemma A.1, this means that $(s^{n})$ is norm-bounded and converges componentwise to $s$. Then, for every $i\in\mathcal{I}$ we have
$\lim_{n\rightarrow\infty}\Gamma_{i}(s^{n})$ $=\lim_{n\rightarrow\infty}\mu_{i}([\gamma_{ij}(s^{n}_{j})]_{j\in\mathcal{I}_{i}})=\mu_{i}\left(\lim_{n\rightarrow\infty}[\gamma_{ij}(s^{n}_{j})]_{j\in\mathcal{I}_{i}}\right)$
$=\mu_{i}\Bigl(\Bigl[\lim_{n\rightarrow\infty}\gamma_{ij}(s^{n}_{j})\Bigr]_{j\in\mathcal{I}_{i}}\Bigr)=\mu_{i}\Bigl(\Bigl[\gamma_{ij}\Bigl(\lim_{n\rightarrow\infty}s^{n}_{j}\Bigr)\Bigr]_{j\in\mathcal{I}_{i}}\Bigr)$
$=\mu_{i}\left(\left[\gamma_{ij}(s_{j})\right]_{j\in\mathcal{I}_{i}}\right)=\Gamma_{i}(s).$
Here, the second equality follows from (M3). The third equality follows from finiteness of $\mathcal{I}_{i}$, and the fourth from continuity of $\gamma_{ij}$. Hence, $(\Gamma(s^{n}))_{n\in\mathbb{N}}$ converges to $\Gamma(s)$ componentwise. Moreover, if $B>0$ is a bound on $\|s^{n}\|$, then $s^{n}\leq B\mathds{1}$ and thus $\Gamma(s^{n})\leq\Gamma(B\mathds{1})$ due to monotonicity, implying $\|\Gamma(s^{n})\|\leq\|\Gamma(B\mathds{1})\|$ for all $n$. Consequently, $\Gamma(s^{n})\stackrel{{\scriptstyle\star}}{{\rightharpoonup}}\Gamma(s)$. ∎
Remark 2.1.
The sequential continuity of $\Gamma$ in the weak-topology has the following important implication: If a norm-bounded trajectory $(\Gamma^{n}(s))_{n\in\mathbb{N}}$ converges componentwise to $s^{*}\in\ell^{\infty}_{+}(\mathcal{I})$, then $s^{*}$ is a fixed point of $\Gamma$. This will be used many times throughout the paper.
Sometimes, it is useful to bound $\Gamma(s)$ from below. [^4] In the finite-dimensional theory, for this purpose sometimes an unboundedness assumption is imposed on $\Gamma$, namely $\|\Gamma(s)\|\rightarrow\infty$ if $\|s\|\rightarrow\infty$, see [20, Prop. 5.4]. If we require that all $\gamma_{ij}$ are bounded from below by some $\eta\in\mathcal{K}_{\infty}$, and additionally that $\mathcal{I}_{i}\neq\emptyset$ for all $i\in\mathcal{I}$, then (M1) implies that for all $s\in\ell^{\infty}_{+}(\mathcal{I})$ and $ki\in E(\mathcal{G})$
$\Gamma_{i}(s)=\mu_{i}([\gamma_{ij}(s_{j})]_{j\in\mathcal{I}_{i}})\geq\xi(\|[\gamma_{ij}(s_{j})]_{j\in\mathcal{I}_{i}}\|)\geq\xi\circ\eta(s_{k}).$ (2)
For easy reference, we will give this assumption a number.
Assumption 2.1.
There is $\eta\in\mathcal{K}_{\infty}$ such that $\gamma_{ij}\geq\eta$ for all $ji\in E(\mathcal{G})$ and, in addition, $\mathcal{I}_{i}\neq\emptyset$ for all $i\in\mathcal{I}$.
Any function $\varphi\in\mathcal{K}_{\infty}$ acts as a continuous monotone operator on $\ell^{\infty}_{+}(\mathcal{I})$ by
$\varphi(s):=(\varphi(s_{i}))_{i\in\mathcal{I}},\quad\varphi:\ell^{\infty}_{+}(\mathcal{I})\rightarrow\ell^{\infty}_{+}(\mathcal{I}).$
This operator has the following properties:
  • It is invertible with inverse $\varphi^{-1}(s)=(\varphi^{-1}(s_{i}))_{i\in\mathcal{I}}$.
  • $\|\varphi(s)\|=\varphi(\|s\|)$ for all $s\in\ell^{\infty}_{+}(\mathcal{I})$.
  • $\varphi(s^{1}\oplus s^{2})=\varphi(s^{1})\oplus\varphi(s^{2})$ for all $s^{1},s^{2}\in\ell^{\infty}_{+}(\mathcal{I})$.
Given a gain operator $\Gamma$ and $\rho\in\mathcal{K}_{\infty}$, we define two enlarged gain operators
[…]
It is not hard to see that $\Gamma_{\rho}$ is the gain operator associated with the gains $\{\gamma_{ij}\}$ and MAFs $\{(\mathrm{id}+\rho)\circ\mu_{i}\}$, [^5] To prove that this family satisfies (M4), one can use that the $\mu_{i}$ are uniformly continuous on bounded sets and $\mathrm{id}+\rho$ is uniformly continuous on compact intervals. while $\Gamma^{\rho}$ is the gain operator associated with the gains $\{\gamma_{ij}\circ(\mathrm{id}+\rho)\}$ and MAFs $\{\mu_{i}\}$. The operators $\Gamma_{\rho}$ and $\Gamma^{\rho}$ are, in fact, conjugate, since
$(\mathrm{id}+\rho)\circ\Gamma^{\rho}=\Gamma_{\rho}\circ(\mathrm{id}+\rho).$ (3)
A mapping $\sigma:\mathbb{R}_{+}\rightarrow\ell^{\infty}_{+}(\mathcal{I})$ is called a path in the cone $\ell^{\infty}_{+}(\mathcal{I})$. A path $\sigma$ is called increasing if $r_{1}\leq r_{2}$ implies $\sigma(r_{1})\leq\sigma(r_{2})$, and strictly increasing if $r_{1}<r_{2}$ implies $\sigma(r_{1})\ll\sigma(r_{2})$ for all $r_{1},r_{2}\in\mathbb{R}_{+}$.
We now introduce the notion of a path of strict decay (which is adopted from [13, Def. II.10]). The existence of a path of strict decay for the gain operator $\Gamma$, together with other assumptions, guarantees that an ISS Lyapunov function for the overall network can be constructed from ISS Lyapunov functions of the subsystems, which is described in detail in [14, Sec. 2].
Definition 2.2.
Given a gain operator $\Gamma$, a path $\sigma$ in the cone $\ell^{\infty}_{+}(\mathcal{I})$ is called a path of strict decay for $\Gamma$ if it satisfies the following properties:
  1. (i)
    There exists $\rho\in\mathcal{K}_{\infty}$ such that $\Gamma_{\rho}(\sigma(r))\leq\sigma(r)$ for all $r\geq 0$.
  2. (ii)
    There exist $\varphi_{\min},\varphi_{\max}\in\mathcal{K}_{\infty}$ such that $\varphi_{\min}(r)\mathds{1}\leq\sigma(r)\leq\varphi_{\max}(r)\mathds{1}$ for all $r\geq 0$.
  3. (iii)
    Each component function $\sigma_{i}$, $i\in\mathcal{I}$, is a $\mathcal{K}_{\infty}$-function.
  4. (iv)
    For every compact interval $K\subset(0,\infty)$, there exist constants $0<l\leq L$ such that for all $r_{1},r_{2}\in K$ and $i\in\mathcal{I}$
    $l|r_{1}-r_{2}|\leq|\sigma_{i}^{-1}(r_{1})-\sigma_{i}^{-1}(r_{2})|\leq L|r_{1}-r_{2}|.$
If all of the above properties hold with the exception that $\rho=0$ in (i), we call $\sigma$ a path of decay for $\Gamma$.
Remark 2.2.
Observe that a path of strict decay for $\Gamma$ is a path of decay for $\Gamma_{\rho}$ with $\rho$ as in (i). Using Proposition 2.2(a) below, it is easy to see that any path of decay is continuous, locally Lipschitz continuous on $(0,\infty)$, strictly increasing, and satisfies $\sigma(0)=0$.
The definition can be slightly simplified, using the next result.
Proposition 2.2.
Regarding Definition 2.2, the following holds:
  1. (a)
    In property (iv), the functions $\sigma_{i}^{-1}$ can be replaced by $\sigma_{i}$.
  2. (b)
    In property (ii), we can assume without loss of generality that $\varphi_{\min}=\mathrm{id}$.
Proof.
(a) Assume that for every compact interval $L\subset(0,\infty)$ there are constants $0<c\leq C$ such that
$c|r_{1}-r_{2}|\leq|\sigma_{i}(r_{1})-\sigma_{i}(r_{2})|\leq C|r_{1}-r_{2}|$
for all $i\in\mathcal{I}$ and $r_{1},r_{2}\in L$. If $K=[a,b]\subset(0,\infty)$ is any compact interval, then
$\sigma_{i}^{-1}([a,b])=[\sigma_{i}^{-1}(a),\sigma_{i}^{-1}(b)]\subset[\varphi_{\max}^{-1}(a),\varphi_{\min}^{-1}(b)]=:L\subset(0,\infty).$
Hence, with the corresponding constants $0<c\leq C$, we obtain
$\frac{1}{C}|r_{1}-r_{2}|\leq|\sigma_{i}^{-1}(r_{1})-\sigma_{i}^{-1}(r_{2})|\leq\frac{1}{c}|r_{1}-r_{2}|\mbox{\quad for all\ }r_{1},r_{2}\in K.$
(b) Given a path of decay $\sigma$, we can choose $\tilde{\varphi}_{\min}\in\mathcal{K}_{\infty}$ such that $\tilde{\varphi}_{\min}\leq\varphi_{\min}$ and both $\tilde{\varphi}_{\min}$, $\tilde{\varphi}_{\min}^{-1}$ are locally Lipschitz continuous on $(0,\infty)$; for instance, we can choose $z_{1}$ from Lemma A.7 with $z_{1}\leq\varphi_{\min}$ and put $\tilde{\varphi}_{\min}(r):\equiv z_{1}(r)(1-\mathrm{e}^{-r})$. This $\mathcal{K}_{\infty}$-function is smooth on $(0,\infty)$ and its derivative vanishes nowhere, implying that also its inverse is smooth on $(0,\infty)$. Then, we consider the reparametrized path $\tilde{\sigma}:=\sigma\circ\tilde{\varphi}_{\min}^{-1}$. Clearly, $\tilde{\sigma}$ satisfies properties (i) and (iii). Moreover, by construction, $r\mathds{1}\leq\tilde{\sigma}(r)\leq\varphi_{\max}\circ\tilde{\varphi}_{\min}^{-1}(r)\mathds{1}$ for all $r\geq 0$. By the choice of the reparametrization, $\tilde{\sigma}$ also satisfies property (iv), and thus is a path of strict decay for $\Gamma$. ∎
The following proposition is the main result of this subsection. With this result at hand, finding a path of strict decay becomes a considerably easier task.
Proposition 2.3.
Given a gain operator $\Gamma$, assume that there exists a continuous and increasing path $\sigma$ in $\ell^{\infty}_{+}(\mathcal{I})$, satisfying properties (i) and (ii) in Definition 2.2. Then there exists a path of strict decay for $\Gamma$.
Proof.
The proof consists of two steps. In the first step, we assume that $\sigma$ is strictly increasing, while in the second step we weaken this assumption.
Step 1: Let $\sigma$ be a path satisfying (i) and (ii) in Definition 2.2, which is continuous and strictly increasing. Consider a compact interval $a,b]\subset(0,\infty)$. We show that the restriction of $\sigma$ to $[a,b]$ can be approximated by a piecewise linear path evolving in $\Psi(\Gamma)$. The fact that such an approximation is possible will yield a piecewise linear path of decay. By Lemma A.6, there exists $\eta\in\mathcal{K}_{\infty}$ with $(\mathrm{id}+\rho)^{-1}=\mathrm{id}-\eta$. By assumption, $\sigma$ is continuous, hence uniformly continuous on the compact interval $[a,b]$. Thus, we can find $\delta>0$ such that $|r_{1}-r_{2}|<\delta$ implies $\|\sigma(r_{1})-\sigma(r_{2})\|<\varepsilon:=\eta(\varphi_{\min}(a))$, whenever $r_{1},r_{2}\in[a,b]$. Now, let $s^{\alpha}:=(1-\alpha)\sigma(r_{1})+\alpha\sigma(r_{2})$ for $\alpha\in[0,1]$ and $r_{1},r_{2}\in[a,b]$ with $r_{1}<r_{2}<r_{1}+\delta$. This implies $\sigma(r_{1})\leq s^{\alpha}\leq\sigma(r_{2})$, and thus
$\Gamma(s^{\alpha})$ $\leq\Gamma(\sigma(r_{2}))\leq(\mathrm{id}+\rho)^{-1}(\sigma(r_{2}))=\sigma(r_{2})-\eta(\sigma(r_{2}))$
$\leq\sigma(r_{2})-\eta(\sigma(a))\leq\sigma(r_{2})-\eta(\varphi_{\min}(a))\mathds{1}=\sigma(r_{2})-\varepsilon\mathds{1}\leq\sigma(r_{1})\leq s^{\alpha}.$
[…]
$\tilde{\sigma}(r):=\left(1-\frac{r-r_{n}}{r_{n+1}-r_{n}}\right)\sigma(r_{n})+\frac{r-r_{n}}{r_{n+1}-r_{n}}\sigma(r_{n+1})\mbox{\quad for all\ }r>0,$
extended by $\tilde{\sigma}(0):=0$. Observe that upper and lower $\mathcal{K}_{\infty}$-bounds for the components of $\tilde{\sigma}$ are obtained via the analogous piecewise linear approximation with $\sigma(r_{n})$ replaced by $\varphi_{\min}(r_{n})$ or $\varphi_{\max}(r_{n})$, respectively. In this way, we obtain a path of decay for $\Gamma$. If we write $\mathrm{id}+\rho=(\mathrm{id}+\rho_{2})\circ(\mathrm{id}+\rho_{1})$ with $\rho_{1},\rho_{2}\in\mathcal{K}_{\infty}$, using Lemma A.6, we can replace $\Gamma$ with $\Gamma_{\rho_{1}}$ in our proof, and thus obtain a path of decay for $\Gamma_{\rho_{1}}$, which is a path of strict decay for $\Gamma$.
Step 2: Now, let $\sigma$ be a continuous and increasing path satisfying (i) and (ii) in Definition 2.2. First, observe that the path $\tilde{\sigma}:=(\mathrm{id}+\rho)^{-1}\circ\sigma$ satisfies
[…]
Moreover, $\tilde{\sigma}$ clearly satisfies all other assumptions of this proposition. Hence, we will assume that $\sigma$ satisfies $\Gamma^{\rho}\circ\sigma\leq\sigma$ instead of $\Gamma_{\rho}\circ\sigma\leq\sigma$. By Step 1, it now suffices to show that there exists a continuous and strictly increasing path $\tilde{\sigma}$ satisfying (i) and (ii) in Definition 2.2. This is done in four substeps.
Step 2.a: For each $r\geq 0$, we factorize $\mathrm{id}+\rho=(\mathrm{id}+\rho_{1}^{r})\circ(\mathrm{id}+\rho_{2}^{r})$ with $\rho_{1}^{r},\rho_{2}^{r}\in\mathcal{K}_{\infty}$, such that the following holds:
  1. (a)
    If $r_{1}<r_{2}$, then $\rho_{2}^{r_{1}}(r)<\rho_{2}^{r_{2}}(r)$ for every $r>0$.
  2. (b)
    The function $\rho_{2}:(r_{1},r_{2})\mapsto\rho_{2}^{r_{1}}(r_{2})$ is continuous.
  3. (c)
    There exists $\tilde{\rho}\in\mathcal{K}_{\infty}$ such that $\tilde{\rho}\leq\rho_{2}^{r}$ for all $r\geq 0$.
To achieve this, first choose a continuous and strictly increasing function $f:\mathbb{R}_{+}\rightarrow(0,1)$. Then define
$\rho_{2}^{r}:=f(r)\rho\mbox{\quad for all\ }r\in\mathbb{R}_{+}.$
[…]
$\rho_{1}^{r}:=(1-f(r))\rho\circ(\mathrm{id}+\rho_{2}^{r})^{-1}.$
Then clearly $\rho_{1}^{r}\in\mathcal{K}_{\infty}$ and
$(\mathrm{id}+\rho_{1}^{r})\circ(\mathrm{id}+\rho_{2}^{r})=\mathrm{id}+f(r)\rho+(1-f(r))\rho=\mathrm{id}+\rho.$
Item (a) above holds, because $f$ is strictly increasing. The continuity of $\rho_{2}$ as a function of both arguments easily follows from the continuity of $\rho$ and $f$, i.e. (b) holds. If we choose $f$ such that $f(r)\geq\alpha>0$ for some $\alpha>0$ (which is easily possible), it follows that $\rho_{2}^{r}\geq\alpha\rho=:\tilde{\rho}$ for all $r$, i.e. (c) holds.
Step 2.b: We define the path $\tilde{\sigma}(r):=(\mathrm{id}+\rho_{2}^{r})(\sigma(r))$ for all $r\geq 0$. Then
$\Gamma_{\tilde{\rho}}(\tilde{\sigma}(r))$ $=(\mathrm{id}+\tilde{\rho})\circ\Gamma\circ(\mathrm{id}+\rho_{2}^{r})\circ\sigma(r)$
$\leq(\mathrm{id}+\rho_{2}^{r})\circ\Gamma\circ(\mathrm{id}+\rho)\circ\sigma(r)=(\mathrm{id}+\rho_{2}^{r})\circ\Gamma^{\rho}\circ\sigma(r)$
$\leq(\mathrm{id}+\rho_{2}^{r})\circ\sigma(r)=\tilde{\sigma}(r).$
Hence, $\tilde{\sigma}$ satisfies property (i) of a path of strict decay. Moreover, it holds that
$\varphi_{\min}(r)\mathds{1}\leq\sigma(r)\leq\tilde{\sigma}(r)\leq(\mathrm{id}+\rho)\circ\sigma(r)\leq(\mathrm{id}+\rho)\circ\varphi_{\max}(r)\mathds{1},$
showing that $\tilde{\sigma}$ also satisfies property (ii) of a path of strict decay.
Step 2.c: Let $r_{1}<r_{2}$ and assume without loss of generality that $r_{1}>0$. We then have $\sigma(r_{1})\geq\varphi_{\min}(r_{1})\mathds{1}\gg 0$. In particular, all components of the vector $\sigma(r_{1})$ are contained in the compact interval $J:=[\varphi_{\min}(r_{1}),\varphi_{\max}(r_{1})]\subset(0,\infty)$. Observe that there exists $\alpha>0$ with $\rho_{2}^{r_{2}}(r)-\rho_{2}^{r_{1}}(r)=(f(r_{2})-f(r_{1}))\rho(r)\geq\alpha$ for all $r\in J$, implying that $\tilde{\sigma}$ is strictly increasing:
$\tilde{\sigma}(r_{1})=(\mathrm{id}+\rho_{2}^{r_{1}})(\sigma(r_{1}))\ll(\mathrm{id}+\rho_{2}^{r_{2}})(\sigma(r_{1}))\leq(\mathrm{id}+\rho_{2}^{r_{2}})(\sigma(r_{2}))=\tilde{\sigma}(r_{2}).$
Step 2.d: It remains to prove the continuity of $\tilde{\sigma}$. To this end, fix $r_{0}\geq 0$ and $\varepsilon>0$. We first observe that for any $r\geq 0$ it holds that
$\|\tilde{\sigma}(r)-\tilde{\sigma}(r_{0})\|\leq\|\sigma(r)-\sigma(r_{0})\|+\sup_{i\in\mathcal{I}}|\rho_{2}(r,\sigma_{i}(r))-\rho_{2}(r_{0},\sigma_{i}(r_{0}))|.$
By continuity of $\sigma$, we can find $\delta>0$ such that $|r-r_{0}|\leq\delta$ implies $\|\sigma(r)-\sigma(r_{0})\|\leq\varepsilon/2$. By uniform continuity of $\rho_{2}$ on compact subsets of $\mathbb{R}_{+}\times\mathbb{R}_{+}$, we see that the second term can also be made smaller than $\varepsilon/2$ if $\delta$ is chosen small enough. Hence, $\tilde{\sigma}$ is continuous. ∎
Motivated by the preceding proposition, we introduce the following definition.
Definition 2.3.
Given a gain operator $\Gamma$, a path $\sigma$ in the cone $\ell^{\infty}_{+}(\mathcal{I})$ is called a $C^{0}$-path of strict decay for $\Gamma$ if it satisfies the following properties:
  1. (i)
    Strict decay: There is $\rho\in\mathcal{K}_{\infty}$ such that $\sigma(r)\in\Psi(\Gamma_{\rho})$ for all $r\geq 0$.
  2. (ii)
    Coercivity: There are $\varphi_{\min},\varphi_{\max}\in\mathcal{K}_{\infty}$ such that $\varphi_{\min}(r)\mathds{1}\leq\sigma(r)\leq\varphi_{\max}(r)\mathds{1}$ for all $r\geq 0$.
  3. (iii)
    Monotonicity: $\sigma$ is increasing.
  4. (iv)
    Continuity: $\sigma$ is continuous.
If all of the above properties hold with the exception that $\rho=0$ in (i), we call $\sigma$ a $C^{0}$-path of decay for $\Gamma$.
We conclude this section with a discussion of the axioms imposed on the graph $\mathcal{G}$ and the families $\{\gamma_{ij}\}$ and $\{\mu_{i}\}$.
Remark 2.3.
Let us summarize our main assumptions on the data defining a gain operator:
  • Finiteness of the sets $\mathcal{I}_{i}$
  • Equicontinuity of $\{\gamma_{ij}\}$
  • (M1) Uniform positive definiteness for $\{\mu_{i}\}$
  • (M2) Monotonicity of each $\mu_{i}$
  • (M3) Finite continuity of each $\mu_{i}$
  • (M4) Equicontinuity of $\{\mu_{i}\}$
Axioms (M2) and (M3) are natural and necessary to obtain a monotone continuous operator and they are also used in the finite-dimensional theory. So there is no need to further discuss these axioms.
Note, however, that we have used (M3) to prove the sequential continuity of $\Gamma$ in the weak-topology, not its norm-continuity. Continuity in the norm-topology is needed, in particular, in the path construction methods to be discussed in Subsection 3.3. More concretely, the proofs of Proposition 3.3 and Theorem 3.3 explicitly make use of the continuity of $\Gamma$. Additionally, the uniform continuity of $\Gamma$ on bounded sets is used in Lemma 2.1, which in turn is needed to relate the uniform NJI condition to UGAS of $\Sigma(\Gamma)$, which is a crucial part of the theory. Now, as an example, consider a max-type gain operator
$\Gamma_{i}(s)=\sup_{j\in\mathcal{I}_{i}}\gamma_{ij}(s_{j}),\quad i\in\mathcal{I}$
with $|\mathcal{I}_{i}|=1$ for all $i\in\mathcal{I}$, say $\mathcal{I}_{i}=\{j_{i}\}$. Let us further assume that this operator is well-defined and norm-continuous. Then, in particular, for every $r_{0}\geq 0$ and every $\varepsilon>0$ there exists $\delta>0$ such that $|r-r_{0}|\leq\delta$ implies
$\|\Gamma(r\mathds{1})-\Gamma(r_{0}\mathds{1})\|=\sup_{i\in\mathcal{I}}|\gamma_{ij_{i}}(r)-\gamma_{ij_{i}}(r_{0})|\leq\varepsilon.$
But this is equivalent to the equicontinuity of $\{\gamma_{ij}\}$ at $r_{0}$. Hence, we see that already in this very simple case the pointwise equicontinuity of $\{\gamma_{ij}\}$ is needed. Once this equicontinuity condition for $\{\gamma_{ij}\}$ is imposed, in the general case, it is clear that the equicontinuity condition for $\{\mu_{i}\}$, i.e. Axiom (M4), is needed to obtain the uniform continuity of $\Gamma$ on bounded sets. The two equicontinuity conditions could be replaced by a joint condition on the functions $\gamma_{ij}$ and $\mu_{i}$; but then we could just simply assume the continuity of $\Gamma$, and would have to find a method to verify it in each specific case. Altogether, it is fair to say that the two equicontinuity assumptions are justified.
The uniform positive definiteness of $\{\mu_{i}\}$ was not used so far, only the requirement that $\mu_{i}(0)=0$ for each $i$. However, we will use this axiom for several central arguments in the theory to be developed. We will need it, whenever we have to bound a trajectory of $\Gamma$ from below, employing Assumption 2.1 in addition. This is the case in Subsection 2.4 about stability properties, when we show that the uniform NJI condition in combination with the cofinality of the decay set $\Psi(\Gamma_{\rho})$ implies strong stability properties, which are necessary for the existence of a path of strict decay (see Corollary 2.1). It is also the case in the path construction method by linear interpolation (see Theorem 3.3). Still, it is unclear whether uniform positive definiteness is needed for a complete characterization of the small-gain condition. We have to leave this question open. However, for the theory developed in this paper, it is certainly required and it is always satisfied in the finite case, where we explicitly use it to characterize the existence of a path of strict decay, see Theorem 4.5.
The assumption that each neighbor set $\mathcal{I}_{i}$ is finite was used to prove the sequential weak continuity of $\Gamma$, which is absolutely crucial for the theory developed in this paper, because it guarantees the existence of fixed points and points of decay in various proofs. Without the finiteness assumption, it is unclear how to even show that $\Gamma$ is well-defined. For instance, consider a sum-type operator. Then an additional condition on the gains $\gamma_{ij}$ needs to guarantee that the vectors $(\gamma_{ij}(s_{j}))_{j\in\mathcal{I}_{i}}$ with $s\in\ell^{\infty}_{+}(\mathcal{I})$ are summable, i.e. contained in $\ell^{1}(\mathcal{I})$. Probably, one could find such a condition and additional ones that guarantee the weak continuity. But currently, due to a lack of possible applications, it is unclear if this is worth the effort.
Possible improvements of the current set of axioms could be obtained by
  • combining Axioms (M3) and (M4) into one axiom
  • combining the uniformity part of Axiom (M1) and Assumption 2.1, since they are always used together
2.3 Associated operators
Given a gain operator $\Gamma$, we introduce further monotone operators associated with $\Gamma$, which are useful for analyzing under which conditions a $C^{0}$-path of strict decay exists. First, we define the augmented gain operator
[…]
Moreover, for each $b\in\ell^{\infty}_{+}(\mathcal{I})$, we define the projected gain operator [^6] Observe that the mapping $s\mapsto b\oplus s$ can be regarded as a projection to the set $\{s\geq b\}$.
$\bar{\Gamma}_{b}(s):=b\oplus\Gamma(s),\quad\bar{\Gamma}_{b}:\ell^{\infty}_{+}(\mathcal{I})\rightarrow\ell^{\infty}_{+}(\mathcal{I}).$
The properties of these operators listed in the following proposition can easily be derived, see [14, Prop. 3.1 and Lem. 3.2].
Proposition 2.4.
The following statements hold for all $b,s\in\ell^{\infty}_{+}(\mathcal{I})$:
  1. (a)
    $\bar{\Gamma}_{b}$ and $\hat{\Gamma}$ are well-defined continuous monotone operators.
  2. (b)
    $\bar{\Gamma}_{b}(0)=b$ and $\hat{\Gamma}(0)=0$.
  3. (c)
    $\bar{\Gamma}_{b}(s)\geq b$ and $\hat{\Gamma}(s)\geq s$.
  4. (d)
    $\Psi(\bar{\Gamma}_{b})\subset\Psi(\Gamma)$ and $\mathrm{Fix}(\hat{\Gamma})=\Psi(\Gamma)$.
  5. (e)
    $\hat{\Gamma}^{n}(s)=\bar{\Gamma}_{s}^{n}(s)$ for all $n\in\mathbb{N}_{0}$.
  6. (f)
    $\bar{\Gamma}_{b}$ and $\hat{\Gamma}$ are sequentially continuous in the weak-topology.
[…]
Proposition 2.5.
The following statements are equivalent:
  1. (a)
[…]
  2. (b)
    There exists a path $\sigma$ with properties (i)–(iii) of a $C^{0}$-path of decay.
  3. (c)
    There exists a path $\sigma$ with properties (i) and (ii) of a $C^{0}$-path of decay.
  4. (d)
    The set $\Psi(\Gamma)$ is uniformly cofinal.
Proof.
(a) $\Rightarrow$ (b): Let $\varphi\in\mathcal{K}_{\infty}$ be chosen such that $\|\hat{\Gamma}^{n}(s)\|\leq\varphi(\|s\|)$. We define $\sigma$ componentwise by
$\sigma_{i}(r):=\lim_{n\rightarrow\infty}\hat{\Gamma}^{n}_{i}(r\mathds{1})\mbox{\quad for all\ }i\in\mathcal{I},\ r\geq 0.$ (4)
[…]
(b) $\Rightarrow$ (c): This is trivial.
(c) $\Rightarrow$ (d): Assume that $\varphi_{1}(r)\mathds{1}\leq\sigma(r)\leq\varphi_{2}(r)\mathds{1}$ for some $\varphi_{1},\varphi_{2}\in\mathcal{K}_{\infty}$. For each $s\in\ell^{\infty}_{+}(\mathcal{I})$, we put $s^{*}:=\sigma(\varphi_{1}^{-1}(\|s\|))$. It follows that
$s\leq\|s\|\mathds{1}=\varphi_{1}(\varphi_{1}^{-1}(\|s\|))\mathds{1}\leq\sigma(\varphi_{1}^{-1}(\|s\|))=s^{*}.$
Moreover, $\Gamma(s^{*})\leq s^{*}$ by construction and $\|s^{*}\|\leq\varphi_{2}\circ\varphi_{1}^{-1}(\|s\|)$.
(d) $\Rightarrow$ (a): Let $\varphi\in\mathcal{K}_{\infty}$ be such that for every $s\in\ell^{\infty}_{+}(\mathcal{I})$ there is $s^{*}\in\Psi(\Gamma)$ with $s\leq s^{*}$ and $\|s^{*}\|\leq\varphi(\|s\|)$. Then $\hat{\Gamma}^{n}(s)\leq\hat{\Gamma}^{n}(s^{*})=s^{*}$ for all $s$ and $n$, implying $\|\hat{\Gamma}^{n}(s)\|\leq\varphi(\|s\|)$ as required. ∎
2.4 Stability properties
We say that a gain operator $\Gamma$ satisfies the no-joint-increase (NJI) condition if
$\Gamma(s)\not\geq s\mbox{\quad for all\ }s>0.$
This is the classical small-gain condition, which in the finite case characterizes the existence of a path of strict decay (when required for $\Gamma_{\rho}$), see Theorem 4.4 and Theorem 4.5. In the general case, the NJI condition is an important ingredient of the following characterization of GATT.
Proposition 2.6.
The following statements are equivalent:
  1. (a)
[…]
  2. (b)
    The trajectories of $\Sigma(\Gamma)$ are norm-bounded and $\Gamma$ satisfies the NJI condition.
Proof.
(a) $\Rightarrow$ (b): A trajectory which weakly converges is norm-bounded by Lemma A.1. Hence, every trajectory of $\Sigma(\Gamma)$ is norm-bounded. The NJI condition holds, because otherwise there would exist $s>0$ with $\Gamma(s)\geq s$ implying $\Gamma^{n}(s)\geq s$ for all $n$, and hence $\Gamma^{n}(s)$ could not converge to zero componentwise.
(b) $\Rightarrow$ (a): For $s\in\ell^{\infty}_{+}(\mathcal{I})$, let $\omega_{*}(s)$ denote the weak $\omega$-limit set of $s$, i.e. the set of points in $\ell^{\infty}_{+}(\mathcal{I})$ which are weak-limits of subsequences of $(\Gamma^{n}(s))_{n\in\mathbb{N}_{0}}$. By assumption, there exists $b\in\ell^{\infty}_{+}(\mathcal{I})$ such that $\Gamma^{n}(s)$ is contained in the closed order interval $[0,b]$ for all $n$. Since closed order intervals are sequentially compact in the weak-topology by Lemma A.3, $\omega_{*}(s)$ is nonempty. Now, fix $s\in\ell^{\infty}_{+}(\mathcal{I})$. Then there exists $u\in\ell^{\infty}_{+}(\mathcal{I})$ with $\omega_{*}(s)\leq u$. From monotonicity and $\Gamma(\omega_{*}(s))=\omega_{*}(s)$, [^7] This is a general property of $\omega$-limit sets, which also holds in this context, since $\Gamma$ is sequentially continuous in the weak-topology. it follows that $\omega_{*}(s)\leq\Gamma^{n}(u)$ for all $n$, implying $\omega_{*}(s)\leq\omega_{*}(u)$. By the same arguments, there is $v\in\ell^{\infty}_{+}(\mathcal{I})$ with $\omega_{*}(u)\leq\omega_{*}(v)$. The set
$F:=\left\{\tilde{s}\in\ell^{\infty}_{+}(\mathcal{I})\ :\ \omega_{*}(s)\leq\tilde{s}\leq\omega_{*}(v)\right\}$
is the intersection of closed order intervals. Hence, it is weakly sequentially compact and convex. Moreover, it contains $\omega_{*}(u)$, and hence is nonempty. Since $\Gamma$ is weakly sequentially continuous, the set $\Gamma(F)$ is weakly sequentially compact. Monotonicity of $\Gamma$ and invariance of $\omega$-limit sets implies $\Gamma(F)\subset F$. It follows from Schauder’s fixed point theorem that there is a fixed point in $F$. From the NJI condition, it follows that the origin is the only fixed point. Hence, $0\in F$, implying $\omega_{*}(s)=\{0\}$, which means that $\Gamma^{n}(s)\stackrel{{\scriptstyle\star}}{{\rightharpoonup}}0$. ∎
Remark 2.4.
The proof of the preceding proposition is essentially taken from the much more general result [6, Lem. 1] about the existence of fixed points for monotone operators with bounded orbits.
We now aim at characterizations of GATT and UGAS. Here, the following lemma will be useful.
Lemma 2.1.
Given a gain operator $\Gamma$, for all $n\in\mathbb{N}$ and $r,\varepsilon>0$ there exists $\delta>0$ such that for every $s\geq 0$ with $\|s\|\leq r$ and every $i\in\mathcal{I}$ the following holds: $\Gamma_{j}(s)\geq s_{j}-\delta$ for all $j\in\mathcal{N}^{-}_{i}(n-1)$ implies $\Gamma^{n}_{i}(s)\geq s_{i}-\varepsilon$.
Proof.
We put $\delta_{n}:=\varepsilon$ and choose $\varepsilon_{i},\delta_{i}>0$, $1\leq i\leq n-1$, according to the following algorithm: We start with $\varepsilon_{n-1}:=\delta_{n}/2$ and continue recursively by
$\|s^{1}-s^{2}\|\leq\delta_{l}\quad\Rightarrow\quad\|\Gamma(s^{1})-\Gamma(s^{2})\|\leq\varepsilon_{l},\quad\varepsilon_{l-1}:=\frac{\delta_{l}}{2}$
whenever $\max\{\|s^{1}\|,\|s^{2}\|\}\leq r$, which is possible by the uniform continuity of $\Gamma$ on bounded sets. We do this for l=n1l=n-1 down to $l=2$. Eventually, we choose $\delta_{1}=:\delta$ satisfying the two requirements $\delta_{1}\leq\min\left\{\varepsilon_{i}:1\leq i\leq n-1\right\}$ and
$\|s^{1}-s^{2}\|\leq\delta_{1}\quad\Rightarrow\quad\|\Gamma(s^{1})-\Gamma(s^{2})\|\leq\varepsilon_{1}.$
Now, let $s\geq 0$ with $\|s\|\leq r$ and $i\in\mathcal{I}$ such that $\Gamma_{j}(s)\geq s_{j}-\delta$ for all $j\in\mathcal{N}^{-}_{i}(n-1)$. Inductively, we get estimates for $\Gamma^{l}_{i}(s)$, $1\leq l\leq n$. First,
$\Gamma_{j}(s)\geq s_{j}-\delta=s_{j}-\delta_{1}\mbox{\quad for all\ }j\in\mathcal{N}^{-}_{i}(n-1)$
by assumption. Using, as the induction hypothesis, that $\Gamma^{l}_{j}(s)\geq s_{j}-\delta_{l}$ for all $j\in\mathcal{N}^{-}_{i}(n-l)$ with $l\in\{1,\ldots,n-1\}$, for each $j\in\mathcal{N}^{-}_{i}(n-l-1)$ we obtain
$\Gamma^{l+1}_{j}(s)$ $=\Gamma_{j}(\Gamma^{l}(s))=\Gamma_{j}([\Gamma^{l}_{k}(s)]_{k\in\mathcal{I}_{j}})\geq\Gamma_{j}([\max\{0,s_{k}-\delta_{l}\}]_{k\in\mathcal{I}_{j}})$
$\geq\Gamma_{j}(s)-\varepsilon_{l}\geq s_{j}-(\delta_{1}+\varepsilon_{l})\geq s_{j}-2\varepsilon_{l}=s_{j}-\delta_{l+1}.$
For l=n1l=n-1, this yields the desired estimate $\Gamma^{n}_{i}(s)\geq s_{i}-\delta_{n}=s_{i}-\varepsilon$. ∎
We now introduce a uniform version of the NJI condition.
Definition 2.4.
We say that a gain operator $\Gamma$ satisfies the uniform no-joint-increase (NJI) condition if the following holds: For all $r,\varepsilon>0$ there are $n\in\mathbb{N}$ and $\delta>0$ such that for all $s\in\ell^{\infty}_{+}(\mathcal{I})$ and $i\in\mathcal{I}$ the following implication holds:
$s_{i}\geq\varepsilon\ \wedge\ \|s\|\leq r\quad\Rightarrow\quad\exists j\in\mathcal{N}^{-}_{i}(n):\ s_{j}\geq\delta\ \wedge\ \Gamma_{j}(s)<s_{j}.$
The next proposition shows that in the finite case the uniform NJI condition reduces to the NJI condition.
Proposition 2.7.
If $|\mathcal{I}|<\infty$, then the following statements are equivalent:
  1. (a)
    $\Gamma$ satisfies the NJI condition.
  2. (b)
    $\Gamma$ satisfies the uniform NJI condition.
Proof.
[…]
$\tilde{s}_{j}:=\left\{\begin{array}[]{cc}s_{j}&\mbox{if }j\in\mathcal{N}^{-}_{i},\\ 0&\mbox{otherwise}.\end{array}\right$.
[…]
The following two propositions relate the uniform NJI condition to GATT.
Proposition 2.8.
If $\Sigma(\Gamma)$ is GATT, then $\Gamma$ satisfies the uniform NJI condition.
Proof.
From the proof of Lemma A.8, we see that global attractivity together with monotonicity implies uniform global attractivity, i.e.
$\forall r>0\ \forall\varepsilon>0\ \exists n\in\mathbb{N}:\ k\geq n\mbox{ and }\|s\|\leq r\quad\Rightarrow\quad\|\Gamma^{k}(s)\|\leq\varepsilon.$
Now, assume to the contrary that there exist $\varepsilon,r>0$ such that for every $n\in\mathbb{N}$ and $\delta>0$ there are $s\in\ell^{\infty}_{+}(\mathcal{I})$ and $i\in\mathcal{I}$ with $s_{i}\geq\varepsilon$ and $\|s\|\leq r$, but $\Gamma_{j}(s)\geq s_{j}$ or $s_{j}<\delta$ for all $j\in\mathcal{N}^{-}_{i}(n)$. We then choose $n$ large enough such that $\|s\|\leq r$ implies $\|\Gamma^{n}(s)\|<\varepsilon/2$, and subsequently, we choose $\delta=\delta(n,r,\varepsilon/2)>0$ according to Lemma 2.1. By our assumption, there exist $s\geq 0$ and $i\in\mathcal{I}$ with
$\varepsilon\leq s_{i}\leq\Gamma^{n}_{i}(s)+\frac{\varepsilon}{2}\leq\|\Gamma^{n}(s)\|+\frac{\varepsilon}{2}<\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon,$
which is a contradiction. To apply the lemma, we have used that both $\Gamma_{j}(s)\geq s_{j}$ and $s_{j}<\delta$ imply $\Gamma_{j}(s)\geq s_{j}-\delta$. ∎
Proposition 2.9.
Let the following assumptions hold for some $\rho\in\mathcal{K}_{\infty}$:
  1. (i)
    The set $\Psi(\Gamma_{\rho})$ is cofinal.
  2. (ii)
    $\Gamma_{\rho}$ satisfies the uniform NJI condition.
  3. (iii)
    There exists $M>0$ with $|\mathcal{I}_{i}|\leq M$ for all $i\in\mathcal{I}$.
Then there exists $\rho^{\prime}\in\mathcal{K}_{\infty}$ such that $\Sigma(\Gamma_{\rho^{\prime}})$ is GATT.
Proof.
For notational simplicity, we write the proof for $\rho^{\prime}=0$, i.e. we show that $\Sigma(\Gamma)$ is GATT. We can do this without loss of generality, since we can decompose $\mathrm{id}+\rho=(\mathrm{id}+\rho_{1})\circ(\mathrm{id}+\rho_{2})$ according to Lemma A.6 and use $\rho^{\prime}=\rho_{1}$, which replaces $\Gamma$ with $\Gamma_{\rho_{1}}$ throughout the proof. Observe that $\Psi(\Gamma_{\rho})\subset\Psi(\Gamma)$, hence assumption (i) implies that $\Psi(\Gamma)$ is cofinal. It then suffices to prove that every trajectory starting in $\Psi(\Gamma)$ converges to zero. To this end, fix $s\in\Psi(\Gamma)$ and $\varepsilon>0$ with $\varepsilon\mathds{1}\leq s$, and consider the decreasing trajectory $s^{n}:=\Gamma^{n}(s)$, $n\in\mathbb{N}_{0}$. Note that $s^{n}\stackrel{{\scriptstyle\star}}{{\rightharpoonup}}0$, because assumption (ii) implies that the origin is the only fixed point of $\Gamma$.
By assumption (ii), there are $m=m(\varepsilon,\|s\|)\in\mathbb{N}$ and $\delta=\delta(\varepsilon,\|s\|)>0$ such that for every $i\in\mathcal{I}$ we have $\Gamma_{j}(s^{n})<(\mathrm{id}+\rho)^{-1}(s^{n}_{j})$ for some $j=j(n)\in\mathcal{N}^{-}_{i}(m)$ with $s_{j}^{n}\geq\delta$ as long as $s^{n}_{i}\geq\varepsilon$. We define
$K_{i}:=\max\{k\in\mathbb{N}:s^{k}_{i}\geq\varepsilon\}\mbox{\quad for all\ }i\in\mathcal{I}.$
Then, for each $i\in\mathcal{I}$ there must exist some $j_{i}\in\mathcal{N}^{-}_{i}(m)$ with
$|\{k\in[0;K_{i}-1]:s^{k}_{j_{i}}\geq\delta\wedge\Gamma_{j_{i}}(s^{k})<(\mathrm{id}+\rho)^{-1}(s^{k}_{j_{i}})\}|\geq\frac{K_{i}}{|\mathcal{N}^{-}_{i}(m)|}.$
If this was not the case, we would end up with the contradiction
$K_{i}$ $=|\{k\in[0;K_{i}-1]:\exists j\in\mathcal{N}^{-}_{i}(m),\ s^{k}_{j}\geq\delta\wedge\Gamma_{j}(s^{k})<(\mathrm{id}+\rho)^{-1}(s^{k}_{j})\}|$
$\leq\sum_{j\in\mathcal{N}^{-}_{i}(m)}|\{k\in[0;K_{i}-1]:s^{k}_{j}\geq\delta\wedge\Gamma_{j}(s^{k})<(\mathrm{id}+\rho)^{-1}(s^{k}_{j})\}|<K_{i}.$
From assumption (iii), it follows that $|\mathcal{N}^{-}_{i}(m)|\leq B_{m}$ with $B_{m}:=\sum_{l=0}^{m}M^{l}$. Now, let $k_{i}\in[0;K_{i}-1]$ be the largest time at which $s^{k}_{j_{i}}\geq\delta$ and $\Gamma_{j_{i}}(s^{k})<(\mathrm{id}+\rho)^{-1}(s^{k}_{j_{i}})$. Then we obtain
$\delta\leq s^{k_{i}}_{j_{i}}\leq(\mathrm{id}+\rho)^{-(\lceil K_{i}/|\mathcal{N}^{-}_{i}(m)|\rceil-1)}(s_{j_{i}})\leq(\mathrm{id}+\rho)^{-(\lceil K_{i}/B_{m}\rceil-1)}(\|s\|).$
This implies that there is a bound on $K_{i}$, which is uniform in $i$, since the right-hand side converges to zero as $K_{i}\rightarrow\infty$. But if $K_{i}\leq K$ for all $i\in\mathcal{I}$, then $\|\Gamma^{K+1}(s)\|<\varepsilon$. The number $\varepsilon>0$ can be chosen arbitrarily small. Hence, we have shown that $\Gamma^{n}(s)$ converges to zero, which completes the proof. ∎
Adding Assumption 2.1, we actually obtain a much stronger conclusion.
Corollary 2.1.
In addition to the assumptions of Proposition 2.9, let Assumption 2.1 hold. Then $\Sigma(\hat{\Gamma}_{\rho^{\prime}})$ is UGS and $\Sigma(\Gamma_{\rho^{\prime}})$ is UGAS.
Proof.
By Lemma A.10, it suffices to show that there exists a sequence $(s^{n})_{n\in\mathbb{N}}$ in $\mathrm{int}(\ell^{\infty}_{+}(\mathcal{I}))\cap\Psi(\Gamma_{\rho^{\prime}})$, converging to the origin. To obtain such a sequence, take any $s\in\mathrm{int}(\ell^{\infty}_{+}(\mathcal{I}))\cap\Psi(\Gamma_{\rho^{\prime}})$ (which exists, because $\Psi(\Gamma_{\rho^{\prime}})\supset\Psi(\Gamma_{\rho})$ is cofinal), and put $s^{n}:=\Gamma_{\rho^{\prime}}^{n}(s)$. Since $\Sigma(\Gamma_{\rho^{\prime}})$ is GATT, we have $s^{n}\rightarrow 0$. Moreover, from Assumption 2.1, it follows that $s^{n}\in\mathrm{int}(\ell^{\infty}_{+}(\mathcal{I}))$ for all $n$ (using (2) repeatedly). Finally, since $\Psi(\Gamma_{\rho^{\prime}})$ is forward-invariant under $\Gamma_{\rho^{\prime}}$, we also have $s^{n}\in\Psi(\Gamma_{\rho^{\prime}})$ for all $n$. It then follows from Proposition 2.5 that $\Sigma(\hat{\Gamma}_{\rho^{\prime}})$ is UGS, which implies that $\Sigma(\Gamma_{\rho^{\prime}})$ is UGS. Hence, we have UGS and GATT for $\Sigma(\Gamma_{\rho^{\prime}})$, which is equivalent to UGAS (using Lemma A.8). ∎
From the above proof, we obtain another useful corollary.
Corollary 2.2.
Let $\Gamma$ be a gain operator satisfying Assumption 2.1. Then the following statements are equivalent:
  1. (a)
    $\Sigma(\Gamma)$ is GATT and $\Psi(\Gamma)$ is cofinal.
  2. (b)
    $\Sigma(\Gamma)$ is UGAS and $\Psi(\Gamma)$ is uniformly cofinal.
In accordance with Proposition 2.6, we would like to replace assumption (i) in Proposition 2.9 with the norm-boundedness of the trajectories of $\Sigma(\Gamma_{\rho})$. However, it is unclear if this is possible. But we can replace it with the norm-boundedness of the trajectories of $\Sigma(\hat{\Gamma}_{\rho})$, as the next proposition shows.
Proposition 2.10.
The following statements are equivalent:
  1. (a)
    The set $\Psi(\Gamma)$ is cofinal.
  2. (b)
    For every $b\in\ell^{\infty}_{+}(\mathcal{I})$, the trajectories of $\Sigma(\bar{\Gamma}_{b})$ are norm-bounded.
  3. (c)
[…]
  4. (d)
    For every $s\geq 0$, there exist $\hat{s}\geq 0$ and $n\in\mathbb{N}$ with
    $s\leq\Gamma^{l}(\hat{s})\mbox{\qquad for\ }0\leq l<n\mbox{\qquad and \qquad}\Gamma^{n}(\hat{s})\leq\hat{s}$.
Proof.
(a) $\Rightarrow$ (b): Pick $b,s\geq 0$. Let $\hat{s}\geq b\oplus s$ such that $\Gamma(\hat{s})\leq\hat{s}$. Then $\bar{\Gamma}_{b}(\hat{s})=b\oplus\Gamma(\hat{s})\leq b\oplus\hat{s}=\hat{s}$. This implies $\bar{\Gamma}_{b}^{n}(s)\leq\bar{\Gamma}^{n}_{b}(\hat{s})\leq\hat{s}$ for all $n$.
(b) $\Rightarrow$ (c): This follows from the identity $\hat{\Gamma}^{n}(s)\equiv\bar{\Gamma}^{n}_{s}(s)$.
(c) $\Rightarrow$ (d): If, for a given $s\geq 0$, the increasing trajectory $(\hat{\Gamma}^{n}(s))_{n\in\mathbb{N}_{0}}$ is norm-bounded, it weakly converges to a fixed point $s^{*}$ of $\hat{\Gamma}$. Then $\Gamma(s^{*})\leq s^{*}$ and $s^{*}\geq s$. Hence, (d) holds with $n=1$.
[…]
From the characterization given in (d), we can conclude that the cofinality of $\Psi(\Gamma)$ follows from a strong version of UGAS, including uniform global exponential stability as a special case.
Proposition 2.11.
Assume that $\Sigma(\Gamma)$ is UGAS with a $\mathcal{K}\mathcal{L}$-function of the form $\beta(r,n)=r\gamma(n)$, $\gamma\in\mathcal{L}$, and let Assumption 2.1 hold. Then $\Psi(\Gamma)$ is cofinal.
Proof.
We verify (d) in Proposition 2.10. To this end, fix $\xi,\eta\in\mathcal{K}_{\infty}$ from (M1) and Assumption 2.1. We can assume that $\xi\circ\eta<\mathrm{id}$. We then fix $n\in\mathbb{N}$ with $\gamma(n)\leq 1$, implying $\beta(r,n)\leq r$ for all $r$. For a given $s\geq 0$, let $\hat{s}:=(\xi\circ\eta)^{-n}(\|s\|)\mathds{1}$. Then $\Gamma^{l}_{i}(\hat{s})\geq(\xi\circ\eta)^{l}\circ(\xi\circ\eta)^{-n}(\|s\|)=(\xi\circ\eta)^{l-n}(\|s\|)\geq\|s\|$ for $0\leq l<n$ and $i\in\mathcal{I}$, which implies $\Gamma^{l}(\hat{s})\geq s$ for $0\leq l<n$. Moreover, $\Gamma^{n}(\hat{s})\leq\beta(\|\hat{s}\|,n)\mathds{1}\leq(\xi\circ\eta)^{-n}(\|s\|)\mathds{1}=\hat{s}$. ∎
It makes sense to combine UGAS with the property expressed in statement (d) of Proposition 2.10 for two reasons. First, in the finite case UGAS of $\Gamma_{\rho}$ implies (d) for a possibly smaller gain operator $\Gamma_{\rho^{\prime}}$, and second, both UGAS of $\Sigma(\Gamma_{\rho})$ and the cofinality of $\Psi(\Gamma_{\rho})$ are implied by the existence of a $C^{0}$-path of strict decay, as we will prove in the next section.
We then have the following equivalence, which immediately follows from the proven results.
Corollary 2.3.
Given a gain operator $\Gamma$, satisfying Assumption 2.1 and $\sup_{i\in\mathcal{I}}|\mathcal{I}_{i}|<\infty$, the following statements are equivalent:
  1. (a)
[…]
  2. (b)
    There exists $\rho\in\mathcal{K}_{\infty}$ such that $\Gamma_{\rho}$ satisfies the uniform NJI condition and the trajectories of $\Sigma(\hat{\Gamma}_{\rho})$ are norm-bounded.
3 Existence of paths of strict decay
In this section, we derive necessary conditions and sufficient conditions for the existence of a path of strict decay for a given gain operator $\Gamma$. Recall that by Proposition 2.3 it suffices to consider $C^{0}$-paths.
3.1 The $\oplus$-MBI property
We first introduce a property that will turn out to be crucial for characterizing the existence of a path of strict decay.
Definition 3.1.
A gain operator $\Gamma$ satisfies the $\oplus$-MBI property if there exists $\varphi\in\mathcal{K}_{\infty}$ such that for all $s,b\in\ell^{\infty}_{+}(\mathcal{I})$ the following implication holds:
$s\leq\bar{\Gamma}_{b}(s)\quad\Rightarrow\quad\|s\|\leq\varphi(\|b\|).$
Remark 3.1.
The letters MBI stand for monotone bounded invertibility. As an explanation for this name, think of $\Gamma$ being a linear operator, $\oplus$ being replaced with addition, and $\varphi$ a linear function. Then the required property could be rewritten as
$(\mathrm{id}-\Gamma)s\leq b\quad\Rightarrow\quad\|s\|\leq K\|b\|,$
with a constant $K>0$. If we assume that $\mathrm{id}-\Gamma$ is an invertible bounded operator with positive inverse, this property is satisfied with $K=\|(\mathrm{id}-\Gamma)^{-1}\|$. Hence, the $\oplus$-MBI property generalizes the bounded invertibility and positivity of the operator $\mathrm{id}-\Gamma$. By [10, Thm. 3.3], this property is equivalent to the spectral radius of $\Gamma$ being strictly less than one. Also recall that the bounded inverse theorem from functional analysis guarantees that an invertible bounded operator between Banach spaces always has a bounded inverse.
Remark 3.2.
It is easy to see that for verifying the $\oplus$-MBI property one only needs to consider vectors of the form $b=r\mathds{1}$ with $r\geq 0$.
The $\oplus$-MBI property first appeared in [20], but was only used for max-type gain operators in the finite case. The next proposition provides several characterizations.
Proposition 3.1.
For a gain operator $\Gamma$, let $A:=\bigcup_{r\geq 0}\mathrm{Fix}(\bar{\Gamma}_{r\mathds{1}})\subset\Psi(\Gamma)$. Then the following statements are equivalent:
  1. (a)
    $\Gamma$ satisfies the $\oplus$-MBI property.
  2. (b)
[…]
  3. (c)
[…]
  4. (d)
    The set $\Psi(\Gamma)$ is cofinal, $\Gamma$ satisfies the NJI condition and $A$ is coercive.
Proof.
[…]
(b) $\Rightarrow$ (c): This follows from Proposition 2.6.
[…]
Corollary 3.1.
Under the additional assumption that each of the operators $\bar{\Gamma}_{b}$ has a unique fixed point, the statements (a)–(c) in Proposition 3.1 are still equivalent after removing the requirement that $A$ is coercive.
Proof.
[…]
Finally, we show that the $\oplus$-MBI property implies that each of the operators $\bar{\Gamma}_{b}$ has a minimal and a maximal fixed point.
Proposition 3.2.
Let $\Gamma$ be a gain operator satisfying the $\oplus$-MBI property. Then, for each $b\in\ell^{\infty}_{+}(\mathcal{I})$, there exist fixed points $s_{*}(b)$ and $s^{*}(b)$ of $\bar{\Gamma}_{b}$ such that the following statements hold:
  1. (a)
    Every $s\in\mathrm{Fix}(\bar{\Gamma}_{b})$ satisfies $s_{*}(b)\leq s\leq s^{*}(b)$.
  2. (b)
    If $b^{1}\leq b^{2}$, then $s_{*}(b_{1})\leq s_{*}(b_{2})$ and $s^{*}(b_{1})\leq s^{*}(b_{2})$.
  3. (c)
    For every $s\leq s_{*}(b)$, we have $\bar{\Gamma}_{b}^{n}(s)\stackrel{{\scriptstyle\star}}{{\rightharpoonup}}s_{*}(b)$.
  4. (d)
    For every $s\geq s^{*}(b)$, we have $\bar{\Gamma}_{b}^{n}(s)\stackrel{{\scriptstyle\star}}{{\rightharpoonup}}s^{*}(b)$.
Proof.
We define $s_{*}(b)$ as the weak limit of the trajectory $s^{n}:=\bar{\Gamma}_{b}^{n}(b)$, $n\in\mathbb{N}_{0}$. Since $\bar{\Gamma}_{b}(b)\geq b$ and $\bar{\Gamma}_{b}$ is monotone, this trajectory is increasing. Since $s^{n}\leq s^{n+1}=b\oplus\Gamma(s^{n})$ implies $\|s^{n}\|\leq\varphi(\|b\|)$ for all $n\in\mathbb{N}_{0}$ by the $\oplus$-MBI property, the trajectory is also bounded. Hence, the weak limit exists and must be a fixed point of $\bar{\Gamma}_{b}$. If $s$ is an arbitrary fixed point of $\bar{\Gamma}_{b}$, we must have $s\geq b$, and thus $s=\bar{\Gamma}_{b}^{n}(s)\geq s^{n}$ for all $n$, implying $s\geq s_{*}(b)$. This shows that $s_{*}(b)$ is the minimal fixed point of $\bar{\Gamma}_{b}$.
We now prove the existence of the maximal fixed point $s^{*}(b)$. To this end, we put $r:=\varphi(\|b\|)$ and consider the point $s^{0}:=s_{*}(r\mathds{1})$ and its trajectory $s^{n}:=\bar{\Gamma}_{b}^{n}(s^{0})$, $n\in\mathbb{N}_{0}$. We have
$s^{1}=b\oplus\Gamma(s_{*}(r\mathds{1}))\leq r\mathds{1}\oplus\Gamma(s_{*}(r\mathds{1}))=s_{*}(r\mathds{1})=s^{0}.$
By monotonicity of $\bar{\Gamma}_{b}$, this implies that the trajectory $(s^{n})_{n\in\mathbb{N}_{0}}$ is decreasing, and thus $s^{n}\stackrel{{\scriptstyle\star}}{{\rightharpoonup}}s^{*}$ for some $s^{*}\in\mathrm{Fix}(\bar{\Gamma}_{b})$. If $s$ is an arbitrary fixed point of $\bar{\Gamma}_{b}$, the $\oplus$-MBI property implies $s\leq\varphi(\|b\|)\mathds{1}\leq s_{*}(\varphi(\|b\|)\mathds{1})=s^{0}$, so that $s=\bar{\Gamma}_{b}^{n}(s)\leq\bar{\Gamma}_{b}^{n}(s^{0})=s^{n}$ for all $n\in\mathbb{N}_{0}$. Hence, $s\leq s^{*}$, showing that $s^{*}=s^{*}(b)$ is the desired maximal fixed point.
It remains to prove (b)–(d). Statement (b) is an easy consequence of monotonicity. Every trajectory starting in $s\leq s_{*}(b)$ is bounded below by the increasing trajectory starting in the origin, which implies (c). Conversely, every trajectory starting in $s\geq s^{*}(b)$ is bounded by a decreasing trajectory as constructed above, implying (d). ∎
Figure 1 illustrates the dynamics of the projected gain operator $\bar{\Gamma}_{b}$ in the weak-topology.
Figure 1: The dynamics of $\bar{\Gamma}_{b}$ in the weak-topology under the $\oplus$-MBI property. The order interval $s_{*}(b),s^{*}(b)]$ is an invariant set which is a weak attractor. (Observe that any $s$ not satisfying $s\leq s_{*}(b)$ or $s\geq s^{*}(b)$ does satisfy $s^{1}\leq s\leq s^{2}$ for some $s^{1}\leq s_{*}(b)$ and $s^{*}(b)\leq s^{2}$.)
3.2 Necessary conditions
From Proposition 2.5, we know that UGS of $\Sigma(\hat{\Gamma}_{\rho})$ for some $\rho\in\mathcal{K}_{\infty}$ is a necessary condition for the existence of a path of strict decay. The following theorem provides further necessary conditions that we obtain from the continuity of the path which we have not used so far.
Theorem 3.1.
If a gain operator $\Gamma$ admits a $C^{0}$-path $\sigma$ of strict decay, the following statements hold for some $\rho\in\mathcal{K}_{\infty}$:
  1. (a)
    $\Sigma(\Gamma_{\rho})$ is UGAS.
  2. (b)
    $\Gamma_{\rho}$ satisfies the $\oplus$-MBI property.
These statements also hold if $\sigma$ only satisfies properties (i), (ii) and (iv) of a $C^{0}$-path of strict decay, i.e. it is not necessary that $\sigma$ is increasing.
Proof.
We let $\tilde{\rho}$ denote the $\mathcal{K}_{\infty}$-function such that $\Gamma_{\tilde{\rho}}\circ\sigma\leq\sigma$. We write $\mathrm{id}+\tilde{\rho}=(\mathrm{id}+\rho^{\prime})\circ(\mathrm{id}+\rho)$ with $\rho,\rho^{\prime}\in\mathcal{K}_{\infty}$, using Lemma A.6. It is then easy to see that $\Gamma_{\rho}(\sigma(r))\leq(\mathrm{id}+\rho^{\prime})^{-1}\circ\sigma(r)$ for all $r\geq 0$. We further write $(\mathrm{id}+\rho^{\prime})^{-1}=\mathrm{id}-\eta$ with $\eta\in\mathcal{K}_{\infty}$, also using Lemma A.6. Then
$\Gamma_{\rho}(\sigma(r))\leq(\mathrm{id}-\eta)(\sigma(r))\mbox{\quad for all\ }r\geq 0.$ (5)
By Proposition 2.2 (b) (which does not require monotonicity), we can assume that for some $\varphi\in\mathcal{K}_{\infty}$ it holds that
$r\mathds{1}\leq\sigma(r)\leq\varphi(r)\mathds{1}\mbox{\quad for all\ }r\geq 0.$ (6)
[…]
$\Gamma_{\rho}(\sigma(r))\leq\sigma(r)-\eta(\sigma(r))\leq\sigma(r)-\eta(r_{*})\mathds{1}=\sigma(r)-\varepsilon\mathds{1}.$
Since $\sigma(r_{n})$ converges to $\sigma(r_{*})$ by continuity of $\sigma$, we can choose $n_{0}$ large enough so that $\|\sigma(r_{n})-\sigma(r_{n+1})\|\leq\varepsilon/2$ for all $n\geq n_{0}$, implying
$s^{n+1}=\Gamma_{\rho}(s^{n})\leq\Gamma_{\rho}(\sigma(r_{n}))\leq\sigma(r_{n})-\varepsilon\mathds{1}\leq\sigma(r_{n+1})-\frac{\varepsilon}{2}\mathds{1}.$
However, the inequality $s^{n+1}+(\varepsilon/2)\mathds{1}\leq\sigma(r_{n+1})$ contradicts the minimality of $r_{n+1}$. Indeed, for any $r<r_{n+1}$ we must have $s^{n+1}_{i}>\sigma_{i}(r)$ for some $i\in\mathcal{I}$. Thus, $\|\sigma(r)-\sigma(r_{n+1})\|\geq\sigma_{i}(r_{n+1})-\sigma_{i}(r)>\varepsilon/2$, which contradicts the continuity of $\sigma$ for $r$ close enough to $r_{n+1}$.
(b) Let $s\in\ell^{\infty}_{+}(\mathcal{I})$ and $r\geq 0$ be chosen such that $s\leq r\mathds{1}\oplus\Gamma_{\rho}(s)$. The case $s=0$ is trivial, so we assume that $s>0$. We define
$r^{\prime}:=\min\{\tilde{r}\geq 0:s\leq\sigma(\tilde{r})\}>0.$
Again, the minimum exists due to continuity. From the monotonicity of $\bar{\Gamma}_{\rho,r\mathds{1}}$ and $s\leq\sigma(r^{\prime})$, it follows that
$s\leq\bar{\Gamma}_{\rho,r\mathds{1}}(s)\leq\bar{\Gamma}_{\rho,r\mathds{1}}(\sigma(r^{\prime}))=r\mathds{1}\oplus\Gamma_{\rho}(\sigma(r^{\prime}))\leq r\mathds{1}\oplus(\mathrm{id}-\eta)(\sigma(r^{\prime})).$ (7)
We pick $\varepsilon\in(0,\eta(r^{\prime}))$. Since $r^{\prime}$ was chosen minimal, there exists an index $i\in\mathcal{I}$ with $\sigma_{i}(r^{\prime})-s_{i}<\varepsilon$. Evaluating (7) at this index yields
$\sigma_{i}(r^{\prime})-\varepsilon<s_{i}\leq\max\{r,\sigma_{i}(r^{\prime})-\eta(\sigma_{i}(r^{\prime}))\}\leq\max\{r,\sigma_{i}(r^{\prime})-\varepsilon\}.$
To avoid a contradiction, the maximum at the end of the line must be attained at $r$. Since $r^{\prime}\leq\sigma_{i}(r^{\prime})$, this implies $r^{\prime}-\varepsilon\leq r$. Sending $\varepsilon$ to zero, we obtain $r^{\prime}\leq r$, and thus $s\leq\sigma(r^{\prime})\leq\varphi(r^{\prime})\mathds{1}\leq\varphi(r)\mathds{1}$, which implies $\|s\|\leq\varphi(r)$. Hence, $\Gamma_{\rho}$ satisfies the $\oplus$-MBI property.∎
Remark 3.3.
In general, it is not clear how the two properties in (a) and (b) are related to each other. What we know for sure is that the $\oplus$-MBI property does not imply UGAS in the general infinite case. This can be seen by looking at max-type gain operators. Proposition 4.1 below shows that for these operators the $\oplus$-MBI property is equivalent to UGS plus GATT, which is strictly weaker than UGAS in the general infinite case.
Remark 3.4.
The proof of (b) has an interesting conclusion: If $s$ is a fixed point of $\bar{\Gamma}_{r\mathds{1}}$ and $\sigma$ is a $C^{0}$-path of strict decay with $\varphi_{\min}=\mathrm{id}$, then $s\leq\sigma(r)$. This yields an upper bound on every fixed point of $\bar{\Gamma}_{r\mathds{1}}$, namely
$s\leq\inf_{\sigma(\cdot)}\sigma(r),$
where the (componentwise) infimum runs over all such $C^{0}$-paths.
3.3 Three path construction methods
So far, we have not provided any method to construct a $C^{0}$-path of strict decay except for the one appearing in the proof of Proposition 2.5, where we did not show continuity, however. In this subsection, we discuss three different approaches.
3.3.1 The minimal path
Assuming that $\Gamma$ satisfies the $\oplus$-MBI property, we define a candidate path by
$\sigma_{*}(r):=s_{*}(r\mathds{1})=\bigoplus_{n=0}^{\infty}\hat{\Gamma}^{n}(r\mathds{1})\mbox{\quad for all\ }r\geq 0.$
Note that this is precisely the path defined in (4). Since $s_{*}(r\mathds{1})$ is the minimal fixed point of $\bar{\Gamma}_{r\mathds{1}}$, it is also the minimal element of $\Psi(\bar{\Gamma}_{r\mathds{1}})$, and hence $\sigma_{*}$ is the minimal $C^{0}$-path of decay satisfying $\sigma(r)\geq r\mathds{1}$ for all $r$, provided that it is a $C^{0}$-path.
From the proof of Proposition 2.5, we already know that $\sigma_{*}$ satisfies properties (i)–(iii) of a $C^{0}$-path of decay with $\varphi_{\min}=\mathrm{id}$. It thus only remains to prove continuity to show that $\sigma_{*}$ is a $C^{0}$-path of decay. The following proposition yields a sufficient condition.
Proposition 3.3.
The candidate path $\sigma_{*}$ has the following properties:
  1. (a)
    If $\sigma_{*}(r)$ is the only fixed point of $\bar{\Gamma}_{r\mathds{1}}$ for each $r\geq 0$, then every component function $\sigma_{*i}$, $i\in\mathcal{I}$, is continuous. In particular, $\sigma_{*}$ is sequentially continuous in the weak-topology.
  2. (b)
    If $\sigma_{*}(r)$ is a globally attractive fixed point of $\bar{\Gamma}_{r\mathds{1}}$ for each $r\geq 0$, then $\sigma_{*}$ is continuous.
Remark 3.5.
The assumption in (b) means that the invariant order interval $s_{*}(r\mathds{1}),s^{*}(r\mathds{1})]$, depicted in Fig. 1, is a single point $\sigma_{*}(r)$ and convergence to $\sigma_{*}(r)$ holds not only in the weak-topology, but in the norm-topology. Since in the finite case these topologies coincide, attractivity is not needed as an assumption in (a) as it is automatically satisfied.
Proof.
(a) From $r\mathds{1}\leq\sigma_{*}(r)\leq\varphi(r)\mathds{1}$, it follows that $\sigma_{*}$ (and hence each $\sigma_{*i}$) is continuous at $r=0$. We now fix $r>0$ and prove that each $\sigma_{*i}$ is continuous at $r$. It suffices to prove left-sided and right-sided continuity separately. Let $(r_{n})_{n\in\mathbb{N}}$ be a sequence which approaches $r$ from the left. Without loss of generality, we can assume that the sequence is increasing. Then $\sigma_{*}(r_{n})\leq\sigma_{*}(r_{n+1})\leq\sigma_{*}(r)$ for all $n$, and thus $s_{i}:=\lim_{n\rightarrow\infty}\sigma_{*i}(r_{n})$ exists for each $i\in\mathcal{I}$. We have
$r_{n}\mathds{1}\oplus\Gamma(\sigma_{*}(r_{n}))=\sigma_{*}(r_{n})\mbox{\quad for all\ }n\in\mathbb{N}.$
Taking the componentwise limit on both sides and using sequential weak continuity yields $r\mathds{1}\oplus\Gamma(s)=s$ for $s=(s_{i})_{i\in\mathcal{I}}$. Hence, $s$ is a fixed point of $\bar{\Gamma}_{r\mathds{1}}$, implying $s=\sigma_{*}(r)$. This shows that $\sigma_{*i}(r_{n})$ converges to $\sigma_{*i}(r)$ for each $i\in\mathcal{I}$. The proof works analogously if $(r_{n})_{n\in\mathbb{N}}$ approaches $r$ from the right.
[…]
Remark 3.6.
Observe that the assumption that $\sigma_{*}(r)$ is a globally attractive fixed point of $\bar{\Gamma}_{r\mathds{1}}$, considered for $r=0$, implies that $\Sigma(\Gamma)$ is UGAS.
We obtain the following theorem as an immediate conclusion of Proposition 3.3 in combination with Proposition 2.3.
Theorem 3.2.
Given a gain operator $\Gamma$, assume that there is $\rho\in\mathcal{K}_{\infty}$ such that $\Gamma_{\rho}$ satisfies the $\oplus$-MBI property and the point $\sigma_{*,\rho}(r)=\bigoplus_{n=0}^{\infty}\hat{\Gamma}_{\rho}^{n}(r\mathds{1})$ is a globally attractive fixed point of $\bar{\Gamma}_{\rho,r\mathds{1}}$ for each $r\geq 0$. Then there exists a path of strict decay for $\Gamma$.
Remark 3.7.
The assumption that $\Gamma_{\rho}$ satisfies the $\oplus$-MBI property can be weakened to the assumption that $\Sigma(\hat{\Gamma}_{\rho})$ is UGS (using Corollary 3.1).
Remark 3.8.
The assumption of the theorem is closely related to the notion of static Input/State characteristics for monotone control systems, introduced in [1]. To understand this connection, we have to regard
$s^{k+1}=b^{k}\oplus\Gamma(s^{k})$ (8)
as a control system on $\ell^{\infty}_{+}(\mathcal{I})$ with inputs $b^{k}\in\ell^{\infty}_{+}(\mathcal{I})$. This control system is monotone in the sense of [1], which means that solutions behave jointly monotonically with respect to initial conditions and inputs, see [1, Def. II.1]. A static Input/State characteristic of (8) is a function $k_{s}:\ell^{\infty}_{+}(\mathcal{I})\rightarrow\ell^{\infty}_{+}(\mathcal{I})$, which assigns to each constant control input $b$ a globally attractive fixed point $k_{s}(b)$, which is just what we require to exist for inputs of the form $b=r\mathds{1}$. We also mention [21], which considers the control system (8) and relates its properties to that of the given continuous-time system (for which ISS is to be proved).
The assumption that $\bar{\Gamma}_{\rho,r\mathds{1}}$ has a unique globally attractive fixed point for each $r$ is probably a restrictive one, which is not necessary for the existence of a path of strict decay. However, as the following proposition shows, the uniqueness of the fixed point is necessary for the continuity of the candidate path, at least under some technical assumptions. A hint that this should be true is already given by Remark 3.4, which shows that if $\sigma_{*,\rho^{\prime}}$ is continuous for some $\rho^{\prime}>\rho$, then any fixed point $s$ of $\bar{\Gamma}_{\rho,r\mathds{1}}$ must satisfy $\sigma_{*,\rho}(r)\leq s\leq\sigma_{*,\rho^{\prime}}(r)$.
Proposition 3.4.
Let the following assumptions hold:
  1. (i)
    $\Gamma$ satisfies the $\oplus$-MBI property.
  2. (ii)
    There exists $\rho\in\mathcal{K}_{\infty}$ such that $\Sigma(\Gamma_{\rho})$ is UGAS.
  3. (iii)
    The family $\{\gamma_{ij}^{-1}:ji\in E(\mathcal{G})\}$ is pointwise equicontinuous.
  4. (iv)
    There exists $\zeta\in\mathcal{K}_{\infty}$ such that for all $ji\in E(\mathcal{G})$ and $s^{1}\leq s^{2}$
    $\mu_{i}(s^{2}_{|\mathcal{I}_{i}})-\mu_{i}(s^{1}_{|\mathcal{I}_{i}})\geq\zeta(s^{2}_{j}-s^{1}_{j}).$
  5. (v)
    $\sigma_{*}$ is continuous.
Then $\bar{\Gamma}_{r\mathds{1}}$ has precisely one fixed point for each $r\geq 0$.
Proof.
We prove the following claim: Whenever $r_{1}<r_{2}$, then
$\exists n\in\mathbb{N}:\ \bar{\Gamma}_{r_{1}\mathds{1}}^{n}(\sigma_{*}(r_{2}))\ll\sigma_{*}(r_{2}).$ (9)
Assume that this claim holds and let $s$ be a fixed point of $\bar{\Gamma}_{r_{1}\mathds{1}}$ different from $\sigma_{*}(r_{1})$, where $r_{1}\geq 0$ is given. By Proposition 3.3 (a), we must have $s>\sigma_{*}(r_{1})$. Now, we put $r_{2}:=\min\{r\geq 0:s\leq\sigma_{*}(r)\}$. The existence of the minimum follows from assumption (v). We then have $r_{2}>r_{1}$ and thus there exists $n$ with
$s=\bar{\Gamma}_{r_{1}\mathds{1}}^{n}(s)\leq\bar{\Gamma}_{r_{1}\mathds{1}}^{n}(\sigma_{*}(r_{2}))\ll\sigma_{*}(r_{2}).$
However, since $r_{2}$ was chosen minimal with $s\leq\sigma_{*}(r_{2})$, we cannot have $s\ll\sigma_{*}(r_{2})$, so the existence of $s$ yields a contradiction.
It remains to prove (9). To do this, we use assumption (ii) and Proposition 2.8. Let us write $(\mathrm{id}+\rho)^{-1}=\mathrm{id}-\eta$ for some $\eta\in\mathcal{K}_{\infty}$, using Lemma A.6. Then
$\exists n\in\mathbb{N}:\ \forall i\in\mathcal{I}\ \exists j\in\mathcal{N}^{-}_{i}(n)\mbox{ with }\Gamma_{j}(\sigma_{*}(r_{2}))<(\mathrm{id}-\eta)(\sigma_{*j}(r_{2})).$
If $j$ is an index as above, we necessarily have $\sigma_{*j}(r_{2})=r_{2}$ and
$\bar{\Gamma}_{r_{1}\mathds{1},j}(\sigma_{*}(r_{2}))$ $=\max\{r_{1},\Gamma_{j}(\sigma_{*}(r_{2}))\}$
$\leq\max\{r_{1},r_{2}-\eta(r_{2})\}=\sigma_{*j}(r_{2})-\underbrace{\min\{r_{2}-r_{1},\eta(r_{2})\}}_{=:\alpha_{1}>0}.$
Now, assume that $ji\in E(\mathcal{G})$. Then
$\Gamma^{2}_{r_{1}\mathds{1},i}(\sigma_{*}(r_{2}))$ $=\max\{r_{1},\Gamma_{i}(\bar{\Gamma}_{r_{1}\mathds{1}}(\sigma_{*}(r_{2})))\}$
$=\max\{r_{1},\mu_{i}([\gamma_{ik}(\bar{\Gamma}_{r_{1}\mathds{1},k}(\sigma_{*}(r_{2})))]_{k\in\mathcal{I}_{i}})\}.$
We have
$\gamma_{ik}(\bar{\Gamma}_{r_{1}\mathds{1},k}(\sigma_{*}(r_{2})))\leq\left\{\begin{array}[]{cc}\gamma_{ik}(\sigma_{*k}(r_{2}))&\mbox{if }k\neq j,\\ \gamma_{ik}(\sigma_{*k}(r_{2})-\alpha_{1})&\mbox{if }k=j.\end{array}\right$.
By assumption (iii) and Lemma A.4, we can find some $\beta>0$ such that
$\gamma_{ij}(\sigma_{*j}(r_{2})-\alpha_{1})\leq\gamma_{ij}(\sigma_{*j}(r_{2}))-\beta,$
no matter how we choose $j$ and subsequently $i$. Assumption (iv) then implies
$\mu_{i}([\gamma_{ik}(\bar{\Gamma}_{r_{1}\mathds{1},k}(\sigma_{*}(r_{2})))]_{k\in\mathcal{I}_{i}})\leq\mu_{i}([\gamma_{ik}(\sigma_{*k}(r_{2}))]_{k\in\mathcal{I}_{i}})-\zeta(\beta).$
Altogether, we obtain
$\bar{\Gamma}^{2}_{r_{1}\mathds{1},i}(\sigma_{*}(r_{2}))$ $\leq\max\{r_{1},\Gamma_{i}(\sigma_{*}(r_{2}))-\zeta(\beta)\}\leq\max\{r_{1},\sigma_{*i}(r_{2})-\zeta(\beta)\}$
$=\sigma_{*i}(r_{2})-\min\{\sigma_{*i}(r_{2})-r_{1},\zeta(\beta)\}$
$\leq\sigma_{*i}(r_{2})-\underbrace{\min\{r_{2}-r_{1},\zeta(\beta)\}}_{=:\alpha_{2}>0}.$
We can continue in this manner, next looking at some index in $\mathcal{N}^{-}_{j}(2)$. After $n$ iterations, we will obtain $\bar{\Gamma}^{n}_{r_{1}\mathds{1}}(\sigma_{*}(r_{2}))\leq\sigma_{*}(r_{2})-\alpha\mathds{1}$ with $\alpha=\min\{\alpha_{1},\ldots,\alpha_{n}\}$, which implies the validity of the claim. ∎
Remark 3.9.
Assumption (iv) of the theorem holds, in particular, for sum-type operators. In this case, $\mu_{i}(s)=\sum_{j\in\mathcal{I}}s_{j}$ for all $i\in\mathcal{I}$, and thus
$\mu_{i}(s^{2}_{|\mathcal{I}_{i}})-\mu_{i}(s^{1}_{|\mathcal{I}_{i}})=\sum_{j\in\mathcal{I}_{i}}(s^{2}_{j}-s^{1}_{j}),$
so that assumption (iv) holds with $\zeta=\mathrm{id}$. However, the assumption does not hold for max-type operators.
3.3.2 The linear interpolation method
In the finite case, under the assumptions that the graph $\mathcal{G}$ is strongly connected and the system $\Sigma(\Gamma)$ is UGAS, one can construct a path of decay essentially by linear interpolation between the elements of a complete orbit $\{s^{n}:n\in\mathbb{Z}\}$ of $\Gamma$ (i.e. $s^{n+1}\equiv\Gamma(s^{n})$). In this section, we explain why this construction works and which assumption is necessary to extend it to the general case.
First, observe that if $\sigma$ is a path of decay for $\Gamma$, then the coercivity property implies $\sigma(r)\geq\varphi_{\min}\circ\varphi_{\max}^{-1}(\|\sigma(r)\|)\mathds{1}$ for all $r\geq 0$, showing that the image of $\sigma$ is a coercive set. Hence, for the linear interpolation method to work, there must exist some $s^{0}\in\mathrm{int}(\ell^{\infty}_{+}(\mathcal{I}))\cap\Psi(\Gamma)$ such that the corresponding forward orbit is a coercive set. Lemma 4.1 shows that in the finite irreducible case this condition is completely unproblematic, as here the whole decay set is coercive.
To prove the main result of this subsection, we need the following lemmas.
Lemma 3.1.
[…]
Proof.
For a given $s\in\ell^{\infty}_{+}(\mathcal{I})$, choose a sequence $(j_{n})$ in $\mathbb{N}$ such that $s_{j_{n}}\rightarrow\|s\|$ for $n\rightarrow\infty$. By assumption, there exist indices $i_{n}$ with $j_{n}\in\mathcal{I}_{i_{n}}$ for all $n$. Hence,
$\|\Gamma(s)\|\geq\sup_{n\in\mathbb{N}}\Gamma_{i_{n}}(s)\geq\sup_{n\in\mathbb{N}}(\xi\circ\eta)(s_{j_{n}})=(\xi\circ\eta)(\|s\|).$
The proof is complete. ∎
Lemma 3.2.
Let $\sigma$ be a path in $\ell^{\infty}_{+}(\mathcal{I})$ satisfying the following properties:
  1. (i)
    $\sigma(r)\in\Psi(\Gamma)$ for all $r\geq 0$.
  2. (ii)
    $\sigma(0)=0$, $\|\sigma(r)\|>0$ for all $r>0$, and $\|\sigma(r)\|\rightarrow\infty$ as $r\rightarrow\infty$.
  3. (iii)
    $\sigma$ is increasing and continuous.
  4. (iv)
    The image of $\sigma$ is coercive with some $\underline{\varphi}\in\mathcal{K}_{\infty}$.
Then $\sigma$ is a $C^{0}$-path of decay for $\Gamma$.
Proof.
We only have to show that $\varphi_{\min}(r)\mathds{1}\leq\sigma(r)\leq\varphi_{\max}(r)\mathds{1}$ for some $\varphi_{\min},\varphi_{\max}\in\mathcal{K}_{\infty}$. If we put $\tilde{\varphi}_{\min}(r):=\underline{\varphi}(\|\sigma(r)\|)$ and $\tilde{\varphi}_{\max}(r):=\|\sigma(r)\|$, we see that $\tilde{\varphi}_{\min}(r)\mathds{1}\leq\sigma(r)\leq\tilde{\varphi}_{\max}(r)\mathds{1}$. Moreover, by assumption (ii), we have $\tilde{\varphi}_{\min}(0)=\tilde{\varphi}_{\max}(0)=0$, both functions are continuous (by assumption (iii)) and converge to $\infty$ as $r\rightarrow\infty$, and finally are positive and increasing on $(0,\infty)$ (also by assumption (iii)). Hence, we can lower-bound $\tilde{\varphi}_{\min}$ by a $\mathcal{K}_{\infty}$-function $\varphi_{\min}$ and upper-bound $\tilde{\varphi}_{\max}$ by a $\mathcal{K}_{\infty}$-function $\varphi_{\max}$, using Lemma A.7. ∎
Lemma 3.3.
[…]
$A^{\prime}:=\bigcup_{s\in A}[\Gamma(s),s].$
If $A$ is weakly sequentially closed, then so is $A^{\prime}$.
Proof.
From Lemma A.9 and $A\subset\Psi(\Gamma)$, it follows that $A^{\prime}\subset\Psi(\Gamma)$. If $s^{\prime}\in A^{\prime}$, then $\Gamma(s)\leq s^{\prime}\leq s$ for some $s\in A$. This implies $\Gamma^{2}(s)\leq\Gamma(s^{\prime})\leq\Gamma(s)$. As $A$ is forward-invariant, we have $\Gamma(s)\in A$, implying $\Gamma(s^{\prime})\in A^{\prime}$. This shows that $A^{\prime}$ is forward-invariant. Now, assume that $s\geq\underline{\varphi}(\|s\|)\mathds{1}$ for all $s\in A$ with some $\underline{\varphi}\in\mathcal{K}_{\infty}$. If $s^{\prime}\in A^{\prime}$ and $\Gamma(s)\leq s^{\prime}\leq s$, $s\in A$, then
$s^{\prime}\geq\Gamma(s)\geq\underline{\varphi}(\|\Gamma(s)\|)\mathds{1}\geq\underline{\varphi}\circ\zeta(\|s\|)\mathds{1}\geq\underline{\varphi}\circ\zeta(\|s^{\prime}\|),$
[…]
Now, we can prove the main result of this subsection.
Theorem 3.3.
Given a gain operator $\Gamma$, let the following assumptions hold:
  1. (i)
    There exists a subset $A\subset\Psi(\Gamma)$, which is weakly closed, forward-invariant, cofinal and coercive.
  2. (ii)
    The system $\Sigma(\Gamma)$ is GATT.
  3. (iii)
[…]
Then there exists $s^{0}\in\mathrm{int}(\ell^{\infty}_{+}(\mathcal{I}))\cap\Psi(\Gamma)$ which admits a complete orbit $\{s^{n}:n\in\mathbb{Z}\}$ such that the path obtained by linear interpolation between the points $s^{n}$ is a $C^{0}$-path of decay for $\Gamma$.
Proof.
The proof proceeds in five steps.
Step 1: Assumption (iii) together with Lemma 3.1 implies
$\|s\|\rightarrow\infty\quad\Rightarrow\quad\|\Gamma(s)\|\rightarrow\infty.$ (10)
We also show:
$\mbox{If }A\subset\ell^{\infty_{+}}\mbox{ is cofinal, then }\Gamma(A)\mbox{ is cofinal}.$ (11)
To this end, let $s\in\ell^{\infty}_{+}$ and choose $\hat{s}\in A$ with $\hat{s}\geq(\xi\circ\eta)^{-1}(\|s\|)\mathds{1}$. Then, for each index $i$, we find $j\in\mathcal{I}_{i}$ and thus have the estimate $\Gamma_{i}(\hat{s})\geq\Gamma_{i}((\xi\circ\eta)^{-1}(\|s\|)\mathds{1})\geq\|s\|$. This implies $\Gamma(\hat{s})\geq\|s\|\mathds{1}\geq s$, which concludes the proof of (11).
Step 2: We prove that the set $\Gamma^{k}(A)$ is weakly sequentially closed for each $k\in\mathbb{N}_{0}$. For $k=0$, this is true by assumption. Now, let $k\geq 1$ and pick a sequence $(s^{n})_{n\in\mathbb{N}}$ in $\Gamma^{k}(A)$, which weakly converges to some $s\in\ell^{\infty}_{+}(\mathcal{I})$. Then there exists a sequence $(\tilde{s}^{n})_{n\in\mathbb{N}}$ in $A$ such that
$s^{n}=\Gamma^{k}(\tilde{s}^{n})\mbox{\quad for all\ }n\in\mathbb{N}.$
[…]
Step 3: We define the set
$A_{\infty}:=\bigcap_{n=0}^{\infty}\Gamma^{n}(A)$
and prove that $\Gamma(A_{\infty})=A_{\infty}$. The inclusion $\Gamma(A_{\infty})\subset A_{\infty}$ is trivial:
$\Gamma(A_{\infty})=\Gamma\Bigl(\bigcap_{n=0}^{\infty}\Gamma^{n}(A)\Bigr)\subset\bigcap_{n=0}^{\infty}\Gamma^{n+1}(A)=A_{\infty}.$
[…]
$s=\Gamma(\hat{s}^{n}),\quad\hat{s}^{n}:=\Gamma^{n-1}(s^{n}).$
[…]
To conclude the proof, it remains to show that $\hat{s}\in A_{\infty}$. This is equivalent to $\hat{s}\in\Gamma^{n}(A)$ for all $n\geq 0$. Since $\Gamma^{n+1}(A)\subset\Gamma^{n}(A)$ for all $n$, this in turn is equivalent to $\hat{s}\in\Gamma^{n_{m}-1}(A)$ for all $m$. By construction, we have
$\hat{s}^{n_{m+k}}\in\Gamma^{n_{m}-1}(A)\mbox{\quad for all\ }m,k.$
Since $\Gamma^{n_{m}-1}(A)$ is weakly sequentially closed by Step 2, we have $\hat{s}\in\Gamma^{n_{m}-1}(A)$ for each $m$, concluding the proof.
Step 4: We prove that $A^{\prime}_{\infty}\cap\mathrm{int}(\ell^{\infty}_{+}(\mathcal{I}))$ is nonempty, where $A^{\prime}$ is constructed as in Lemma 3.3. To this end, we fix a closed order interval of the form
$\mathcal{C}:=[\underline{\varphi}(a)\mathds{1},a\mathds{1}],\quad a>0,$
where $\underline{\varphi}\in\mathcal{K}_{\infty}$ is chosen with $s\geq\underline{\varphi}(\|s\|)\mathds{1}$ for all $s\in A^{\prime}$, and show that it has nonempty intersection with each of the sets $\Gamma^{n}(A)$, $n\in\mathbb{N}_{0}$. To this end, observe that each $\Gamma^{n}(A)$ is cofinal. For $n=0$, this is true by assumption and for $n>0$ it follows from (11). Hence, for each $n$, we can find some $\hat{s}=\Gamma^{n}(s)$ with $s\in A$ and $\hat{s}\geq a\mathds{1}$. We can connect $s$ with 0 by a piecewise linear path $\sigma:[0,1]\rightarrow\ell^{\infty}_{+}(\mathcal{I})$ obtained by linear interpolation between the points of the trajectory starting in $s$ (using GATT). By construction, this path completely evolves in $A^{\prime}$. Then $\hat{\sigma}:=\Gamma^{n}\circ\sigma$ is a continuous path in $\Gamma^{n}(A^{\prime})\subset A^{\prime}$, connecting $\hat{s}$ with 0. Thus,
$\hat{\sigma}(r)\geq\underline{\varphi}(\|\hat{\sigma}(r)\|)\mathds{1}\mbox{\quad for all\ }r\in[0,1].$
Since the auxiliary function $f(r):=\|\hat{\sigma}(r)\|$ is continuous with $f(0)\geq a$ and $f(1)=0$, there exists an $r_{0}$ such that $\|\hat{\sigma}(r_{0})\|=f(r_{0})=a$. It then follows that $\underline{\varphi}(a)\mathds{1}\leq\hat{\sigma}(r_{0})\leq a\mathds{1}$, implying $\sigma(r_{0})\in\mathcal{C}$. Hence, we have found a point in $\mathcal{C}\cap\Gamma^{n}(A^{\prime})$. Now, to conclude Step 4, let $s^{n}\in\mathcal{C}\cap\Gamma^{n}(A^{\prime})$ for each $n$. Since $\mathcal{C}$ is weakly sequentially compact, there exists a weakly convergent subsequence $(s^{n_{m}})_{m\in\mathbb{N}_{0}}$ with limit $s^{*}\in\mathcal{C}$. Since the sets $\Gamma^{n}(A^{\prime})$ are nested and sequentially weakly closed, we obtain $s^{*}\in A^{\prime}_{\infty}$.
Step 5: We construct the actual $C^{0}$-path $\sigma:\mathbb{R}_{+}\rightarrow\ell^{\infty}_{+}(\mathcal{I})$. To this end, fix $s^{0}\in\mathcal{C}\cap A^{\prime}_{\infty}$. By Step 3, we can write $s^{0}=\Gamma(s^{-1})$ for some $s^{-1}\in A^{\prime}_{\infty}$. Again, we can write $s^{-1}=\Gamma(s^{-2})$ with $s^{-2}\in A^{\prime}_{\infty}$. Iterating this process, we obtain a sequence $(s^{-n})_{n\in\mathbb{N}_{0}}$ in $A^{\prime}$ with $\Gamma(s^{n})\equiv s^{n+1}$. We can extend this sequence in the other direction by putting $s^{n}:=\Gamma(s^{0})$ for all $n>0$. The double-sided sequence $(s^{n})_{n\in\mathbb{Z}}$ obtained in this way satisfies $s^{n+1}\leq s^{n}$ for all $n\in\mathbb{Z}$, which directly follows from $s^{n+1}=\Gamma(s^{n})\leq s^{n}$. Moreover, assumption (ii) yields $s^{n}\rightarrow 0$ for $n\rightarrow\infty$. Assume to the contrary that $\|s^{n}\|$ is bounded. Then $s^{n}\stackrel{{\scriptstyle\star}}{{\rightharpoonup}}s^{*}$ for some $s^{*}$ as $n\rightarrow$-$\infty$. Since $s^{n+1}\equiv\Gamma(s^{n})$, this implies $s^{*}=\Gamma(s^{*})$, contradicting global attractivity. Hence,
$\lim_{n\rightarrow-\infty}\|s^{n}\|=\infty.$
Interpolating linearly between the points $s^{n}$ then yields a path $\sigma:(0,\infty)\rightarrow\ell^{\infty}_{+}(\mathcal{I})$ that we can extend continuously by $\sigma(0):=0$. The extended path $\sigma$ has the following properties:
  • $\sigma(r)\in\Psi(\Gamma)$ for all $r\geq 0$ by the choice of $s^{n}$ and Lemma A.9.
  • $\sigma(0)=0$, $\|\sigma(r)\|>0$ for all $r>0$ by the choice of $s^{0}$ and assumption (iii), and $\|\sigma(r)\|\rightarrow\infty$ as $r\rightarrow\infty$.
  • $\sigma$ is increasing and continuous.
Hence, by Lemma 3.2 and assumption (i), $\sigma$ is a $C^{0}$-path of decay for $\Gamma$. ∎
Remark 3.10.
We can easily see that assumption (i) of the theorem is necessary if we want to construct a $C^{0}$-path by linear interpolation between the points of a complete orbit. This is so, because if a corresponding orbit $A:=\{s^{n}:n\in\mathbb{Z}\}$ exists, then $A\cup\{0\}$ is weakly closed, forward-invariant by construction and cofinal as well as coercive by requirement. Hence, a set as required in assumption (i) must exist.
3.3.3 A combined method
In this subsection, we introduce a method for path construction that combines the ideas used in the two previous methods. We assume again that $\Gamma$ satisfies the $\oplus$-MBI property, which, in particular, implies that the candidate path $\sigma_{*}:\mathbb{R}_{+}\rightarrow\ell^{\infty}_{+}(\mathcal{I})$ is well-defined. We recall that $\sigma_{*}(r)$ is the minimal fixed point of the operator $\bar{\Gamma}_{r\mathds{1}}$. We fix a function $\varphi\in\mathcal{K}_{\infty}$ such that
$s\leq\bar{\Gamma}_{b}(s)\quad\Rightarrow\quad\|s\|\leq\varphi(\|b\|)\mbox{\quad for all\ }s,b\in\ell^{\infty}_{+}(\mathcal{I}).$
From Proposition 3.2, we know that $\bar{\Gamma}_{r\mathds{1}}$ also has a maximal fixed point
$\sigma^{*}(r):=s^{*}(r\mathds{1}).$
The new path construction method starts by fixing a strictly increasing double-sided sequence of real numbers $(r_{k})_{k\in\mathbb{Z}}$ with
[…]
We also consider the associated sequence of points $s^{k}:=\sigma^{*}(r_{k})$, which has the following properties for all $k\in\mathbb{Z}$:
  • $s^{k}\leq s^{k+1}$.
  • $r_{k}\mathds{1}\leq s^{k}\leq\varphi(r_{k})\mathds{1}$.
  • $s^{k}\in\Psi(\Gamma)$, since $\mathrm{Fix}(\bar{\Gamma}_{r\mathds{1}})\subset\Psi(\bar{\Gamma}_{r\mathds{1}})\subset\Psi(\Gamma)$.
Our goal is to construct a $C^{0}$-path of decay $\sigma:\mathbb{R}_{+}\rightarrow\ell^{\infty}_{+}(\mathcal{I})$ which satisfies $\sigma(r_{k})=s^{k}$ for all $k\in\mathbb{Z}$. Hence, we have to define $\sigma$ on the open intervals $(r_{k},r_{k+1})$ and at $r=0$. Here, we use statements (b) and (d) of Proposition 3.2, which show that $s^{k,n}:=\bar{\Gamma}_{r_{k}\mathds{1}}^{n}(s^{k+1})\stackrel{{\scriptstyle\star}}{{\rightharpoonup}}s^{k}$. We define the path on $(r_{k},r_{k+1})$ by linear interpolation between the points $s^{k+1}=s^{k,0}$, $s^{k,1}$, $s^{k,2}$, and so on, and put $\sigma(0):=0$. In this way, we obtain a path $\sigma:\mathbb{R}_{+}\rightarrow\ell^{\infty}_{+}(\mathcal{I})$, satisfying the properties listed in the following proposition. Note that the parametrization of the path on the intervals $(r_{k},r_{k+1})$ is not unique.
Proposition 3.5.
The path $\sigma$ satisfies properties (i)–(iii) of a $C^{0}$-path of decay. Additionally, it has the following continuity properties:
  1. (C1)
    It is norm-continuous from below, i.e., if $(\rho_{n})_{n\in\mathbb{N}}$ is a sequence in $\mathbb{R}_{+}$ converging to some $\rho$ from below, then $\|\sigma(\rho_{n})-\sigma(\rho)\|\rightarrow 0$.
  2. (C2)
    It is sequentially weakly continuous from above, i.e., if $(\rho_{n})_{n\in\mathbb{N}}$ is a sequence in $\mathbb{R}_{+}$ converging to some $\rho$ from above, then $\sigma(\rho_{n})\stackrel{{\scriptstyle\star}}{{\rightharpoonup}}\sigma(\rho)$.
Proof.
Property (i) (decay): The points $s^{k}=\sigma(r_{k})$ are points of decay of $\bar{\Gamma}_{r_{k}\mathds{1}}$, hence also points of decay of $\Gamma$. Observe that $s^{k+1}$ is also a point of decay for $\bar{\Gamma}_{r_{k}\mathds{1}}$. By Lemma A.9, then also the points $\sigma(r)$ with $r\in(r_{k},r_{k+1})$ are contained in the decay set of $\bar{\Gamma}_{r_{k}\mathds{1}}$, and thus in $\Psi(\Gamma)$.
Property (ii) (coercivity): The piecewise constant functions
$\tilde{\varphi}_{\min}(r):=\left\{\begin{array}[]{rl}r_{k}&\mbox{if }r\in[r_{k},r_{k+1})\\ 0&\mbox{if }r=0\end{array}\right$.
and
$\tilde{\varphi}_{\max}(r):=\left\{\begin{array}[]{rl}\varphi(r_{k+1})&\mbox{if }r\in[r_{k},r_{k+1})\\ 0&\mbox{if }r=0\end{array}\right$.
can be used to lower-bound and upper-bound the components of $\sigma$, respectively. By Lemma A.7, the existence of $\varphi_{\min},\varphi_{\max}\in\mathcal{K}_{\infty}$ with $\varphi_{\min}\leq\tilde{\varphi}_{\min}$ and $\tilde{\varphi}_{\max}\leq\varphi_{\max}$ follow.
Property (iii) (monotonicity): This is clear by construction.
Property (C1): Norm-continuity of $\sigma$ is obvious at any $\rho\notin\{r_{k}:k\in\mathbb{Z}\}$. If $(\rho_{n})_{n\in\mathbb{N}}$ converges to some $r_{k}$ from below (i.e. from the left), then $\rho_{n}\in(r_{k-1},r_{k}]$ for all sufficiently large $n$. But on $(r_{k-1},r_{k}]$, $\sigma$ is piecewise linear, hence norm-continuous.
Property (C2): Again, we only have to consider the case $\rho=r_{k}$ for some $k\in\mathbb{Z}$. If $(\rho_{n})_{n\in\mathbb{N}}$ converges to $r_{k}$ from above (i.e. from the right), for all sufficiently large $n$ we can find $m_{n}$ such that $\sigma(r_{k})\leq\sigma(\rho_{n})\leq\bar{\Gamma}_{r_{k}\mathds{1}}^{m_{n}}(s^{k+1})$, and $m_{n}\rightarrow\infty$ as $n\rightarrow\infty$. Since $\bar{\Gamma}_{r_{k}\mathds{1}}^{m_{n}}(s^{k+1})\stackrel{{\scriptstyle\star}}{{\rightharpoonup}}\sigma(r_{k})$, then also $\sigma(\rho_{n})\stackrel{{\scriptstyle\star}}{{\rightharpoonup}}\sigma(r_{k})$. ∎
If we additionally assume that the fixed point $\sigma^{*}(r)$ is semi-attractive from above, i.e. that the convergence $\bar{\Gamma}_{r\mathds{1}}^{n}(s)\rightarrow\sigma^{*}(r)$ for all $s\geq\sigma^{*}(r)$ holds in the norm-topology, then $\sigma$ is obviously continuous.
We have thus proved the following theorem.
Theorem 3.4.
Let $\Gamma$ be a gain operator with the following properties:
  1. (i)
    $\Gamma$ satisfies the $\oplus$-MBI property.
  2. (ii)
    The fixed point $\sigma^{*}(r)$ of $\bar{\Gamma}_{r\mathds{1}}$ is semi-attractive from above for all $r\geq 0$.
Then there exists a $C^{0}$-path of decay for $\Gamma$.
Remark 3.11.
Observe that assumptions (a) and (b) together imply that $\Sigma(\Gamma)$ is UGAS. Indeed, (a) implies UGS and (b) implies GATT by letting $r=0$.
From the construction of $\sigma$, we obtain a new characterization of the $\oplus$-MBI property.
Theorem 3.5.
For a gain operator $\Gamma$, the following statements are equivalent:
  1. (a)
    There exists $\rho\in\mathcal{K}_{\infty}$ such that $\Gamma_{\rho}$ satisfies the $\oplus$-MBI property.
  2. (b)
    There exist $\rho\in\mathcal{K}_{\infty}$ and a path $\sigma:\mathbb{R}_{+}\rightarrow\ell^{\infty}_{+}(\mathcal{I})$ satisfying the properties listed in Proposition 3.5 for the gain operator $\Gamma_{\rho}$ (in place of $\Gamma$).
Proof.
We already know that (a) implies (b). To prove the converse, we can essentially use the same arguments as in the proof of Theorem 3.1. Let $s>0$ (without loss of generality) and $r\geq 0$ be chosen such that $s\leq r\mathds{1}\oplus\Gamma_{\rho}(s)$ (for a $\rho$ smaller than that in (b)). We define
[…]
$s\leq\sigma(r^{\prime})-\varepsilon\mathds{1}\leq\sigma(r_{n}),$
which contradicts the minimality of $r^{\prime}$. Hence, for every $\varepsilon>0$ there exists an index $i\in\mathcal{I}$ with $\sigma_{i}(r^{\prime})-s_{i}<\varepsilon$. The rest of the proof is identical to that of statement (b) in Theorem 3.1. ∎
We immediately obtain the following characterization in the finite case.
Theorem 3.6.
In the finite case, the following statements are equivalent:
  1. (a)
    There exists $\rho\in\mathcal{K}_{\infty}$ such that $\Gamma_{\rho}$ satisfies the $\oplus$-MBI property.
  2. (b)
    There exists a path of strict decay for $\Gamma$.
Let us now assume that $\mathcal{I}$ is infinite and $\mathcal{J}\subset\mathcal{I}$ is a nonempty finite subset. This subset corresponds to a finite sub-network with associated gain operator
$\Gamma^{\langle\mathcal{J}\rangle}:\ell^{\infty}_{+}(\mathcal{J})\rightarrow\ell^{\infty}_{+}(\mathcal{J}),\quad\Gamma^{\langle\mathcal{J}\rangle}_{i}(s):=\mu_{i}([\gamma_{ij}(s_{j})]_{j\in\mathcal{I}_{i}\cap\mathcal{J}}),$
where we use the restrictions of the MAFs $\mu_{i}$ to $\ell^{\infty}_{+}(\mathcal{I},\mathcal{J})=\ell^{\infty}_{+}(\mathcal{J})$, which trivially satisfy (M1)–(M4). The associated graph $\mathcal{G}^{\mathcal{J}}$ has vertex set $V(\mathcal{G}^{\mathcal{J}})=\mathcal{J}$ and edge set $E(\mathcal{G}^{\mathcal{J}})=\{ji:i\in\mathcal{J},\ j\in\mathcal{I}_{i}\cap\mathcal{J}\}$.
Now, assume that $\Gamma$ admits a path $\sigma:\mathbb{R}_{+}\rightarrow\ell^{\infty}_{+}(\mathcal{I})$ satisfying all properties listed in Proposition 3.5. Then, we can consider the restricted path
$\sigma^{\mathcal{J}}(r):=(\sigma_{i}(r))_{i\in\mathcal{J}},\quad\sigma^{\mathcal{J}}:\mathbb{R}_{+}\rightarrow\ell^{\infty}_{+}(\mathcal{J}).$
Since $\mathcal{J}$ is finite, this path is continuous, and moreover, it is increasing and inherits the coercivity estimates of $\sigma$. Finally, for every $i\in\mathcal{J}$ we have
$\Gamma^{\langle\mathcal{J}\rangle}_{i}(\sigma^{\mathcal{J}}(r))$ $=\mu_{i}([\gamma_{ij}(\sigma_{j}(r))]_{j\in\mathcal{I}_{i}\cap\mathcal{J}})\leq\mu_{i}([\gamma_{ij}(\sigma_{j}(r))]_{j\in\mathcal{I}_{i}})$
$=\Gamma_{i}(\sigma(r))\leq(\mathrm{id}+\rho)^{-1}(\sigma_{i}(r)).$
This implies $\Gamma^{\langle\mathcal{J}\rangle}_{\rho}\circ\sigma^{\mathcal{J}}\leq\sigma^{\mathcal{J}}$, hence $\sigma^{\mathcal{J}}$ is a path of strict decay for $\Gamma^{\langle\mathcal{J}\rangle}$. We thus have the following result.
Corollary 3.2.
Assume that $\Gamma_{\rho}$ satisfies the $\oplus$-MBI property for some $\rho\in\mathcal{K}_{\infty}$. Then the gain operator of every finite sub-network admits a path of strict decay.
4 Particular classes of gain operators
In this section, we apply our results and techniques to three particular classes of gain operators.
4.1 Max-type operators
Recall that a gain operator $\Gamma$ is called a max-type operator if all MAFs $\mu_{i}$ are of the form $\mu_{i}(s)=\sup_{j\in\mathcal{I}}s_{j}=\|s\|$. In this case, $\Gamma$ satisfies
$\Gamma(s^{1}\oplus s^{2})=\Gamma(s^{1})\oplus\Gamma(s^{2})\mbox{\quad for all\ }s^{1},s^{2}\in\ell^{\infty}_{+}(\mathcal{I}).$
An operator with this property is also called max-preserving. Gain operators of this type are particularly well-behaved. Using our general results, it is easy to provide a full characterization of the existence of a path of strict decay.
Theorem 4.1.
For a max-type gain operator $\Gamma$, the following statements are equivalent:
  1. (a)
    There exists a path of strict decay for $\Gamma$.
  2. (b)
    There exists $\rho\in\mathcal{K}_{\infty}$ such that $\Sigma(\Gamma_{\rho})$ is UGAS.
If Assumption 2.1 holds and there exists $M>0$ with $|\mathcal{I}_{i}|\leq M$ for all $i\in\mathcal{I}$, then another equivalent statement is the following:
  1. (c)
    There exists $\rho\in\mathcal{K}_{\infty}$ such that the trajectories of $\Sigma(\Gamma_{\rho})$ are norm-bounded and $\Gamma_{\rho}$ satisfies the uniform NJI condition.
Proof.
(a) $\Rightarrow$ (b): We only have to apply Theorem 3.1.
(b) $\Rightarrow$ (a): We first prove that $\Gamma_{\rho}$ satisfies the $\oplus$-MBI property. Hence, let $s\leq b\oplus\Gamma_{\rho}(s)=\bar{\Gamma}_{\rho,b}(s)$ for some $b,s\in\ell^{\infty}_{+}(\mathcal{I})$. This implies $s\leq\bar{\Gamma}_{\rho,b}^{n}(s)$ for all $n\in\mathbb{N}$. From the special form of the MAFs $\mu_{i}$, it follows that $\Gamma_{\rho}$ is also a max-type gain operator associated with the gains $(\mathrm{id}+\rho)\circ\gamma_{ij}$. From the fact that $\Gamma_{\rho}$ is max-preserving, we can easily derive that
$\bar{\Gamma}_{\rho,b}^{n}(s)=\Gamma_{\rho}^{n}(s)\oplus\bigoplus_{k=0}^{n-1}\Gamma^{k}_{\rho}(b)\mbox{\quad for all\ }n\geq 0.$ (12)
By assumption, $\Gamma_{\rho}^{n}(s)\rightarrow 0$. Hence, $s\leq\bigoplus_{k=0}^{\infty}\Gamma^{k}_{\rho}(b)$, which implies
$\|s\|\leq\sup_{k\in\mathbb{N}_{0}}\|\Gamma^{k}_{\rho}(b)\|\leq\sup_{k\in\mathbb{N}_{0}}\beta(\|b\|,k)=\beta(\|b\|,0)=:\varphi(\|b\|).$
Thus, $\Gamma_{\rho}$ satisfies the $\oplus$-MBI property. Next, we consider the operator $\bar{\Gamma}_{\rho,r\mathds{1}}$. Putting $b=r\mathds{1}$ in (12) implies that $\bar{\Gamma}_{\rho,r\mathds{1}}$ only has one fixed point, namely $\bigoplus_{k}\Gamma_{\rho}^{k}(r\mathds{1})$, which is globally attractive, since $\Gamma_{\rho}^{n}(s)$ converges to zero. The proof can then be concluded by an application of Theorem 3.2.
(b) $\Rightarrow$ (c): This follows from Proposition 2.8.
(c) $\Rightarrow$ (b): By Corollary 2.1, it suffices to prove that $\Psi(\Gamma_{\rho})$ is cofinal. From (12) and the norm-boundedness of the $\Sigma(\Gamma_{\rho})$-trajectories, it follows that also the trajectories of $\Sigma(\bar{\Gamma}_{\rho,b})$ are norm-bounded. Hence, the required cofinality follows from Proposition 2.10. ∎
Remark 4.1.
The theorem improves [13, Thm. VI.1], where uniform Lipschitz conditions on $\gamma_{ij}$ and $\gamma_{ij}^{-1}$ were required to obtain a path of strict decay. Moreover, [13, Thm. VI.1] only provided a sufficient condition and no relation to an NJI condition.
For max-type operators, we also have the following result characterizing the $\oplus$-MBI property, see [14, Prop. 3.12]. This result also follows easily from Proposition 3.1 and Corollary 3.1.
Proposition 4.1.
For a max-type gain operator $\Gamma$, the following statements are equivalent:
  1. (a)
    $\Gamma$ satisfies the $\oplus$-MBI property.
  2. (b)
[…]
  3. (c)
[…]
The condition that the trajectories of $\Sigma(\Gamma_{\rho})$ are norm-bounded in statement (c) of Theorem 4.1 can be written as
$\sup_{n\in\mathbb{N}_{0}}\sup_{j_{n}\cdots j_{2}j_{1}\in E^{n}(\mathcal{G})}\gamma_{j_{1}j_{2}}^{\rho}\circ\gamma_{j_{2}j_{3}}^{\rho}\circ\cdots\circ\gamma_{j_{n-1}j_{n}}^{\rho}(r)\leq M$
where $M=M(r)>0$ and $\gamma_{ij}^{\rho}:=(\mathrm{id}+\rho)\circ\gamma_{ij}$ for all $ji\in E(\mathcal{G})$. In the finite case, each sufficiently long chain of gains as above necessarily contains cycles and the number of gains not contained in a cycle can be effectively bounded. Moreover, the NJI condition implies that every cycle satisfies $\gamma_{k_{1}k_{2}}^{\rho}\circ\gamma_{k_{2}k_{3}}^{\rho}\circ\cdots\circ\gamma_{k_{l-1}k_{1}}^{\rho}<\mathrm{id}$. [^9] In the literature, this is expressed by saying that all cycles are contractions. Since this terminology is not compatible with the usual notion of a contraction, it should not be used in this context. This property easily yields the norm-boundedness of the $\Sigma(\Gamma_{\rho})$-trajectories, which we have to assume here, because in the infinite case there might be no cycles. In [13, Prop. VII.2], we have used a compactification strategy and the introduction of virtual gains “at infinity” to provide a sufficient condition for UGAS of $\Sigma(\Gamma_{\rho})$. This condition is somewhat similar to the cyclic condition (without really using cycles). However, in this paper we try to avoid compactification, since it introduces additional artefacts into the network.
4.2 Homogeneous, subadditive operators
We call a gain operator $\Gamma$ homogeneous if $\Gamma(\alpha s)=\alpha\Gamma(s)$ for all $s\in\ell^{\infty}_{+}(\mathcal{I})$ and $\alpha\geq 0$. We call $\Gamma$ subadditive if $\Gamma(s^{1}+s^{2})\leq\Gamma(s^{1})+\Gamma(s^{2})$ for all $s^{1},s^{2}\in\ell^{\infty}_{+}(\mathcal{I})$. In this subsection, we consider gain operators which are both homogeneous and subadditive. Examples are sum-type and max-type operators with linear interconnection gains. In the first case, $\Gamma$ is the restriction of a linear operator. In the second case, we speak of a max-linear operator. The next theorem is known, see [17, Prop. 8], but we give an alternative proof.
Theorem 4.2.
For a homogeneous and subadditive gain operator $\Gamma$, the following statements are equivalent:
  1. (a)
    There exists a path of strict decay for $\Gamma$.
  2. (b)
    The following “spectral radius” condition holds:
    $\inf_{n\in\mathbb{N}}\|\Gamma^{n}(\mathds{1})\|<1.$ (13)
  3. (c)
    The system $\Sigma(\Gamma)$ is UGAS.
If Assumption 2.1 holds and there exists $M>0$ with $|\mathcal{I}_{i}|\leq M$ for all $i\in\mathcal{I}$, then another equivalent statement is the following:
  1. (d)
    There exists $\rho\in\mathcal{K}_{\infty}$ such that the trajectories of $\Sigma(\hat{\Gamma}_{\rho})$ are norm-bounded and $\Gamma_{\rho}$ satisfies the uniform NJI condition.
Proof.
The equivalence of (b) and (c) was shown in [17, Prop. 8]. The same proposition also yields the equivalence to (a), but with different methods than the ones used in the present paper. So, we will give another proof. For the implication “(a) $\Rightarrow$ (c)” we can simply apply Theorem 3.1, so it remains to conclude (a) from (b) or (c).
First, we refer again to [17, Prop. 9], which also shows that (b) is equivalent to uniform global exponential stability (UGES), which is UGAS with a $\mathcal{K}\mathcal{L}$-function of the form $\beta(r,n)=Ma^{n}r$, where $M>0$ and $a\in(0,1)$. From the homogeneity of $\Gamma$, it follows that we can choose a linear $\mathcal{K}_{\infty}$-function $\rho$ (now regarded as a positive real number) such that $\Gamma_{\rho}=(1+\rho)\cdot\Gamma$ is still UGES. In particular, this implies that (b) holds with $\Gamma_{\rho}$ in place of $\Gamma$. We refer to [14, Prop. 3.7], which shows that this, in turn, implies that $\Sigma(\hat{\Gamma}_{\rho})$ is UGS. We claim that for any $r\geq 0$, points $s^{1}\leq s^{2}$ and $n\in\mathbb{N}_{0}$, we have
$\bar{\Gamma}^{n}_{r\mathds{1}}(s^{2})-\bar{\Gamma}^{n}_{r\mathds{1}}(s^{1})\leq\Gamma^{n}(s^{2}-s^{1}).$ (14)
For $n=0$, this trivially holds. Assuming that it holds for a fixed $n$, with Lemma A.5 we obtain
$\bar{\Gamma}^{n+1}_{r\mathds{1}}(s^{2})-\bar{\Gamma}^{n+1}_{r\mathds{1}}(s^{1})$ $=r\mathds{1}\oplus\Gamma(\bar{\Gamma}_{r\mathds{1}}^{n}(s^{2}))-r\mathds{1}\oplus\Gamma(\bar{\Gamma}_{r\mathds{1}}^{n}(s^{1}))$
$\leq\Gamma(\bar{\Gamma}_{r\mathds{1}}^{n}(s^{2}))-\Gamma(\bar{\Gamma}_{r\mathds{1}}^{n}(s^{1}))\leq\Gamma(\bar{\Gamma}_{r\mathds{1}}^{n}(s^{2})-\bar{\Gamma}_{r\mathds{1}}^{n}(s^{1}))$
$\leq\Gamma(\Gamma^{n}(s^{2}-s^{1}))=\Gamma^{n+1}(s^{2}-s^{1}).$
Here, the second inequality follows from subadditivity of $\Gamma$ and the last one from the induction hypothesis. The same estimates hold with $\Gamma_{\rho}$ in place of $\Gamma$, in case that $\rho$ is linear. Hence,
$\|\bar{\Gamma}^{n}_{\rho,r\mathds{1}}(s^{2})-\bar{\Gamma}^{n}_{\rho,r\mathds{1}}(s^{1})\|\leq Ma^{n}\|s^{2}-s^{1}\|$
for appropriate constants $M>0$ and $a\in(0,1)$. This implies that $\bar{\Gamma}_{\rho,r\mathds{1}}$ can only have one fixed point and every trajectory converges to this fixed point at exponential speed. Hence, Theorem 3.2 yields a path of strict decay for $\Gamma$.
Finally, for the equivalence of the first three statements to statement (d), we use Corollary 2.3 together with Proposition 2.11, Proposition 2.10 and the fact that UGAS implies UGES in our special case. ∎
Remark 4.2.
Under the assumption that $\Sigma(\Gamma)$ is UGAS, the proof of the theorem shows that the candidate path $\sigma_{*,\rho}$ is a $C^{0}$-path of strict decay for some linear $\rho$. We can easily see that this path is linear, and hence a path of strict decay (i.e. we do not need a piecewise linear approximation). Indeed, we can multiply both sides of the identity $\sigma_{*,\rho}(r)=r\mathds{1}\oplus\Gamma_{\rho}(\sigma_{*,\rho}(r))$ with a scalar $\alpha\geq 0$ and obtain $\alpha\sigma_{*,\rho}(r)=(\alpha r)\mathds{1}\oplus\Gamma_{\rho}(\alpha\sigma_{*,\rho}(r))$, using the homogeneity of $\Gamma_{\rho}$. This implies $\sigma_{*,\rho}(r)=r\sigma_{*,\rho}(1)$ for all $r\geq 0$.
We further have the following result.
Proposition 4.2.
For a homogeneous and subadditive gain operator $\Gamma$, consider the following statements:
  1. (a)
    The system $\Sigma(\Gamma)$ is UGAS.
  2. (b)
    $\Gamma$ satisfies the $\oplus$-MBI property.
  3. (c)
[…]
Then (a) $\Rightarrow$ (b) $\Leftrightarrow$ (c).
Proof.
[…]
We finally prove the following theorem, which is somewhat surprising, since it shows that the $\oplus$-MBI property is sufficient to obtain a path of strict decay for homogeneous subadditive operators.
Theorem 4.3.
Let $\Gamma$ be a homogeneous and subadditive gain operator and assume that $\Gamma_{\rho}$ satisfies the $\oplus$-MBI property with a linear $\rho\in\mathcal{K}_{\infty}$. Then there exists a path of strict decay for $\Gamma$.
Proof.
First, observe that $\Gamma_{\rho}$ is a subadditive and homogeneous gain operator if $\rho$ is linear. As we have shown in the proof of Theorem 4.2,
$\bar{\Gamma}^{n}_{\rho,r\mathds{1}}(s^{2})-\bar{\Gamma}^{n}_{\rho,r\mathds{1}}(s^{1})\leq\Gamma_{\rho}^{n}(s^{2}-s^{1})$
holds, whenever $s^{1}\leq s^{2}$, $r\geq 0$ and $n\in\mathbb{N}_{0}$. This implies that $\bar{\Gamma}_{\rho,r\mathds{1}}$ has a unique fixed point. We consider the candidate path $\sigma_{*}$, which maps $r$ to the unique fixed point of $\bar{\Gamma}_{\rho,r\mathds{1}}$ for each $r\geq 0$. Then, we have the identity
$\sigma_{*}(r)\equiv r\mathds{1}\oplus\Gamma_{\rho}(\sigma_{*}(r)).$
Multiplying both sides by a nonnegative scalar $\alpha$ yields
$\alpha\sigma_{*}(r)\equiv(\alpha r)\mathds{1}\oplus\Gamma_{\rho}(\alpha\sigma_{*}(r)).$
By uniqueness of the fixed point, this implies $\sigma_{*}(\alpha r)=\alpha\sigma_{*}(r)$. We thus have $\sigma_{*}(r)\equiv r\sigma_{*}(1)$, showing that $\sigma_{*}$ is linear, and hence continuous. Consequently, $\sigma_{*}$ is a path of strict decay. ∎
Remark 4.3.
We note that several further characterizations of the equivalent conditions in Theorem 4.2 are available, especially for linear operators, see [10]. This includes, in particular, the spectral radius condition $r(\Gamma)<1$, which is an immediate consequence of (13) and Gelfand’s formula.
4.3 Finite-dimensional operators
In this subsection, we study the finite case, where $\ell^{\infty}(\mathcal{I})$ can be identified with the Euclidean space $\mathbb{R}^{N}$, $N:=|\mathcal{I}|$, and $\ell^{\infty}_{+}(\mathcal{I})$ with $\mathbb{R}^{N}_{+}:=\{s\in\mathbb{R}^{N}:s_{i}\geq 0,\ 1\leq i\leq N\}$. In particular, we clarify how our results are related to previous results from the literature.
First, let us clarify which assumptions we have to impose on the functions $\gamma_{ij}$ and $\mu_{i}$ to obtain a gain operator in the sense of our definition.
Proposition 4.3.
Let $\mathcal{G}=(V(\mathcal{G}),E(\mathcal{G}))$ be a finite directed graph without self-loops (i.e. edges from a vertex to itself) and put $\mathcal{I}:=V(\mathcal{G})$. Consider families $\{\gamma_{ij}:ji\in E(\mathcal{G})\}$ and $\{\mu_{i}:1\leq i\leq|\mathcal{I}|\}$ with $\gamma_{ij}\in\mathcal{K}_{\infty}$ and $\mu_{i}:\ell^{\infty}_{+}(\mathcal{I})\rightarrow\mathbb{R}_{+}$ satisfying the following properties for all $i\in\mathcal{I}$:
  1. (i)
    $\mu_{i}(0)=0$ and $\mu_{i}(s)>0$ if $s>0$.
  2. (ii)
    $0\leq s^{1}\leq s^{2}$ implies $\mu_{i}(s^{1})\leq\mu_{i}(s^{2})$.
  3. (iii)
    $\|s\|\rightarrow\infty$ implies $\mu_{i}(s)\rightarrow\infty$.
  4. (iv)
    $\mu_{i}$ is continuous.
Then the operator $\Gamma:\ell^{\infty}_{+}(\mathcal{I})\rightarrow\ell^{\infty}_{+}(\mathcal{I})$, defined by $\Gamma_{i}(s):=\mu_{i}([\gamma_{ij}(s_{j})]_{j\in\mathcal{I}_{i}})$ for all $i\in\mathcal{I}$, is a gain operator in the sense of our definition.
Proof.
Since $\{\gamma_{ij}\}$ is a finite family of functions, it is pointwise equicontinuous. It remains to verify (M1)–(M4) for $\{\mu_{i}\}$. This, in turn, reduces to showing that $\mu_{i}(s)\geq\xi(\|s\|)$ for some $\xi\in\mathcal{K}_{\infty}$, since the rest trivially follows from our assumptions. To this end, we define
$\tilde{\xi}(r):=\min_{(i,j)\in\mathcal{I}\times\mathcal{I}}\mu_{i}(re^{j}).$
From assumption (i), we obtain $\tilde{\xi}(0)=0$. Clearly, $\tilde{\xi}(r)\geq 0$ for all $r\in\mathbb{R}_{+}$. From assumption (iv), it follows, moreover, that $\tilde{\xi}$ is continuous. If $r_{1}\leq r_{2}$, then $r_{1}e^{j}\leq r_{2}e^{j}$ for each $j$, and hence assumption (ii) implies $\tilde{\xi}(r_{1})\leq\tilde{\xi}(r_{2})$, showing that $\tilde{\xi}$ is non-decreasing. If $r\rightarrow\infty$, then $\|re^{j}\|\rightarrow\infty$, and hence, following from assumption (iii), $\tilde{\xi}(r)\rightarrow\infty$. Finally, assumption (i) implies $\tilde{\xi}(r)>0$ if $r>0$. By Lemma A.7, we can then bound $\tilde{\xi}$ from below by a $\mathcal{K}_{\infty}$-function $\xi$. To show that $\mu_{i}(s)\geq\xi(\|s\|)$, pick $s\in\ell^{\infty}_{+}(\mathcal{I})$ and choose an index $j$ such that $s_{j}=\|s\|$. Then $s\geq s_{j}e^{j}=\|s\|e^{j}$, implying $\mu_{i}(s)\geq\mu_{i}(\|s\|e^{j})\geq\tilde{\xi}(\|s\|)\geq\xi(\|s\|)$. This concludes the proof. ∎
Remark 4.4.
The assumptions of the proposition can also be found in the literature on finite-dimensional gain operators. It is furthermore clear that in finite dimensions there always exists a uniform lower bound $\eta\in\mathcal{K}_{\infty}$ on the gains $\gamma_{ij}$ as required in Assumption 2.1.
4.3.1 The irreducible case
We first consider the case in which the interconnection graph $\mathcal{G}$ is strongly connected. That is, for each pair $(i,j)\in\mathcal{I}\times\mathcal{I}$, there exists a directed path from $i$ to $j$. This case is much simpler than the general one. We first make the following observation.
Lemma 4.1.
If $\mathcal{G}$ is strongly connected, then $\Psi(\Gamma)$ is coercive.
Proof.
By monotonicity, the inequality $\Gamma(s)\leq s$ implies $\Gamma^{n}(s)\leq s$ for all $n\in\mathbb{N}_{0}$. Since $\mathcal{G}$ is strongly connected, for each index pair $(i,j)$ we find an $n$ such that $j\in\mathcal{N}^{-}_{i}(n)$. Moreover, since there are only finitely many indices, $n$ can be chosen independently of $(i,j)$. This implies $s_{i}\geq\Gamma_{i}^{l}(s)\geq(\xi\circ\eta)^{l}(s_{j})$ for some $l\in\{0,1,\ldots,n\}$. If we choose $j$ such that $s_{j}=\|s\|$, then
$s_{i}\geq\min_{0\leq l\leq n}(\xi\circ\eta)^{l}(\|s\|)=(\xi\circ\eta)^{n}(\|s\|).$
This holds for all $i$, and thus the assertion holds with $\underline{\varphi}:=(\xi\circ\eta)^{n}$. ∎
The next proposition shows that most properties that we have investigated are equivalent in the considered scenario.
Proposition 4.4.
If $|\mathcal{I}|<\infty$ and $\mathcal{G}$ is strongly connected, the following statements are equivalent:
  1. (a)
    $\Gamma$ satisfies the NJI condition.
  2. (b)
    $\Gamma$ satisfies the uniform NJI condition.
  3. (c)
    $\Gamma$ satisfies the $\oplus$-MBI property.
  4. (d)
    $\Sigma(\Gamma)$ is UGAS.
  5. (e)
[…]
Proof.
(a) $\Rightarrow$ (c): Consider arbitrary $s\in\ell^{\infty}_{+}(\mathcal{I})$ and $r>0$ satisfying $s\leq\bar{\Gamma}_{r\mathds{1}}(s)$. By monotonicity, the trajectory $s^{n}:=\bar{\Gamma}_{r\mathds{1}}^{n}(s)$, $n\in\mathbb{N}_{0}$, is increasing, and thus $s^{n}>0$ for all $n$. Hence, by the NJI condition, for every $n$ there exists $i=i(n)\in\mathcal{I}$ with $\Gamma_{i}(s^{n})<s^{n}_{i}$. Since $\mathbb{N}_{0}$ is infinite and $\mathcal{I}$ is finite, there must be an $i^{*}\in\mathcal{I}$ with $\Gamma_{i^{*}}(s^{n})<s^{n}_{i^{*}}$ for infinitely many $n\in\mathbb{N}_{0}$. For all such $n$, we obtain
$s^{n}_{i^{*}}\leq s^{n+1}_{i^{*}}=\max\{r,\Gamma_{i^{*}}(s^{n})\}=r.$
The last identity holds, since otherwise we would obtain the contradiction $s^{n}_{i^{*}}<s^{n}_{i^{*}}$. Clearly, this yields $s^{n}_{i^{*}}=r$ for infinitely many $n$, and since the trajectory is increasing, $s^{n}_{i^{*}}=r$ for all $n\in\mathbb{N}$. Using uniform positive definiteness of MAFs, we find that for all $n\in\mathbb{N}$ and $j\in\mathcal{I}_{i^{*}}$ it holds that
$(\xi\circ\eta)(s^{n}_{j})\leq\Gamma_{i^{*}}(s^{n})\leq s^{n+1}_{i^{*}}=r.$
Using the same argument repetitively, we can show that $(\xi\circ\eta)^{l}(s^{n}_{j})\leq r$ whenever there exists a path from $j$ to $i^{*}$ of length $l$ in $\mathcal{G}$. Since we assume that $\mathcal{G}$ is strongly connected, there exists such a path for every $j$ and its length can be bounded by some $m\in\mathbb{N}$ due to the finite number of vertices. This implies
$\|s\|\leq\max_{0\leq l\leq m}(\xi\circ\eta)^{-l}(r).$
We have thus proven that $\Gamma$ satisfies the $\oplus$-MBI property.
[…]
(d) $\Rightarrow$ (b): Apply Proposition 2.8.
(b) $\Rightarrow$ (a): This is trivial.
(c) $\Rightarrow$ (e): Apply Proposition 3.1.
(e) $\Rightarrow$ (a): This is trivial. ∎
From Theorem 3.1, we now know that the existence of a path of strict decay implies all the properties listed in the preceding proposition. For the converse, observe that UGAS implies that assumptions (i) and (ii) in Theorem 3.3 with $A=\Psi(\Gamma)$ hold. Moreover, assumption (iii) follows from the strong connectivity of $\mathcal{G}$. Hence, we have the following theorem.
Theorem 4.4.
If $|\mathcal{I}|<\infty$ and $\mathcal{G}$ is strongly connected, the following statements are equivalent:
  1. (a)
    There exists a path of strict decay for $\Gamma$.
  2. (b)
    There exists $\rho\in\mathcal{K}_{\infty}$ such that $\Gamma_{\rho}$ satisfies the NJI condition.
Alternatively, we could have used Theorem 3.6 which shows that the existence of a path of strict decay is equivalent to the $\oplus$-MBI property of $\Gamma_{\rho}$.
Remark 4.5.
Let us compare the theorem to the state-of-the-art for finite networks. In [20, Thm. 5.5], we find the result that there exists a piecewise linear path $\sigma:\mathbb{R}_{+}\rightarrow\ell^{\infty}_{+}(\mathcal{I})$ satisfying $\sigma_{i}\in\mathcal{K}_{\infty}$ for all $i$ and $\Gamma(\sigma(r))<\sigma(r)$ for all $r>0$ under the following assumptions:
  • $\mathcal{G}$ is strongly connected.
  • The MAFs satisfy
    • $s^{1}\ll s^{2}$ $\Rightarrow$ $\mu_{i}(s^{1})<\mu_{i}(s^{2})$.
    • $\|s\|\rightarrow\infty$ implies $\mu_{i}(s)\rightarrow\infty$.
  • $\Gamma$ satisfies the NJI condition.
These assumptions are partially weaker, partially stronger than ours. The conclusion is weaker, since there might be no $\rho\in\mathcal{K}_{\infty}$ with $\Gamma_{\rho}\circ\sigma\leq\sigma$.
4.3.2 The general case
For the general finite case, without the assumption that $\mathcal{G}$ is strongly connected, we need an additional uniform continuity assumption for the MAFs that we have not used before. We can formulate this assumption in terms of a comparison function as follows: There exists a $\mathcal{K}_{\infty}$-function $\beta$ such that $0\leq s^{1}\leq s^{2}$ implies
$\mu_{i}(s^{2})-\mu_{i}(s^{1})\leq\beta(\|s^{2}-s^{1}\|)\mbox{\quad for all\ }i\in\mathcal{I}.$ (15)
Observe that this is indeed equivalent to the uniform continuity of each $\mu_{i}$. First, since there are only finitely many MAFs, the same $\beta$ can be used for all $i\in\mathcal{I}$. Second, if $s^{1}$ and $s^{2}$ are arbitrary elements of $\ell^{\infty}_{+}(\mathcal{I})$ (not necessarily ordered), it follows from the assumption that
$\mu_{i}(s^{2})-\mu_{i}(s^{1})\leq\mu_{i}(s^{1}\oplus s^{2})-\mu_{i}(s^{1})\leq\beta(\|s^{1}\oplus s^{2}-s^{1}\|)\leq\beta(\|s^{2}-s^{1}\|)$
and analogously for $\mu_{i}(s^{1})-\mu_{i}(s^{2})$. Hence, $|\mu_{i}(s^{2})-\mu_{i}(s^{1})|\leq\beta(\|s^{2}-s^{1}\|)$, which is the characterization of uniform continuity in terms of comparison functions.
In the literature on finite-dimensional small-gain theory (see [8, 9, 20, 21]), instead the assumption of subadditivity is used, which is a special case of the above. To see this, note that a MAF $\mu$ is subadditive if $\mu(s^{1}+s^{2})\leq\mu(s^{1})+\mu(s^{2})$ for all $s^{1},s^{2}\in\ell^{\infty}_{+}(\mathcal{I})$. If $0\leq s^{1}\leq s^{2}$, this implies
$\mu(s^{2})-\mu(s^{1})=\mu(s^{1}+(s^{2}-s^{1}))-\mu(s^{1})\leq\mu(s^{2}-s^{1})\leq\mu(\|s^{2}-s^{1}\|\mathds{1})$
and we can define $\beta(r):=\mu(r\mathds{1})$, which can be bounded from above by a $\mathcal{K}_{\infty}$-function.
Theorem 4.5.
Consider the finite case and assume that the MAFs $\mu_{i}$ are uniformly continuous. Then the following statements are equivalent:
  1. (a)
    There exists $\rho\in\mathcal{K}_{\infty}$ such that $\Gamma_{\rho}$ satisfies the NJI condition.
  2. (b)
    There exists $\rho\in\mathcal{K}_{\infty}$ such that $\Gamma_{\rho}$ satisfies the uniform NJI condition.
  3. (c)
    There exists $\rho\in\mathcal{K}_{\infty}$ such that $\Gamma_{\rho}$ satisfies the $\oplus$-MBI property.
  4. (d)
    There exists a path of strict decay for $\Gamma$.
  5. (e)
    There exists $\rho\in\mathcal{K}_{\infty}$ such that $\Sigma(\Gamma_{\rho})$ is UGAS.
Proof.
The equivalences (a) $\Leftrightarrow$ (b) and (c) $\Leftrightarrow$ (d) were proved in Proposition 2.7 and Theorem 3.6, respectively. The implication (d) $\Rightarrow$ (e) follows from Theorem 3.1 and (e) $\Rightarrow$ (a) is trivial.
It remains to prove that (a) $\Rightarrow$ (c). To this end, consider $s\in\ell^{\infty}_{+}(\mathcal{I})$ and $r\geq 0$ with $s\leq r\mathds{1}\oplus\Gamma(s)$. Without loss of generality, we can assume that $s>0$. We split the index set $\mathcal{I}$ into the two subsets
$\mathcal{I}^{1}:=\left\{i\in\mathcal{I}:s_{i}\leq r\right\},\quad\mathcal{I}^{2}:=\mathcal{I}\backslash\mathcal{I}^{1}.$
By assumption, there exists at least one index $i$ with $\Gamma_{i}(s)<s_{i}$, which implies $s_{i}\leq r$. Hence, $\mathcal{I}^{1}$ is nonempty. We define the $\mathcal{K}_{\infty}$-function
$\varphi_{1}:=(\rho\circ\xi\circ\eta)^{-1}\circ\beta\circ\left(\max_{ji\in E(\mathcal{G})}\gamma_{ij}\right),$
where $\rho$ comes from (a), $\xi$ comes from (M1), $\eta\leq\gamma_{ij}$ for all $ji\in E(\mathcal{G})$ with $\eta\in\mathcal{K}_{\infty}$, and $\beta$ comes from (15). Without loss of generality, we can assume that
$\varphi_{1}(r)>r\mbox{\quad and \quad}(\rho\circ\xi\circ\eta)^{-1}(r)>r\mbox{\quad for all\ }r>0.$ (16)
Let us assume to the contrary that $s_{|\mathcal{I}^{2}}\geq\varphi_{1}(r)\mathds{1}$. We want to show that this implies
[…]
which in turn implies $\Gamma_{\rho}(s_{|\mathcal{I}^{2}})\geq s_{|\mathcal{I}^{2}}$ in contradiction to (a). We fix $i\in\mathcal{I}^{2}$ and define
$a_{i}:=[\gamma_{ij}(s_{j})]_{j\in\mathcal{I}_{i}\cap\mathcal{I}^{2}}.$
We can assume without loss of generality that $a_{i}\neq 0$. Indeed, $a_{i}=0$ is only possible if $\mathcal{I}_{i}\cap\mathcal{I}^{2}=\emptyset$, implying
[…]
in contradiction to our assumption hat $s_{|\mathcal{I}^{2}}\geq\varphi_{1}(r)\mathds{1}$. The inequality to be shown can then be written as
$\rho(\mu_{i}(a_{i}))\geq\mu_{i}([\gamma_{ij}(r)]_{j\in\mathcal{I}_{i}\cap\mathcal{I}^{1}}+a_{i})-\mu_{i}(a_{i}).$
Using (15), we see that a sufficient condition for this inequality to hold is
$\rho(\mu_{i}(a_{i}))\geq\beta(\|[\gamma_{ij}(r)]_{j\in\mathcal{I}_{i}\cap\mathcal{I}^{1}}\|).$
This last inequality follows from the definition of $\varphi_{1}$, since
$\rho(\mu_{i}(a_{i}))$ $\geq\rho(\xi(\|a_{i}\|))\geq\rho\circ\xi\circ\eta\circ\varphi_{1}(r)$
$=\beta\circ\left(\max_{ji\in E(\mathcal{G})}\gamma_{ij}\right)(r)\geq\beta(\|[\gamma_{ij}(r)]_{j\in\mathcal{I}_{i}\cap\mathcal{I}^{1}}\|).$
Hence, there exists at least one index $i\in\mathcal{I}^{2}$ such that $s_{i}<\varphi_{1}(r)$. We can split the index set $\mathcal{I}$ again into
$\mathcal{I}^{1}:=\left\{i\in\mathcal{I}:s_{i}\leq\max\{r,\varphi_{1}(r)\}\right\},\quad\mathcal{I}^{2}:=\mathcal{I}\backslash\mathcal{I}^{1},$
observing that the new $\mathcal{I}^{1}$ contains at least one more element than the old one. With the same arguments as before, we can define another $\mathcal{K}_{\infty}$-function $\varphi_{2}$ such that $s_{i}<\varphi_{2}(r)$ must hold for at least one $i\in\mathcal{I}^{2}$. In this way, we can continue inductively until $\mathcal{I}^{2}=\emptyset$ (using that there are only finitely many indices). We thus find a $\mathcal{K}_{\infty}$-function $\varphi$ such that $\|s\|\leq\varphi(r)$. Note that we have proved that $\Gamma$ satisfies the $\oplus$-MBI property (and not $\Gamma_{\rho}$). By choosing a uniformly continuous $\mathcal{K}_{\infty}$-function $\rho^{\prime}<\rho$, the proof also works for $\Gamma_{\rho^{\prime}}$ in place of $\Gamma$. ∎
Remark 4.6.
In the literature, we find a very similar result, see [18, Thm. C.41]. However, here the $\oplus$-MBI property does not appear, but instead the stronger MBI property, which requires that $s\leq b+\Gamma(s)$ implies $\|s\|\leq\varphi(\|b\|)$. The proof of this result uses different methods, however, and in particular uses the subadditivity of MAFs for the path construction.
5 Summary and open questions
In this paper, we introduced gain operators for countable (finite or infinite) networks of systems, where each subsystem is only influenced by finitely many other subsystems. In contrast to the existing theory for finite networks, our definition does not require the subadditivity of the monotone aggregation functions involved.
We also introduced the notion of a path of strict decay for a given gain operator, whose existence was used in [13] as a small-gain condition for constructing an ISS Lyapunov function from ISS Lyapunov functions of the subsystems. In Proposition 2.3, we showed that the existence of such a path is equivalent to the existence of a path $\sigma:\mathbb{R}_{+}\rightarrow\ell^{\infty}_{+}(\mathcal{I})$, which is continuous and increasing, satisfies coercivity estimates and evolves in the decay set of an enlarged gain operator $\Gamma_{\rho}=(\mathrm{id}+\rho)\circ\Gamma$. Such a path was called a $C^{0}$-path (of strict decay).
Subsequently, we investigated a number of properties that a gain operator or its associated dynamical system can have. The two main properties, both implied by the existence of $C^{0}$-path (see Theorem 3.1), were the uniform global asymptotic stability (UGAS) of $\Sigma(\Gamma_{\rho})$ and the $\oplus$-MBI property of $\Gamma_{\rho}$. The relation between these two properties could not be completely clarified. However, in the finite case they are equivalent under the additional assumption that the MAFs are uniformly continuous (see Theorem 4.5), and in the specific infinite cases, studied in Section 4, UGAS implies the $\oplus$-MBI property (see Proposition 4.1 and Proposition 4.2). Concerning the construction of a $C^{0}$-path, the $\oplus$-MBI property turned out to be more constitutive. In particular, the $\oplus$-MBI property implies the existence of a path satisfying all properties of a $C^{0}$-path except for norm-continuity from above, which is weakened to sequential weak continuity. In the finite case, this yields a characterization of the existence of a $C^{0}$-path in terms of the $\oplus$-MBI property, and in general, it implies that every finite sub-network of an infinite network admits a $C^{0}$-path (see Subsection 3.3.3).
It is tempting to conjecture that UGAS is the additional ingredient that is needed to show that the weak $C^{0}$-path obtained from the $\oplus$-MBI property is actually a $C^{0}$-path. But we were not able to prove this except in the cases of max-type and homogeneous, subadditive gain operators, respectively.
Concerning UGAS, we showed that it implies a uniform version of the no-joint-increase (NJI) condition (see Proposition 2.8), used extensively in the small-gain theory of finite networks. In the finite case, the uniform NJI condition reduces to the NJI condition (see Proposition 2.7), so it can be regarded as a proper extension of the latter to the infinite case. We also proved that the uniform NJI condition together with the cofinality of the decay set, both assumed for an enlarged gain operator, imply UGAS (see Proposition 2.9).
We also generalized some of the known results for finite networks. For instance, we proved that the existence of a path of strict decay in the finite case is equivalent to the $\oplus$-MBI property of an enlarged gain operator, which was not known before (see Theorem 3.6). We also gave a new explanation why the construction of a path of strict decay via linear interpolation between points of a complete orbit works in the finite irreducible case. The “magic” ingredient here is the coercivity of the decay set $\Psi(\Gamma)$, which we cannot expect in more general setups (see Lemma 4.1). Finally, we proved that the NJI condition implies the $\oplus$-MBI property under the assumption that the MAFs are uniformly continuous.
Many open problems remain for future work. In particular, the following topics are relevant:
  1. 1.
    The ultimate goal, to obtain a complete characterization of the existence of a path of strict decay in terms of testable conditions, was only obtained in special cases. One major step towards this goal would be to clarify when the weak-convergence of $\bar{\Gamma}_{r\mathds{1}}^{n}(s)$ to $\sigma^{*}(r)$ for $s\geq\sigma^{*}(r)$ can be strengthened to norm-convergence. In particular, it must be understood if this follows from some form of UGAS or if stronger conditions are necessary.
  2. 2.
    With regard to Corollary 2.3, it is important to understand under which conditions the trajectories of the augmented gain operator are norm-bounded. Can we characterize the norm-boundedness in terms of conditions that can be checked by computation?
  3. 3.
    An important underlying object of a gain operator, that we did not pay the appropriate attention to, is the interconnection graph $\mathcal{G}$. This is an infinite directed graph with finite indegrees. In many of our results, we make additional assumptions on this graph for purely technical reasons without paying attention to the question how these assumptions reduce the possible cases of application. So, this is certainly an important topic for future research.
  4. 4.
    A further topic for future research concerns the question of scalable ISS, see [4]. Here, we ask the question under which conditions stability measures are preserved when a network is enlarged. If we can cover a large-but-finite network that may grow in time with one infinite network which is ISS, we should be able to say that stability measures are preserved when further subsystems are added. The stability measures are related to the $\mathcal{K}_{\infty}$-functions appearing in the definition of a path of strict decay, so it is important to take a closer look at how these functions are related to the fundamental components of the network. A first step in this direction is Corollary 3.2.
  5. 5.
    Besides answering the open questions within the setup of this paper, one can ask if this setup can be generalized. For instance, one can ask if it makes sense to build a similar theory on $\ell^{p}$-spaces for finite $p$, if one can relax the assumption that the sets $\mathcal{I}_{i}$ all have to be finite, or even if the countability of the index set $\mathcal{I}$ can be dropped.
Appendix A Some technical lemmas
The following three lemmas describe properties of the weak-topology of $\ell^{\infty}(\mathcal{I})$.
Lemma A.1.
A sequence $(s^{n})_{n\in\mathbb{N}}$ in $\ell^{\infty}(\mathcal{I})$ converges in the weak-topology if and only if it is norm-bounded and componentwise convergent.
Proof.
As a general fact, the sequence $(s^{n})$ converges to some $s\in\ell^{\infty}(\mathcal{I})$ if and only if $s^{n}(t)$ converges to $s(t)$ for every $t\in\ell^{1}(\mathcal{I})$. By (1), this means
$\lim_{n\rightarrow\infty}\sum_{i\in\mathcal{I}}s^{n}_{i}t_{i}=\sum_{i\in\mathcal{I}}s_{i}t_{i}\mbox{\quad for all\ }t\in\ell^{1}(\mathcal{I}).$
If we let $t$ range through the standard basis of $\ell^{1}(\mathcal{I})$, consisting of the unit vectors $e^{i}$, $i\in\mathcal{I}$, this implies
$\lim_{n\rightarrow\infty}s^{n}_{i}=s_{i}\mbox{\quad for all\ }i\in\mathcal{I}.$
Hence, we have shown that weak convergence implies componentwise convergence. Now, we show that $(s^{n})$ is norm-bounded. By [15, Cor. 2.6.8], a set $A\subset\ell^{\infty}(\mathcal{I})$ is norm-bounded if and only if the set $\{s(t):s\in A\}$ is bounded in $\mathbb{R}$ for each $t\in\ell^{1}(\mathcal{I})$. Applying this to $A:=\{s^{n}:n\in\mathbb{N}\}$ shows norm-boundedness, since $s^{n}(t)$ is convergent (and thus bounded) for each $t$ by assumption.
Conversely, let $(s^{n})_{n\in\mathbb{N}}$ be a sequence in $\ell^{\infty}(\mathcal{I})$ which is norm-bounded, say $\|s^{n}\|\leq B$ for all $n$, and converges to some $s\in\ell^{\infty}(\mathcal{I})$ componentwise. Pick $t\in\ell^{1}(\mathcal{I})$ and $\varepsilon>0$. Obviously, we can assume that $t\neq 0$. Choose a subset $\mathcal{J}\subset\mathcal{I}$ such that $\mathcal{I}\backslash\mathcal{J}$ is finite and
$\sum_{i\in\mathcal{J}}|t_{i}|<\frac{\varepsilon}{4B}.$
The existence of such $\mathcal{J}$ immediately follows from $t\in\ell^{1}(\mathcal{I})$. Now, we can choose $N$ large enough so that $|s^{n}_{i}-s_{i}|<\varepsilon/(2\|t\|_{1})$ for $i\in\mathcal{I}\backslash\mathcal{J}$ and $n\geq N$. The choice of $\mathcal{J}$ and $N$ then implies that for all $n\geq N$
$\Bigl|\sum_{i\in\mathcal{I}}(s^{n}_{i}-s_{i})t_{i}\Bigr|$ $\leq\sum_{i\in\mathcal{I}}|s^{n}_{i}-s_{i}|\cdot|t_{i}|$
$=\sum_{i\in\mathcal{I}\backslash\mathcal{J}}|s^{n}_{i}-s_{i}|\cdot|t_{i}|+\sum_{i\in\mathcal{J}}|s^{n}_{i}-s_{i}|\cdot|t_{i}|$
$\leq\frac{\varepsilon}{2\|t\|_{1}}\sum_{i\in\mathcal{I}}|t_{i}|+\sum_{i\in\mathcal{J}}(|s^{n}_{i}|+|s_{i}|)|t_{i}|\leq\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon.$
This shows that $(s^{n})$ converges to $s$ in the weak-topology. ∎
Lemma A.2.
Every norm-bounded sequence $(s^{n})_{n\in\mathbb{N}}$ in $\ell^{\infty}(\mathcal{I})$ has a subsequence which is componentwise convergent.
Proof.
By the theorem of Banach-Alaoglu and the assumption of norm-boundedness, $(s^{n})_{n\in\mathbb{N}}$ is contained in a weakly compact subset of $\ell^{\infty}(\mathcal{I})$. By [15, Cor. 2.6.20], this subset is metrizable. Hence, its compactness is equivalent to its sequential compactness. The rest follows from Lemma A.1. ∎
Lemma A.3.
In $\ell^{\infty}(\mathcal{I})$, every order interval of the form $[a,b]$ with $a\leq b$ is weakly sequentially compact.
Proof.
Let $(s^{n})$ be a sequence in $[a,b]$. Then clearly, $(s^{n})$ is norm-bounded. By Lemma A.2, there exists a subsequence $(s^{n_{m}})$ which is componentwise convergent. Let $s$ denote its limit. The inequalities $a_{i}\leq s^{n_{m}}_{i}\leq b_{i}$ which hold for all $i$ and $m$, clearly carry over to the limit when $m$ goes to infinity, so $s\in[a,b]$. ∎
Lemma A.4.
Let $\{f_{i}:i\in I\}\subset\mathcal{K}_{\infty}$. Assume that both $\{f_{i}\}$ and $\{f_{i}^{-1}\}$ are pointwise equicontinuous. Then for every compact interval $K\subset\mathbb{R}_{+}$ and $\alpha>0$ there is $\beta>0$ such that $|r-\tilde{r}|\geq\alpha$ with $r,\tilde{r}\in K$ implies
$|f_{i}^{-1}(r)-f_{i}^{-1}(\tilde{r})|\geq\beta\mbox{\qquad for all\ }i\in I.$
Proof.
Fix a compact interval $K$ and observe that there exists another compact interval $L$ with $f_{i}^{-1}(K)\subset L$ for all $i\in I$, following from equicontinuity of $\{f_{i}^{-1}\}$ (see the proof of [13, Thm. VI.1] for a precise argument). Now assume to the contrary that there exists $\alpha>0$ such that for every $\beta>0$ there are $r,\tilde{r}\in K$ with $|r-\tilde{r}|\geq\alpha$ and $i\in I$ with $|f_{i}^{-1}(r)-f_{i}^{-1}(\tilde{r})|<\beta$. Let $\beta$ be small enough so that $|f_{i}(r)-f_{i}(\tilde{r})|<\alpha$ whenever $r,\tilde{r}\in L$ with $|r-\tilde{r}|<\beta$ and $i\in I$. Then $|f_{i}^{-1}(r)-f_{i}^{-1}(\tilde{r})|<\beta$ implies $|r-\tilde{r}|<\alpha$, a contradiction. ∎
The proof of the following simple lemma is left to the reader.
Lemma A.5.
Let $a\geq c\geq 0$ and $b\geq d\geq 0$ for vectors $a,b,c,d\in\ell^{\infty}_{+}(\mathcal{I})$. Then
$a\oplus b-c\oplus d\leq(a-c)\oplus(b-d).$
Lemma A.6.
[18, Lem. A.6] For any $\rho\in\mathcal{K}_{\infty}$, the following statements hold:
  1. (a)
    There exists $\eta\in\mathcal{K}_{\infty}$ such that $(\mathrm{id}+\rho)^{-1}=\mathrm{id}-\eta$.
  2. (b)
    There are $\rho_{1},\rho_{2}\in\mathcal{K}_{\infty}$ such that $\mathrm{id}+\rho=(\mathrm{id}+\rho_{1})\circ(\mathrm{id}+\rho_{2})$.
Lemma A.7.
[18, Prop. A.16] Let $z:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ be non-decreasing, continuous at zero and satisfy $z(0)=0$, $z(r)>0$ for all $r>0$. Then there exist $z_{1},z_{2}\in\mathcal{K}$, smooth on $(0,\infty)$, such that $z_{1}(r)\leq z(r)\leq z_{2}(r)$ for all $r\geq 0$. If $z(r)\rightarrow\infty$ as $r\rightarrow\infty$, then $z_{1},z_{2}$ can be chosen in $\mathcal{K}_{\infty}$.
Lemma A.8.
Let $T:\ell^{\infty}_{+}\rightarrow\ell^{\infty}_{+}$ be a monotone operator with $T(0)=0$. If the system $\Sigma(T)$ is UGS and GATT, then it is UGAS.
Proof.
Fix $r,\varepsilon>0$ and choose $n$ such that $\|T^{k}(r\mathds{1})\|\leq\varepsilon$ for all $k\geq n$. Then $\|s\|\leq r$ implies $s\leq r\mathds{1}$, and thus $T^{k}(s)\leq T^{k}(r\mathds{1})\leq\varepsilon\mathds{1}$ for all $k\geq n$. Hence, $\Sigma(T)$ is UGATT. By [19, Thm. 4.2], UGS and UGATT together imply UGAS. ∎
Lemma A.9.
Let $T$ be a monotone operator on $\ell^{\infty}_{+}(\mathcal{I})$. If $s\in\Psi(T)$ for some $s\in\ell^{\infty}_{+}(\mathcal{I})$, then $T(s),s]\subset\Psi(T)$.
Proof.
If $T(s)\leq s^{\prime}\leq s$, then monotonicity implies $T(s^{\prime})\leq T(s)\leq s^{\prime}$. ∎
Lemma A.10.
For a set $A\subset\ell^{\infty}_{+}(\mathcal{I})$, the following statements are equivalent:
  1. (a)
    The set $A$ is uniformly cofinal.
  2. (b)
    The set $A$ is cofinal and there exists a sequence $(\hat{s}^{n})_{n\in\mathbb{N}}$ in $A\cap\mathrm{int}(\ell^{\infty}_{+}(\mathcal{I}))$ with $\hat{s}^{n}\rightarrow 0$.
Proof.
(a) $\Rightarrow$ (b): Clearly, if $A$ is uniformly cofinal, then it is cofinal. Now, take a sequence $(s^{n})_{n\in\mathbb{N}}$ in $\mathrm{int}(\ell^{\infty}_{+}(\mathcal{I}))$, converging to the origin. By assumption, there exists another sequence $(\hat{s}^{n})_{n\in\mathbb{N}}$ with $\hat{s}^{n}\in A$ and $\hat{s}^{n}\geq s^{n}$ such that $\|\hat{s}^{n}\|\leq\varphi(\|s^{n}\|)$. Since $\varphi(\|s^{n}\|)$ converges to zero, the sequence $\hat{s}^{n}$ must converge to the origin. Moreover, $s^{n}\in\mathrm{int}(\ell^{\infty}_{+}(\mathcal{I}))$ implies $\hat{s}^{n}\in\mathrm{int}(\ell^{\infty}_{+}(\mathcal{I}))$.
(b) $\Rightarrow$ (a): From the assumption, it easily follows that there exist two-sided sequences $(r_{n})_{n\in\mathbb{Z}}$ in $(0,\infty)$ and $(\hat{s}^{n})_{n\in\mathbb{Z}}$ in $A$ satisfying
  • $r_{n}\mathds{1}\leq\hat{s}^{n}\leq r_{n+1}\mathds{1}$ for all $n\in\mathbb{Z}$.
  • $(r_{n})_{n\in\mathbb{Z}}$ is strictly increasing.
  • $r_{n}\rightarrow 0$ as $n\rightarrow$-$\infty$ and $r_{n}\rightarrow\infty$ as $n\rightarrow\infty$.
Let us define a function $\tilde{\varphi}$$\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ by $\tilde{\varphi}(0):=0$ and
$\tilde{\varphi}(r):=r_{n+1}\mbox{\ with\ }n=\min\{m:r\leq r_{m}\}\mbox{\quad for all\ }r>0.$
For an arbitrary $0\neq s\in\ell^{\infty}_{+}$, there exists a minimal $n\in\mathbb{Z}$ with $\|s\|\leq r_{n}$, implying $s\leq\|s\|\mathds{1}\leq r_{n}\mathds{1}\leq\hat{s}^{n}\leq r_{n+1}\mathds{1}$. This, in turn, implies $\|\hat{s}^{n}\|\leq r_{n+1}=\tilde{\varphi}(\|s\|)$. We observe that $\tilde{\varphi}(r)\rightarrow 0$ as $r\rightarrow 0$. This implies the existence of a $\mathcal{K}_{\infty}$-function $\varphi$ with $\tilde{\varphi}\leq\varphi$. For instance, we can define $\varphi$ by linear interpolation between the values of $\tilde{\varphi}$. This completes the proof. ∎

### v6

@@
 Moreover, we can conclude that the $\oplus$-MBI property for $\Gamma_{\rho}$ implies that each finite sub-network admits a path of strict decay.
-In Section 4, we apply our results to three classes of gain operators.
+In Section 4, we apply our results to four classes of gain operators.
 Subsection 4.1: We study max-type gain operators, which have the property that they preserve the $\oplus$-operation on $\ell^{\infty}_{+}(\mathcal{I})$, i.e.
@@
 We prove our results assuming the considerably weaker property of uniform continuity.
+Subsection 4.4: We study gain operators which are uniformly continuous on their whole domain of definition.
+Using ideas from Subsection 2.4, we provide conditions under which such an operator satisfies the $\oplus$-MBI property.
 In Section 5, we summarize the main contributions of the paper and present a list of open questions and problems for future work.
@@
 We also show:
-$\mbox{If }A\subset\ell^{\infty_{+}}\mbox{ is cofinal, then }\Gamma(A)\mbox{ is cofinal}.$ (11)
-To this end, let $s\in\ell^{\infty}_{+}$ and choose $\hat{s}\in A$ with $\hat{s}\geq(\xi\circ\eta)^{-1}(\|s\|)\mathds{1}$.
+$\mbox{If }A\subset\ell^{\infty}_{+}(\mathcal{I})\mbox{ is cofinal, then }\Gamma(A)\mbox{ is cofinal}.$ (11)
+To this end, let $s\in\ell^{\infty}_{+}(\mathcal{I})$ and choose $\hat{s}\in A$ with $\hat{s}\geq(\xi\circ\eta)^{-1}(\|s\|)\mathds{1}$.
 Then, for each index $i$, we find $j\in\mathcal{I}_{i}$ and thus have the estimate $\Gamma_{i}(\hat{s})\geq\Gamma_{i}((\xi\circ\eta)^{-1}(\|s\|)\mathds{1})\geq\|s\|$.
@@
 4 Particular classes of gain operators
-In this section, we apply our results and techniques to three particular classes of gain operators.
+In this section, we apply our results and techniques to four particular classes of gain operators.
 4.1 Max-type operators
@@
 The proof of this result uses different methods, however, and in particular uses the subadditivity of MAFs for the path construction.
+4.4 Uniformly continuous operators
+From Proposition 2.1, we know that any gain operator is uniformly continuous on bounded sets.
+In this subsection, we will show that the assumption of uniform continuity on the whole domain of definition allows for stronger conclusions.
+In particular, under this assumption we will obtain more clarity concerning the relation between the two necessary conditions for the existence of a path of strict decay, namely UGAS of $\Sigma(\Gamma_{\rho})$ and the $\oplus$-MBI property of $\Gamma_{\rho}$.
+Lemma 4.2.
+Given a uniformly continuous gain operator $\Gamma$, for each $n\in\mathbb{N}$ there exists a $\varphi\in\mathcal{K}_{\infty}$ such that for all $\varepsilon>0$ and $s\in\ell^{\infty}_{+}(\mathcal{I})$, $i\in\mathcal{I}$ the following holds: $\Gamma_{j}(s)\geq s_{j}-\varphi(\varepsilon)$ for all $j\in\mathcal{N}^{-}_{i}(n-1)$ implies $\Gamma_{i}^{n}(s)\geq s_{i}-\varepsilon$.
+Proof.
+We use the characterization of uniform continuity in terms of comparison functions. Let $\gamma\in\mathcal{K}_{\infty}$ be chosen such that
+$\|\Gamma(s^{1})-\Gamma(s^{2})\|\leq\gamma(\|s^{1}-s^{2}\|)$
+holds for all $s^{1},s^{2}\in\ell^{\infty}_{+}(\mathcal{I})$.
+Given $n\in\mathbb{N}$ and $\varepsilon>0$, we then choose $\varepsilon_{i},\delta_{i}>0$, $1\leq i\leq n-1$ as in the proof of Lemma 2.1, now using the function $\gamma$.
+That is, we put $\delta_{n}:=\varepsilon$, $\varepsilon_{n-1}:=\delta_{n}/2$ and continue recursively by
+$\delta_{l}:=\gamma^{-1}(\varepsilon_{l}),\quad\varepsilon_{l-1}:=\frac{\delta_{l}}{2}$
+We do this for l=n1l=n-1 down to $l=2$. Eventually, we choose $\delta_{1}=:\delta$ as
+$\delta_{1}:=\min\left\{\gamma^{-1}(\varepsilon_{1}),\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{n-1}\right\}.$
+Without loss of generality, we can actually assume that $\gamma^{-1}<\mathrm{id}$, which implies $\delta_{1}=\gamma^{-1}(\varepsilon_{1})$.
+This obviously defines a $\mathcal{K}_{\infty}$-function $\varphi$, only depending on $n$, such that $\delta=\varphi(\varepsilon)$.
+The rest of the proof is completely analogous to that of Lemma 2.1.
+∎
+What makes Lemma 4.2 superior to Lemma 2.1 is that its statement is not only meaningful for $\varepsilon$ close to zero. The following proposition shows how this can be used.
+Proposition 4.5.
+Let $\Gamma$ be a uniformly continuous gain operator. Then, under each of the following assumptions, $\Gamma$ satisfies the $\oplus$-MBI property:
+  1. (i)
+    $|\mathcal{I}|<\infty$ and $\Sigma(\Gamma)$ is GATT.
+  2. (ii)
+    $\Sigma(\Gamma)$ is UGAS with a $\mathcal{K}\mathcal{L}$-function of the form $\beta(r,n)=r\gamma(n)$, $\gamma\in\mathcal{L}$.
+Proof.
+(i) Assuming that $\Sigma(\Gamma)$ is GATT, we first show that the following strong version of the uniform NJI condition holds: There exists $\varphi\in\mathcal{K}_{\infty}$ such that for all $s\geq 0$, $i\in\mathcal{I}$ and $\varepsilon>0$:
+$s_{i}\geq\varepsilon\quad\Rightarrow\quad\exists j\in\mathcal{N}^{-}_{i}:\ s_{j}\geq\varphi(\varepsilon)\wedge\Gamma_{j}(s)<s_{j}.$
+To this end, we pick $m\in\mathbb{N}$ such that $\mathcal{N}^{-}_{i}=\mathcal{N}^{-}_{i}(m+k)$ for all $i\in\mathcal{I}$ and $k\in\mathbb{N}_{0}$, which is possible due to the finiteness of $\mathcal{I}$.
+Then, by Lemma 4.2, we can choose $\tilde{\varphi}=\tilde{\varphi}(m)\in\mathcal{K}_{\infty}$ accordingly.
+We put $\varphi(r):\equiv\tilde{\varphi}(r/2)$ and assume towards a contradiction that there are $s\geq 0$, $i\in\mathcal{I}$ and $\varepsilon>0$ with $s_{i}\geq\varepsilon$, but $s_{j}<\varphi(\varepsilon)$ or $\Gamma_{j}(s)\geq s_{j}$ for all $j\in\mathcal{N}^{-}_{i}$.
+Observe that in both cases $\Gamma_{j}(s)\geq s_{j}-\varphi(\varepsilon)$.
+Hence, Lemma 4.2 implies $\Gamma_{i}^{n}(s)\geq s_{i}-\varepsilon/2$ for all $n\geq m$.
+We then choose $n\geq m$ large enough so that $\|\Gamma^{n}(s)\|<\varepsilon/2$, leading to the contradiction
+$\varepsilon\leq s_{i}\leq\Gamma_{i}^{n}(s)+\frac{\varepsilon}{2}\leq\|\Gamma^{n}(s)\|+\frac{\varepsilon}{2}<\varepsilon$
+To complete the proof, let $s\leq r\mathds{1}\oplus\Gamma(s)$ for some $s\in\ell^{\infty}_{+}(\mathcal{I})$ and $r\geq 0$.
+Assume to the contrary that $s_{i}\geq\varphi^{-1}(r)+\delta$ for some $i\in\mathcal{I}$ and a small $\delta>0$.
+Then, according to the strong version of the uniform NJI condition, shown above, there exists $j\in\mathcal{N}^{-}_{i}$ with $s_{j}\geq\varphi(\varphi^{-1}(r)+\delta)>r$ and $\Gamma_{j}(s)<s_{j}$.
+The latter implies $s_{j}\leq r$, since otherwise $s_{j}\leq\max\{r,\Gamma_{j}(s)\}$ leads to $s_{j}=\Gamma_{j}(s)<s_{j}$.
+Hence, we obtain a contradiction, namely $r<s_{j}\leq r$.
+Altogether, we see that $s_{i}<\varphi^{-1}(r)+\delta$.
+Since this holds for all $i$ and $\delta$, we obtain $\|s\|\leq\varphi^{-1}(r)$, showing that $\Gamma$ satisfies the $\oplus$-MBI property.
+(ii) We fix an $m\in\mathbb{N}$ with $\gamma(m)<1/4$, choose $\tilde{\varphi}=\tilde{\varphi}(m)\in\mathcal{K}_{\infty}$ according to Lemma 4.2, and put $\varphi(r):\equiv\tilde{\varphi}(r/2)$.
+As above, we show a strong version of the uniform NJI condition: for all $s\geq 0$ and $i\in\mathcal{I}$:
+$s_{i}\geq\frac{\|s\|}{2}\quad\Rightarrow\quad\exists j\in\mathcal{N}^{-}_{i}(m):\ s_{j}\geq\varphi\left(\frac{\|s\|}{2}\right)\wedge\Gamma_{j}(s)<s_{j}.$
+Assume towards a contradiction that there are $s\geq 0$ and $i\in\mathcal{I}$ with $s_{i}\geq\|s\|/2$, but $s_{j}<\varphi(\|s\|/2)$ or $\Gamma_{j}(s)\geq s_{j}$ for all $j\in\mathcal{N}^{-}_{i}(m)$; in both cases, $\Gamma_{j}(s)\geq s_{j}-\varphi(\|s\|/2)$ holds.
+Hence, Lemma 4.2 implies $\Gamma_{i}^{m}(s)\geq s_{i}-\|s\|/4$.
+Then
+$\frac{\|s\|}{2}\leq s_{i}\leq\Gamma_{i}^{m}(s)+\frac{\|s\|}{4}\leq\|\Gamma^{m}(s)\|+\frac{\|s\|}{4}\leq\|s\|\gamma(m)+\frac{\|s\|}{4}<\frac{\|s\|}{2},$
+which is a contradiction.
+Now, towards a contradiction, assume that $s\leq r\mathds{1}\oplus\Gamma(s)$ and $\|s\|>2\varphi^{-1}(r)$.
+Pick any $i\in\mathcal{I}$ with $s_{i}\geq\|s\|/2$.
+Then, by the above, there exists $j\in\mathcal{N}^{-}_{i}(m)$ with $s_{j}\geq\varphi(\|s\|/2)$ and $\Gamma_{j}(s)<s_{j}$.
+The latter implies $s_{j}\leq r$.
+Altogether, $r<\varphi(\|s\|/2)\leq s_{j}\leq r$, which is a contradiction.
+Hence, we have shown that $s\leq r\mathds{1}\oplus\Gamma(s)$ implies $\|s\|\leq 2\varphi^{-1}(r)$.
+∎
+Remark 4.7.
+We can compare part (i) of the proposition with Theorem 4.5, where we showed under the assumption of uniform continuity of the MAFs $\mu_{i}$ that the NJI condition for $\Gamma_{\rho}$ implies the $\oplus$-MBI property of $\Gamma_{\rho^{\prime}}$, for some $\rho^{\prime}$ smaller than $\rho$.
+In both results, we needed a uniform continuity assumption which is not automatically satisfied under our general axioms for gain operators.
+This lets us suspect that without such an assumption UGAS does not always imply the $\oplus$-MBI property.
+Part (ii) has a stronger conclusion than Proposition 2.11, which also used the assumption that $\Sigma(\Gamma)$ is UGAS with $\beta(r,n)=r\gamma(n)$.
+Additionally, Proposition 2.11 used Assumption 2.1, and only showed that $\Psi(\Gamma)$ is cofinal.
+Example 4.1.
+As an example, consider a subadditive and homogeneous gain operator $\Gamma$. Such an operator is uniformly continuous. Indeed, $0\leq s^{1}\leq s^{2}$ implies
+$\Gamma(s^{2})-\Gamma(s^{1})=\Gamma(s^{1}+(s^{2}-s^{1}))-\Gamma(s^{1})\leq\Gamma(s^{2}-s^{1})$
+and thus, by continuity at $s=0$, for every $\varepsilon>0$ we find $\delta>0$ such that $\|s^{2}-s^{1}\|\leq\delta$ implies
+$\|\Gamma(s^{2})-\Gamma(s^{1})\|\leq\|\Gamma(s^{2}-s^{1})\|=\|\Gamma(s^{2}-s^{1})-\Gamma(0)\|\leq\varepsilon.$
+For general $s^{1},s^{2}\in\ell^{\infty}_{+}(\mathcal{I})$ (not necessarily ordered), we can use the same trick as in Subsection 4.3.2 used for MAFs to prove the corresponding implication.
+As we have shown in Subsection 4.2, UGAS of $\Sigma(\Gamma)$ implies UGES, i.e.
+UGAS with $\beta$ of the form $\beta(r,n)\equiv Mr\mathrm{e}^{-\gamma n}$.
+Hence, assumption (ii) of Proposition 4.5 is satisfied.
+This yields another explanation for the implication “(a) $\Rightarrow$ (b)” in Proposition 4.2.
 5 Summary and open questions